### v3

# Commutation relations for two-sided radial SLE

Abstract
We study the commutation relation for 2-radial SLE in the unit disc starting from two boundary points. We follow the framework introduced by Dubédat [12]. Under an additional requirement of the interchangeability of the two curves, we classify all locally commuting 2-radial SLEκ for $\kappa\in(0,8)$: it is either a two-sided radial SLEκ with spiral of constant spiraling rate or a chordal SLEκ weighted by a power of the conformal radius of its complement. Namely, for fixed $\kappa$ and starting points, we have exactly two one-parameter continuous families of locally commuting 2-radial SLE. We also discuss the semiclassical limit of the commutation relation as $\kappa\to 0$. In particular, we show that the limit for the second family with an appropriately chosen power of conformal radius is a chord that minimizes a modified chordal Loewner energy, which is unique only when the endpoints are not antipodal.
Keywords: Schramm–Loewner evolution, radial Loewner chain, commutation relation, resampling property.
MSC: 60J67.
1 Introduction
1.1 Background on radial SLE
In 1999, O. Schramm [37] introduced the Schramm-Loewner evolution (SLE) as the non-self-crossing random curve driven by a multiple of Brownian motion using Loewner’s transform. This definition is motivated by a quest to describe mathematically the random interfaces in two-dimensional critical lattice models, which satisfy the conformal invariance and the domain Markov property. These properties impose that the chordal Loewner driving function of such a random interface has to be a multiple of Brownian motion, hence justifying the definition of chordal SLE. Indeed, SLEs are proved to be the scaling limit of interfaces in many conformally invariant statistical mechanics models, e.g., [38, 27, 39, 10, 42, 9], and play a central role in random conformal geometry.
However, to characterize the other natural variant — radial SLE — one needs an additional condition on the reflection symmetry. As we are mainly concerned with radial SLE in the present article, let us briefly describe its definition and characterization. We will describe the radial Loewner chain in $\mathbb{D}=\{z\in\mathbb{C}\,|\,|z|<1\}$ targeting at 0. Radial Loewner chain in other simply connected domain $D\subset\mathbb{C}$ targeting at an interior point $z_{0}\in D$ is defined via a conformal map from $\mathbb{D}$ onto $D$ sending 0 to $z_{0}$.
Conformal radius and capacity. For any compact subset $K$ (not necessarily connected) of $\overline{\mathbb{D}}$ such that $\mathbb{D}\setminus K$ is simply connected and contains 0, we let $g_{K}$ be the unique conformal map $\mathbb{D}\setminus K\to\mathbb{D}$ such that $g_{K}(0)=0$ and $g^{\prime}_{K}(0)>0$ (called the radial mapping-out function of $K$). The conformal radius of $\mathbb{D}\setminus K$ is
$\mathrm{CR}(\mathbb{D}\setminus K):=(g^{\prime}_{K}(0))^{-1}.$
The capacity of $K$ is
$\operatorname{cap}(K)=\log g^{\prime}_{K}(0)=-\log\mathrm{CR}(\mathbb{D}\setminus K).$
Radial Loewner chain. For $\theta\in[0,2\pi)$, suppose $\eta:[0,T]\to\overline{\mathbb{D}}$ is a continuous non-self-crossing curve such that $\eta_{0}=\mathrm{e}^{\mathfrak{i}\theta}$ and $\eta_{(0,T)}\subset\mathbb{D}\setminus\{0\}$. Let $D_{t}$ be the connected component of $\mathbb{D}\setminus\eta_{[0,t]}$ containing the origin. Let $g_{t}:D_{t}\to\mathbb{D}$ be the unique conformal map with $g_{t}(0)=0$ and $g_{t}^{\prime}(0)>0$. We say that the curve is parameterized by capacity if $g_{t}^{\prime}(0)=\mathrm{e}^{t}$. Then $g_{t}$ satisfies the radial Loewner equation
$\partial_{t}g_{t}(z)=g_{t}(z)\frac{\mathrm{e}^{\mathfrak{i}\xi_{t}}+g_{t}(z)}{\mathrm{e}^{\mathfrak{i}\xi_{t}}-g_{t}(z)},\quad g_{0}(z)=z,$ (1.1)
where $t\mapsto\xi_{t}\in\mathbb{R}/2\pi\mathbb{Z}$ is continuous and called the driving function of $\eta$. We note that if $z=\mathrm{e}^{\mathfrak{i}V_{0}}\in\partial\mathbb{D}$, then taking a continuous branch $V_{t}:=\arg g_{t}(\mathrm{e}^{\mathfrak{i}\theta})$, we have
$\partial_{t}V_{t}(z)=\cot((V_{t}-\xi_{t})/2).$ (1.2)
Characterization of radial SLE. Consider a family $(\mathbb{P}_{\theta})_{\theta\in\mathbb{R}/2\pi\mathbb{Z}}$ of probability distributions on curves $\eta:[0,\infty)\to\overline{\mathbb{D}}$ with $\eta_{0}=\mathrm{e}^{\mathfrak{i}\theta}$ and parametrized by capacity that satisfies the following properties:
  1. 1.
    (Conformal invariance) For all $a\in\mathbb{R}/2\pi\mathbb{Z}$, let $\rho_{a}(z)=\mathrm{e}^{\mathfrak{i}a}z$ be the rotation map $\mathbb{D}\to\mathbb{D}$. For all $a,\theta\in\mathbb{R}/2\pi\mathbb{Z}$, the pullback measure $\rho_{a}^{*}\,\mathbb{P}_{\theta}=\mathbb{P}_{\theta-a}$. From this, we may extend the definition of $\mathbb{P}_{\theta}$ to any simply connected domain $D$ with an interior marked point $z_{0}$ by pulling back using a uniformizing conformal map $D\to\mathbb{D}$ sending $z_{0}$ to 0.
  2. 2.
    (Domain Markov property) For any $t>0$, $\theta\in\mathbb{R}/2\pi\mathbb{Z}$, let $\eta\sim\mathbb{P}_{\theta}$. Conditioning on $\eta|_{[0,t]}$, $g_{t}(\eta_{[t,\infty)})\sim\mathbb{P}_{\xi_{t}}$ where $\xi_{\cdot}$ is the driving function of $\eta$. See Figure 1.
  3. 3.
    (Reflection symmetry) Let $\iota:z\mapsto\overline{z}$ be the complex conjugation, then $\mathbb{P}_{\theta}\sim\iota^{*}\mathbb{P}_{-\theta}$.
Then there exists $\kappa\geq 0$ such that for all $\theta$, the driving function of $\eta\sim\mathbb{P}_{\theta}$ is $t\mapsto\theta+\sqrt{\kappa}B_{t}$ modulo $2\pi\mathbb{Z}$, where $B_{\cdot}$ is the standard Brownian motion. This follows from the fact that $(\sqrt{\kappa}B)_{\kappa\geq 0}$ are the only continuous Lévy processes $W$ that have the same law under the transformation $W\mapsto-W$. In this case, as $t\to\infty$, $\eta_{t}\to 0$ almost surely. The distribution $\mathbb{P}_{\theta}$ is the law of the radial $\operatorname{SLE}_{\kappa}$ in $\mathbb{D}$ starting from $\mathrm{e}^{\mathfrak{i}\theta}$. We also call it radial $\operatorname{SLE}_{\kappa}$ in $(\mathbb{D};\mathrm{e}^{\mathfrak{i}\theta};0)$ for short.
Figure 1: Domain Markov property of radial SLE.
The third assumption on the reflection symmetry is natural for conformally invariant and achiral statistical mechanics models. However, one may also wonder what happens without the reflection symmetry, this means that the law of the driving function is no longer required to be invariant with respect to $W\mapsto-W$. From the classification of continuous Lévy processes, we obtain that there exists $\kappa\geq 0$ and $\mu\in\mathbb{R}$ such that for all $\theta$, the driving function of $\eta\sim\mathbb{P}_{\theta}$ is
$t\mapsto\theta+\sqrt{\kappa}B_{t}+\mu t\qquad\text{ (mod) }2\pi\mathbb{Z}.$
The curve generated is called radial SLEκ with spiraling rate $\mu$ starting from $\mathrm{e}^{\mathfrak{i}\theta}$ and we denote it as radial $\operatorname{SLE}_{\kappa}^{\mu}$ in $(\mathbb{D};\mathrm{e}^{\mathfrak{i}\theta};0)$. It was shown in [33] that almost surely, $\eta_{t}\to 0$ as $t\to\infty$ for all $\kappa>0$ and $\mu\in\mathbb{R}$.
1.2 Locally commuting 2-radial SLE
The random radial curve satisfying conformal invariance and domain Markov property are easily characterized thanks to the fact all simply connected domains (excluding $\mathbb{C}$) with one marked interior point and one marked boundary point are conformally equivalent. If we consider the radial Loewner chain of two curves growing from two distinct boundary points $\mathrm{e}^{\mathfrak{i}\theta_{1}},\mathrm{e}^{\mathfrak{i}\theta_{2}}\in\partial\mathbb{D}$, then we have a one-dimensional moduli space of the boundary data that we need to take into account.
The goal of this work is to give an axiomatic characterization of locally commuting 2-radial SLEs. One example of such curves of particular interest, is the two-sided radial SLE analyzed in [25, 16, 28, 17] which arises as the chordal SLE in $\mathbb{D}$ connecting two boundary points conditioned on passing through 0.
More precisely, locally commuting 2-radial SLE is a family of local laws $\mathbb{P}_{(\theta_{1},\theta_{2})}$ of the pair of continuous non-self-crossing curves $(\eta^{(1)},\eta^{(2)})$ in $\mathbb{D}$ starting from all possible choices of $\mathrm{e}^{\mathfrak{i}\theta_{1}},\mathrm{e}^{\mathfrak{i}\theta_{2}}\in\partial\mathbb{D}$ under the following axioms.
We first parameterize $\eta^{(1)},\eta^{(2)}$ by their intrinsic radial capacities and consider the associated radial Loewner chains $g_{\cdot}^{(1)}$ and $g_{\cdot}^{(2)}$ respectively. For all $\mathbf{t}=(t_{1},t_{2})\in\mathbb{R}_{>0}^{2}$, let $g_{\mathbf{t}}$ be the radial mapping-out function of $\eta_{\mathbf{t}}:=\eta^{(1)}_{[0,t_{1}]}\cup\eta^{(2)}_{[0,t_{2}]}$.
  • (CI)
    Conformal invariance: For all $a\in\mathbb{R}/2\pi\mathbb{Z}$ and $\theta_{1},\theta_{2}\in\mathbb{R}/2\pi\mathbb{Z}$, the pullback measure $\rho_{a}^{*}\,\mathbb{P}_{(\theta_{1},\theta_{2})}=\mathbb{P}_{(\theta_{1}-a,\theta_{2}-a)}$. From this, we may extend the definition of $\mathbb{P}_{\theta_{1},\theta_{2}}$ to any simply connected domain $D$ with an interior marked point $z_{0}$ by pulling back using a uniformizing conformal map $D\to\mathbb{D}$ sending $z_{0}$ to 0.
  • (DMP)
    Domain Markov property: Conditioning on $\eta_{\mathbf{t}}$,
    $\left(g_{\mathbf{t}}(\eta^{(1)}\setminus\eta^{(1)}_{[0,t_{1}]}),g_{\mathbf{t}}(\eta^{(2)}\setminus\eta^{(2)}_{[0,t_{2}]})\right)\sim\mathbb{P}_{\left(\theta^{(1)}_{\mathbf{t}},\theta^{(2)}_{\mathbf{t}}\right)}$,
    where $\left(\exp(\mathfrak{i}\theta^{(1)}_{\mathbf{t}}),\exp(\mathfrak{i}\theta^{(2)}_{\mathbf{t}})\right)=\left(g_{\mathbf{t}}(\eta^{(1)}_{t_{1}}),g_{\mathbf{t}}(\eta^{(2)}_{t_{2}})\right)$.
  • (MARG)
    Marginal laws: There exists $\kappa>0$ such that the marginal local law of $\eta^{(j)}$ is “absolutely continuous” with respect to an $\operatorname{SLE}_{\kappa}$ for $j=1,2$.
  • (INT)
    Interchangeability condition: Let $\tau$ be the map which swaps $\eta^{(1)}$ and $\eta^{(2)}$. We have $\mathbb{P}_{(\theta_{1},\theta_{2})}=\tau^{*}\mathbb{P}_{(\theta_{2},\theta_{1})}$.
We note that, by “radial” we mean that the conformal mapping-out functions are normalized at the interior point $0\in\mathbb{D}$ and the precise axioms are written in neighborhoods of the starting points of the curves, see Section 2.1. In particular, they do not necessarily imply that the curves end at 0.
We also note that since $g_{\mathbf{t}}$ is parametrized by two times $t_{1},t_{2}$, the condition (DMP) implies that $g_{\mathbf{t}}$ may be realized by first mapping out $\eta^{(1)}_{[0,t_{1}]}$ using $g_{t_{1}}^{(1)}$, then mapping out $g_{t_{1}}^{(1)}(\eta^{(2)}_{[0,t_{2}]})$, or vice versa. The image has the same law regardless of the order in which we map out the curves. This observation gives a commutation relation on the infinitesimal generators of the two curves (Proposition 2.2). Therefore, we call the family of the laws $(\mathbb{P}_{(\theta_{1},\theta_{2})})$ as an interchangeable and locally commuting 2-radial SLEκ.
These conditions and our analysis are very close to the commutation relation of SLE studied by Dubédat in [12], which focuses on classifying all locally commuting chordal SLEs. Dubédat also derived the commutation relation in the radial setting in terms of the generators of radial SLE. One of our contributions is to find all solutions to the radial commutation relation and identify all interchangeable and locally commuting 2-radial SLEκ. It is also natural to use SLE/GFF couplings, particularly, [33], to find examples of commuting $2$-radial SLEs. See Remark 3.9. Nevertheless, our approach is to give an axiomatic characterization by studying the BPZ equations satisfied by the partition functions, and it is unclear to the authors that all commuting $2$-radial SLEs can be coupled to GFF a priori.
1.3 Main result: A classification
We classify all locally commuting 2-radial SLE with the interchangeability condition (INT). Similar to the analysis in [12], we also show that the conditions (CI), (DMP), and (MARG) imply that there exists $\mathcal{Z}:\{(\theta_{1},\theta_{2})\in\mathbb{R}^{2}\,|\,\theta_{1}<\theta_{2}<\theta_{1}+2\pi\}\to\mathbb{R}_{>0}$, called partition function which encodes the family of distributions $\mathbb{P}_{(\theta_{1},\theta_{2})}$. See Section 2 for more details.
Theorem 1.1.
Fix $\kappa\in(0,8)$. Suppose $\mathcal{Z}$ is the partition function for an interchangeable and locally commuting 2-radial $\operatorname{SLE}_{\kappa}$. Then $\mathcal{Z}$ is one of the following two functions:
  1. 1.
    There exists $\mu\in\mathbb{R}$ such that, up to a multiplicative constant, $\mathcal{Z}$ is the same as
    $\mathcal{G}_{\mu}(\theta_{1},\theta_{2})=\left(\sin\left((\theta_{2}-\theta_{1})/2\right)\right)^{2/\kappa}\exp\left(\frac{\mu}{\kappa}(\theta_{1}+\theta_{2})\right).$ (1.3)
    In this case, the law of the corresponding locally commuting 2-radial $\operatorname{SLE}_{\kappa}$ is the same as two-sided radial $\operatorname{SLE}_{\kappa}$ with spiraling rate $\mu$, see Section 3.2 and Figure 2.
  2. 2.
    There exists $\alpha<1-\kappa/8$ such that, up to a multiplicative constant, $\mathcal{Z}$ is the same as
    $\mathcal{Z}_{\alpha}(\theta_{1},\theta_{2})=\left(\sin\left((\theta_{2}-\theta_{1})/2\right)\right)^{(\kappa-6)/\kappa}\phi_{\alpha}\left(\left(\sin\left((\theta_{2}-\theta_{1})/4\right)\right)^{2}\right),$ (1.4)
    where $\phi_{\alpha}$ is the unique solution to the following Euler’s hypergeometric differential equation
    $\begin{cases}u(1-u)\phi^{\prime\prime}(u)-\frac{3\kappa-8}{2\kappa}(2u-1)\phi^{\prime}(u)+\frac{8\alpha}{\kappa}\phi(u)=0,\quad u\in(0,1);\\ \phi(1/2)=1,\quad\phi^{\prime}(1/2)=0.\end{cases}$ (1.5)
    In this case, the law of the corresponding locally commuting 2-radial $\operatorname{SLE}_{\kappa}$ is the same as $\gamma$ chordal $\operatorname{SLE}_{\kappa}$ in $\mathbb{D}$ weighted by $\mathrm{CR}(\mathbb{D}\setminus\gamma)^{-\alpha}$, where $\mathrm{CR}(\mathbb{D}\setminus\gamma)$ denotes the conformal radius of the connected component of $\mathbb{D}\setminus\gamma$ containing the origin, see Section 3.4.
The first class above — two-sided radial SLE with spiral — also appeared in [33] in the context of SLE/GFF coupling. We will discuss other related literature in Section 1.6. The authors are unaware of studies on the second class of curves.
Both $\mathcal{G}_{\mu}$ in (1.3) and $\mathcal{Z}_{\alpha}$ in (1.4) are solutions to radial Belavin–Polyakov–Zamolodchikov (BPZ) equations (or called BPZ–Cardy equations derived in [23, 24], see also [4, 12] for the BPZ equation for multichordal SLE):
$\frac{\kappa}{2}\frac{\partial_{11}\mathcal{Z}}{\mathcal{Z}}+\cot\left((\theta_{2}-\theta_{1})/2\right)\frac{\partial_{2}\mathcal{Z}}{\mathcal{Z}}-\frac{(6-\kappa)/(4\kappa)}{\left(\sin((\theta_{2}-\theta_{1})/2)\right)^{2}}$ $=F,$
$\frac{\kappa}{2}\frac{\partial_{22}\mathcal{Z}}{\mathcal{Z}}-\cot\left((\theta_{2}-\theta_{1})/2\right)\frac{\partial_{1}\mathcal{Z}}{\mathcal{Z}}-\frac{(6-\kappa)/(4\kappa)}{\left(\sin((\theta_{2}-\theta_{1})/2)\right)^{2}}$ $=F,$
where $\partial_{i}$ is the partial derivative with respect to $\theta_{i}$ and $F$ is the constant:
$F$ $=\frac{\mu^{2}-3}{2\kappa}\geq-\frac{3}{2\kappa},\quad$ $\text{when }\mathcal{Z}=\mathcal{G}_{\mu},$
$F$ $=\frac{(6-\kappa)(\kappa-2)}{8\kappa}-\alpha>-\frac{3}{2\kappa},\quad$ $\text{when }\mathcal{Z}=\mathcal{Z}_{\alpha}.$
In general, the partition function $\mathcal{Z}_{\alpha}$ in (1.4) involves hypergeometric function and is not explicit; but in the following two special cases, such function has a simple explicit expression:
  • When $\kappa=4$, $\mathcal{Z}_{\alpha}$ can be written as trigonometric functions and hyperbolic functions in $\theta=\theta_{2}-\theta_{1}$, see Remark 2.6.
  • When $\alpha=\alpha_{1}(\kappa)$ which is the one-arm exponent for the conformal loop ensemble, $\mathcal{Z}_{\alpha_{1}(\kappa)}$ can be written as trigonometric functions in $\theta=\theta_{2}-\theta_{1}$, see Remark 2.7.
Figure 2: Simulation by Minjae Park. Two-sided radial SLE around the origin with $\kappa=0.01$ and $\mu=5$ in the left panel. Two-sided radial SLE around the origin with $\kappa=2$ and $\mu=5$ in the right panel.
The interchangeability condition (INT) is natural, given the reversibility of chordal SLE. It also simplifies our classification and in particular, the notation, see Section 2.3 and, in particular, Lemma 2.5. In Remark 3.15, we give an example of locally commuting radial $2$-SLE without (INT). In Proposition 3.16, we show the classification without the condition (INT), where the second one-parameter family of solution (1.4) becomes a two-parameter family.
1.4 Resampling property
Let $D\subset\mathbb{C}$ be a simply connected domain and let $x_{1},x_{2}$ be distinct prime ends of $\partial D$. We denote by $\mathfrak{X}(D;x_{1},x_{2})$ the set of continuous non-self-crossing curves in $D$ from $x_{1}$ to $x_{2}$. For other simply connected domain $(D;x_{1},x_{2})$, chordal SLE in $(D;x_{1},x_{2})$ is a random curve in $\mathfrak{X}(D;x_{1},x_{2})$ defined by mapping chordal SLE in $({\mathbb{H}};0,\infty)$ conformally from $\mathbb{H}$} onto $D$ sending 0 to $x_{1}$ and $\infty$ to $x_{2}$. Let $D\subset\mathbb{D}$ be a simply connected domain containing 0 and let $x_{1},x_{2}$ be distinct prime ends of $\partial D$. We denote by $\mathfrak{X}(D;x_{1},x_{2};0)$ the set of pairs of continuous simple curves $(\eta^{(1)},\eta^{(2)})$ in $D$ such that $\eta^{(1)}$ goes from $x_{1}$ to 0 and $\eta^{(2)}$ goes from $x_{2}$ to 0 and $\eta^{(1)}\cap\eta^{(2)}=\{0\}$. We say that a law on $(\eta^{(1)},\eta^{(2)})\in\mathfrak{X}(D;x_{1},x_{2};0)$ satisfies resampling property if
  • the conditional law of $\eta^{(2)}$ given $\eta^{(1)}$ is a chordal $\operatorname{SLE}_{\kappa}$ in $(D\setminus\eta^{(1)};x_{2},0)$,
  • and the conditional law of $\eta^{(1)}$ given $\eta^{(2)}$ is a chordal $\operatorname{SLE}_{\kappa}$ in $(D\setminus\eta^{(2)};x_{1},0)$.
In Theorem 3.8 we show that for each $\mu\in\mathbb{R}$, two-sided radial $\operatorname{SLE}_{\kappa}$ with spiraling rate $\mu$ satisfies resampling property for $\kappa\in(0,4]$. This result follows directly from the expression of the partition function $\mathcal{G}_{\mu}$. Another proof can be found in [33] using the coupling with GFF. Combining with Theorem 1.1, we find that commutation relation implies resampling property.
Corollary 1.2.
Fix $\kappa\in(0,4]$ and $\theta_{1}<\theta_{2}<\theta_{1}+2\pi$. If an interchangeable and locally commuting 2-radial $\operatorname{SLE}_{\kappa}$ is in $\mathfrak{X}(\mathbb{D};\mathrm{e}^{\mathfrak{i}\theta_{1}},\mathrm{e}^{\mathfrak{i}\theta_{2}};0)$, then it is a two-sided radial $\operatorname{SLE}_{\kappa}$ with spiral, hence, it satisfies the resampling property. The converse is not true: a linear combination of two-sided radial $\operatorname{SLE}_{\kappa}$ with different spiraling rates is in $\mathfrak{X}(\mathbb{D};\mathrm{e}^{\mathfrak{i}\theta_{1}},\mathrm{e}^{\mathfrak{i}\theta_{2}};0)$ and satisfies the resampling property but is not a locally commuting 2-radial $\operatorname{SLE}_{\kappa}$.
1.5 Semiclassical limits of partition functions
[…]
We also discuss the commutation relation when $\kappa=0$, the corresponding deterministic pair of curves, and the semiclassical limit $\kappa\to 0{\scriptstyle+}$ in Section 4. In particular, the semiclassical limits of partition functions in Theorem 1.1 have explicit formulas.
Proposition 1.3.
Fix $\theta_{1}<\theta_{2}<\theta_{1}+2\pi$ and denote $\theta=\theta_{2}-\theta_{1}$.
  • For the partition function $\mathcal{G}_{\mu}$ in (1.3), fix $\mu\in\mathbb{R}$, we have
    $\lim_{\kappa\to 0}\kappa\log\mathcal{G}_{\mu}(\theta_{1},\theta_{2})=2\log\sin\big((\theta_{2}-\theta_{1})/2\big)+\mu(\theta_{1}+\theta_{2}).$ (1.6)
  • For the partition function $\mathcal{Z}_{\alpha}$ in (1.4), if $\alpha\sim-\lambda/\kappa$ for some $\lambda\geq 0$, then the following limit exists
    $\mathcal{U}^{\lambda}(\theta)=$ $\lim_{\kappa\to 0}\kappa\log\mathcal{Z}_{\alpha}(\theta_{1},\theta_{2});$
    and for $\theta\in(0,\pi]$,
    $\mathcal{U}^{\lambda}(\theta)=\mathcal{U}^{\lambda}(2\pi-\theta)=-2\log\sin(\theta/2)+\int_{\theta}^{\pi}\sqrt{2\lambda+4\cot^{2}(u/2)}\,\mathrm{d}u.$ (1.7)
We prove Proposition 1.3 in Section 4. We mention that a similar semiclassical limit of the partition functions for multichordal SLE was considered in [34, 3] and that multiradial $\operatorname{SLE}_{0}$ was considered in [2, 45] (see also Section 1.6).
Proposition 1.4.
We use the same notations as in Proposition 1.3. The function $\mathcal{U}^{\lambda}$ can be expressed as
$\mathcal{U}^{\lambda}(\theta)=-6\log\sin(\theta/2)-\inf_{\gamma}\Big(I(\gamma)-\lambda\log\mathrm{CR}(\mathbb{D}\setminus\gamma)\Big)+C$
where the infimum is attained and taken over all chords connecting $\mathrm{e}^{\mathfrak{i}\theta_{1}}$ and $\mathrm{e}^{\mathfrak{i}\theta_{2}}$, $I(\cdot)$ is the chordal Loewner energy, and $C$ is a normalizing constant only depending on $\lambda$ such that $\mathcal{U}^{\lambda}(\pi)=0$. Moreover, the minimizer is unique when $\theta\neq\pi$ and there are exactly two minimizers when $\theta=\pi$.
See Lemma 4.8 for details and Figure 3 for a simulation of those energy minimizers. A related relation between the semiclassical limit of the partition function for multichordal SLE and the minimizers of the multichordal Loewner energy was proved in [34]. We point out in particular that $\mathcal{U}^{\lambda}$ in (1.7) when $\lambda>0$ is not differentiable at $\theta=\pi$ since the minimizer of $I(\cdot)-\lambda\log\mathrm{CR}(\mathbb{D}\setminus\cdot)$ is non-unique when $\theta=\pi$. Such non-differentiable functions do not appear in earlier literature about the semiclassical limits of SLE partition functions [34, 3, 2, 45].
1.6 Comments and further developments
Let us first make some comments on the related works to two-sided radial SLE with spiral. Apart from [33] mentioned above in the context of imaginary geometry, we believe it to be also related to the mixed multifractal spectrum introduced by Binder [6]. More precisely, the points on an SLE curve are classified in terms of the asymptotic behavior of the uniformizing conformal maps in the complement of the curve, according to its Hölder exponent and its winding exponent. The Hausdorff dimension (i.e., mixed multifractal spectrum) of the set of points with a given behavior is computed in [14]. It was shown in [28] that conditioning a chordal SLE to pass through an interior point 0 gives the two-sided radial SLE (without spiral). Since the points with zero winding exponent have the maximal spectrum, this conditioning is equivalent to conditioning on the event where the 0 is a point on the SLE curve with zero winding exponent. We speculate that “conditioning” on the rare event where 0 is a point with winding rate $\mu$, the curve obtained should satisfy the commutation relation and hence has to be the two-sided radial SLE that with the spiraling rate $\mu$. In the terminology of the conformal field theory, this would correspond to inserting a curve-generating operator at 0 with complex charges and conformal weights [5].
There are also a few recent works built upon this article. A follow-up paper [19] generalizes the first class in Theorem 1.1—two-sided radial SLE with spiral—to the multi-sided case and derives its resampling and boundary perturbation properties. Another follow-up paper [15] generalizes the second class in Theorem 1.1 to the multi-chordal case and connects the corresponding partition functions with $\alpha=\alpha_{1}(\kappa)$ to random-cluster model in polygon conditional on the one-arm event. Recent papers [45, 46] of J. Zhang continue the investigation of the radial commutation by studying locally commuting $n$-radial SLE and the corresponding radial BPZ equations.Zhang constructs solutions to these radial BPZ equations using Coulomb gas formalism and relates these solutions to quantum Calogero–Sutherland system in conformal field theory.
Regarding related works on semiclassical limits of SLEs, apart from those mentioned in Section 1.5, we also mention that the resampling property in the deterministic $\kappa=0$ case is equivalent to the geodesic property studied in [29, 22] (and the two-sided radial SLE0 is a geodesic pair in their terminology). See also [8, 34, 7] for other occurrences of geodesic property. Zhang also studied the $\kappa=0$ dynamics of multiple radial SLE0 [45] in relation to the classical Calogero–Sutherland system (see also [3] in the multichordal case).
2 Partition functions of locally commuting 2-radial SLE
In this section, we give the precise definition of locally commuting $2$-radial SLE. We also define and classify the partition functions.
2.1 Definition of locally commuting 2-radial SLE
Let $D\subset\mathbb{D}$ be a simply connected domain containing 0 and let $x_{1},x_{2}$ be distinct prime ends of $D$. Let $U_{1},U_{2}$ be, respectively, closed neighborhoods of $x_{1}$ and $x_{2}$ in $D$ that do not contain 0 and such that $U_{1}\cap U_{2}=\emptyset$. We will consider probability measures $\mathbb{P}_{(D;\,x_{1},x_{2})}^{(U_{1},U_{2})}$ on pairs of unparametrized continuous curves in $U_{1}$ and $U_{2}$ starting from $x_{1}$ and $x_{2}$, and exiting $U_{1}$ and $U_{2}$ almost surely. We call that such a family of measures indexed by different choices of $(U_{1},U_{2})$ compatible if for all $U_{1}\subset U_{1}^{\prime}$ and $U_{2}\subset U_{2}^{\prime}$, we have $\mathbb{P}_{(D;\,x_{1},x_{2})}^{(U_{1},U_{2})}$ is obtained from restricting the curves under $\mathbb{P}_{(D;\,x_{1},x_{2})}^{(U_{1}^{\prime},U_{2}^{\prime})}$ to the part before first exiting the subdomains $U_{1}$ and $U_{2}$.
The locally commuting $2$-radial $\operatorname{SLE}_{\kappa}$ is a compatible family of measures $\mathbb{P}_{(D;\,x_{1},x_{2})}^{(U_{1},U_{2})}$ on pairs of continuous non-self-crossing curves $(\eta^{(1)},\eta^{(2)})$ for all $D$, $(x_{1},x_{2})$, and $(U_{1},U_{2})$ as above that satisfy additionally the (CI),(DMP),(MARG) conditions below. We say that a locally commuting $2$-radial $\operatorname{SLE}_{\kappa}$ is interchangeable if it satisfies further the condition (INT).
  1. (CI)
    Conformal invariance: If $\varphi:D\to D^{\prime}$ is a conformal map fixing 0, then the pullback measure
    $\varphi^{*}\mathbb{P}_{(D^{\prime};\,\varphi(x_{1}),\varphi(x_{2}))}^{(\varphi(U_{1}),\varphi(U_{2}))}=\mathbb{P}_{(D;\,x_{1},x_{2})}^{(U_{1},U_{2})}.$
    Therefore, it suffices to describe the measure when $(D;\,x_{1},x_{2})=(\mathbb{D};\,\mathrm{e}^{\mathfrak{i}\theta_{1}},\mathrm{e}^{\mathfrak{i}\theta_{2}})$.
  2. (DMP)
    Domain Markov property: Let $(\eta^{(1)},\eta^{(2)})\sim\mathbb{P}_{(\mathbb{D};\,\mathrm{e}^{\mathfrak{i}\theta_{1}},\mathrm{e}^{\mathfrak{i}\theta_{2}})}^{(U_{1},U_{2})}$ and we parametrize $\eta^{(1)}$ and $\eta^{(2)}$ by their own capacity in $\mathbb{D}$. Let $\mathbf{t}=(t_{1},t_{2})$, such that $t_{j}$ is a stopping time for $\eta^{(j)}$ and $\eta^{(j)}_{[0,t_{j}]}$ is contained in the interior of $U_{j}$. Let
    $\tilde{U_{j}}=U_{j}\setminus\eta^{(j)}_{[0,t_{j}]},\quad\tilde{\eta}^{(j)}=\eta^{(j)}\setminus\eta^{(j)}_{[0,t_{j}]},\quad j=1,2;\quad\tilde{D}=\mathbb{D}\setminus(\eta^{(1)}_{[0,t_{1}]}\cup\eta^{(2)}_{[0,t_{2}]})$.
    Then conditionally on $\eta^{(1)}_{[0,t_{1}]}\cup\eta^{(2)}_{[0,t_{2}]}$, we have
    $(\tilde{\eta}^{(1)},\tilde{\eta}^{(2)})\sim\mathbb{P}_{(\tilde{D};\,\eta^{(1)}_{t_{1}},\eta^{(2)}_{t_{2}})}^{(\tilde{U}_{1},\tilde{U}_{2})}.$
    See Figure 4.
    Figure 4: (CI) and (DMP) imply that a locally commuting 2-radial SLE satisfies $(\tilde{\eta}^{(1)},\tilde{\eta}^{(2)})\sim\mathbb{P}_{(\tilde{D};\,\eta^{(1)}_{t_{1}},\eta^{(2)}_{t_{2}})}^{(\tilde{U}_{1},\tilde{U}_{2})}\sim g_{\mathbf{t}}^{*}\,\mathbb{P}_{(\mathbb{D};\mathrm{e}^{\mathfrak{i}\theta^{(1)}_{\mathbf{t}}},\mathrm{e}^{\mathfrak{i}\theta^{(2)}_{\mathbf{t}}})}^{(g_{\mathbf{t}}(\tilde{U}_{1}),g_{\mathbf{t}}(\tilde{U}_{2}))}.$
  3. (MARG)
    Marginal laws: Let $(\eta^{(1)},\eta^{(2)})\sim\mathbb{P}_{(\mathbb{D};\,\mathrm{e}^{\mathfrak{i}\theta_{1}},\mathrm{e}^{\mathfrak{i}\theta_{2}})}^{(U_{1},U_{2})}$. We assume that there exist $C^{2}$ functions $b_{j}:S^{1}\times S^{1}\setminus\Delta\to\mathbb{R}$ where $S^{1}=\mathbb{R}/2\pi\mathbb{Z}$, and $\Delta$ is the diagonal $\{(\theta,\theta)\,|\,\theta\in S^{1}\}$ such that the capacity parametrized Loewner driving function $t\mapsto\xi^{(1)}_{t}$ of $\eta^{(1)}$ satisfies
    $\begin{cases}\xi^{(1)}_{0}=\theta_{1},\quad V^{(2)}_{0}=\theta_{2},\\ \mathrm{d}\xi^{(1)}_{t}=\sqrt{\kappa}\,\mathrm{d}B_{t}^{(1)}+b_{1}\left(\xi^{(1)}_{t},V^{(2)}_{t}\right)\mathrm{d}t,\\ \mathrm{d}V^{(2)}_{t}=\cot\left((V^{(2)}_{t}-\xi^{(1)}_{t})/2\right)\mathrm{d}t,\end{cases}$ (2.1)
    where $B^{(1)}$ is one-dimensional standard Brownian motion. In other words, the radial Loewner chain $g^{(1)}_{t}$ associated with $\eta^{(1)}$ maps the tip $\eta^{(1)}_{t}$ to $\exp(\mathfrak{i}\xi^{(1)}_{t})$ and $\mathrm{e}^{\mathfrak{i}\theta_{2}}$ to $\exp(\mathfrak{i}V^{(2)}_{t})$ by (1.2).
    Similarly, the capacity parametrized Loewner driving function $t\mapsto\theta^{(2)}_{t}$ of $\eta^{(2)}$ satisfies
    $\begin{cases}V^{(1)}_{0}=\theta_{1},\quad\xi^{(2)}_{0}=\theta_{2},\\ \mathrm{d}\xi^{(2)}_{t}=\sqrt{\kappa}\,\mathrm{d}B_{t}^{(2)}+b_{2}(V^{(1)}_{t},\xi^{(2)}_{t})\mathrm{d}t,\\ \mathrm{d}V^{(1)}_{t}=\cot\left((V^{(1)}_{t}-\xi^{(2)}_{t})/2\right)\mathrm{d}t.\end{cases}$ (2.2)
    In other words, the radial Loewner chain $g^{(2)}_{t}$ associated with $\eta^{(2)}$ maps the tip $\eta^{(2)}_{t}$ to $\exp(\mathfrak{i}\xi^{(2)}_{t})$ and $\mathrm{e}^{\mathfrak{i}\theta_{1}}$ to $\exp(\mathfrak{i}V^{(1)}_{t})$.
  4. (INT)
    Interchangeability condition: The two curves are unordered. In other words, let $\tau:(\eta^{(1)},\eta^{(2)})\mapsto(\eta^{(2)},\eta^{(1)})$,
    $\mathbb{P}_{(D;\,x_{1},x_{2})}^{(U_{1},U_{2})}\sim\tau^{*}\mathbb{P}_{(D;\,x_{2},x_{1})}^{(U_{2},U_{1})}.$
Remark 2.1.
Despite the heavy notation, the only purpose of introducing the neighborhoods $U_{1},U_{2}$ is to give a precise meaning of “local law” of the $2$-radial SLE near their starting points. In particular, if we have a random pair of curves in $\mathfrak{X}(\mathbb{D};x_{1},x_{2};0)$, or a random simple curve in $\mathfrak{X}(\mathbb{D};x_{1},x_{2})$, we obtain a compatible family of probability measures $\mathbb{P}_{(\mathbb{D};x_{1},x_{2})}^{(U_{1},U_{2})}$ by restricting the random curve (or the pair of random curve) to all possible pairs of neighborhoods $(U_{1},U_{2})$. Note that a priori, a compatible family of local laws does not necessarily imply they can be coupled as the restriction of a random curve or pair of curves in all $(U_{1},U_{2})$, but this will be a consequence of our classification Theorem 1.1.
To simplify notations, we also denote $\mathbb{P}_{(\mathbb{D};\,\mathrm{e}^{\mathfrak{i}\theta_{1}},\mathrm{e}^{\mathfrak{i}\theta_{2}})}^{(U_{1},U_{2})}$ by $\mathbb{P}_{(\theta_{1},\theta_{2})}$ as in Section 1.2.
These axioms allow us to characterize the local law of the $2$-radial SLE by considering the infinitesimal generators of the two-time driving function which is the subject of the next sections.
2.2 Commutation relations
In this section, we assume that $\kappa\in(0,\infty)$ and do not assume (INT). We consider a locally commuting $2$-radial $\operatorname{SLE}_{\kappa}$, which is a compatible family of measures $\mathbb{P}_{(\theta_{1},\theta_{2})}$ satisfying the conditions (CI), (DMP), and (MARG). Let $b_{1},b_{2}:S^{1}\times S^{1}\setminus\Delta\to\mathbb{R}$ be $C^{2}$ functions as in the condition (MARG).
We now derive the infinitesimal form of the radial commutation relation. Let
$\begin{split}\mathcal{L}_{1}&=\frac{\kappa}{2}\partial_{11}+b_{1}(\theta_{1},\theta_{2})\,\partial_{1}+\cot\left(\frac{\theta_{2}-\theta_{1}}{2}\right)\partial_{2}\\ \mathcal{L}_{2}&=\frac{\kappa}{2}\partial_{22}+b_{2}(\theta_{1},\theta_{2})\,\partial_{2}+\cot\left(\frac{\theta_{1}-\theta_{2}}{2}\right)\partial_{1}\end{split}$ (2.3)
be the diffusion generators associated with (2.1) and (2.2).
Proposition 2.2.
The diffusion generators (2.3) of a locally commuting 2-radial $\operatorname{SLE}_{\kappa}$ satisfies the infinitesimal commutation relation
$\left[\mathcal{L}_{1},\mathcal{L}_{2}\right]:=\mathcal{L}_{1}\mathcal{L}_{2}-\mathcal{L}_{2}\mathcal{L}_{1}=\frac{\mathcal{L}_{2}-\mathcal{L}_{1}}{\left(\sin\left((\theta_{2}-\theta_{1})/2\right)\right)^{2}}.$ (2.4)
The proof follows exactly the same steps as in [12] for chordal SLEs. The radial commutation relation was stated briefly in [12] but with an opposite sign. For the reader’s convenience, we derive it here. We introduce the following notations. They will also be used in Section 3.2.
Fix $\theta_{1}<\theta_{2}<\theta_{1}+2\pi$. We will describe the growth of a pair of continuous non-self-crossing curves $(\eta^{(1)},\eta^{(2)})$ in $\mathbb{D}$ such that $\eta^{(1)}_{0}=\mathrm{e}^{\mathfrak{i}\theta_{1}}$ and $\eta^{(2)}_{0}=\mathrm{e}^{\mathfrak{i}\theta_{2}}$. For $\mathbf{t}=(t_{1},t_{2})\in\mathbb{R}_{+}^{2}$, suppose $\eta^{(1)}_{[0,t_{1}]}$ and $\eta^{(2)}_{[0,t_{2}]}$ are disjoint. We consider the following mapping-out functions:
  • $g^{(j)}_{t_{j}}:\mathbb{D}\setminus\eta^{(j)}_{[0,t_{j}]}\to\mathbb{D}$ is conformal with $g^{(j)}_{t_{j}}(0)=0$ and $\left(g^{(j)}_{t_{j}}\right)^{\prime}(0)=\exp(t_{j})>0$, for $j=1,2$.
  • $g_{\mathbf{t}}:\mathbb{D}\setminus\left(\eta^{(1)}_{[0,t_{1}]}\cup\eta^{(2)}_{[0,t_{2}]}\right)\to\mathbb{D}$ is conformal with $g_{\mathbf{t}}(0)=0$ and $g_{\mathbf{t}}^{\prime}(0)>0$.
  • $g_{\mathbf{t},1}:\mathbb{D}\setminus g^{(1)}_{t_{1}}\left(\eta^{(2)}_{[0,t_{2}]}\right)\to\mathbb{D}$ is conformal with $g_{\mathbf{t},1}(0)=0$ and $g_{\mathbf{t},1}^{\prime}(0)>0$.
  • $g_{\mathbf{t},2}:\mathbb{D}\setminus g^{(2)}_{t_{2}}\left(\eta^{(1)}_{[0,t_{1}]}\right)\to\mathbb{D}$ is conformal with $g_{\mathbf{t},2}(0)=0$ and $g_{\mathbf{t},2}^{\prime}(0)>0$.
Using such notations, we have $g_{\mathbf{t}}=g_{\mathbf{t},j}\circ g^{(j)}_{t_{j}}$ for $j=1,2$. Let $\phi^{(j)}_{t_{j}},\phi_{\mathbf{t}},\phi_{\mathbf{t},j}$ be the covering maps of $g^{(j)}_{t_{j}},g_{\mathbf{t}},g_{\mathbf{t},j}$ respectively, i.e., the continuous function such that $g_{\cdot}(\mathrm{e}^{\mathfrak{i}\theta})=\mathrm{e}^{\mathfrak{i}\phi_{\cdot}(\theta)}$ and $\phi_{0}(\theta)=\theta$. Denote by $(\xi^{(j)}_{t_{j}},t_{j}\geq 0)$ the driving function of $\eta^{(j)}$ as a radial Loewner chain for $j=1,2$. Let
$\theta^{(1)}_{\mathbf{t}}=\phi_{\mathbf{t},1}\left(\xi^{(1)}_{t_{1}}\right),\qquad\theta^{(2)}_{\mathbf{t}}=\phi_{\mathbf{t},2}\left(\xi^{(2)}_{t_{2}}\right)$.
The pair $\mathbf{t}\mapsto(\theta^{(1)}_{\mathbf{t}},\theta^{(2)}_{\mathbf{t}})$ may be viewed as the two-time driving function of the pair $(\eta^{(1)},\eta^{(2)})$. See Figure 5.
Figure 5: We have $g_{\mathbf{t}}=g_{\mathbf{t},1}\circ g_{t_{1}}^{(1)}=g_{\mathbf{t},2}\circ g_{t_{2}}^{(2)}$.
Proof of Proposition 2.2.
The strategy of the proof consists of mapping out $\left(\eta^{(1)}_{[0,\varepsilon]},\eta^{(2)}_{[0,\varepsilon]}\right)$ in two ways, where both curves are parametrized by their intrinsic capacity seen in $\mathbb{D}$. We can either
  • first map out $\eta^{(1)}_{[0,\varepsilon]}$ by the conformal map $g_{\varepsilon}^{(1)}$, then by $g_{(\varepsilon,\varepsilon),1}$,
  • or first map out $\eta^{(2)}_{[0,\varepsilon]}$ by the conformal map $g_{\varepsilon}^{(2)}$, then by $g_{(\varepsilon,\varepsilon),2}$.
We then compare the expansions of $(\theta_{(\varepsilon,\varepsilon)}^{(1)},\theta_{(\varepsilon,\varepsilon)}^{(2)})$ in $\varepsilon$ which are expressed in terms of $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$.
More precisely, we first follow the Loewner flow $t\mapsto g_{t}^{(1)}$ until $t=\varepsilon$. Under this flow, the radial driving function is $t\mapsto\xi_{t}^{(1)}$. And $t\mapsto V_{t}^{(2)}$ satisfies
$\partial_{t}V_{t}^{(2)}=\cot\left((V_{t}^{(2)}-\xi_{t}^{(1)})/2\right).$
Hence, for any smooth test function $F\colon(\theta_{1},\theta_{2})\mapsto\mathbb{R}$, we have
$\mathbb{E}_{(\theta_{1},\theta_{2})}\left[F(\xi_{\varepsilon}^{(1)},V_{\varepsilon}^{(2)})\right]=\left(1+\varepsilon\mathcal{L}_{1}+\frac{\varepsilon^{2}\mathcal{L}_{1}^{2}}{2}\right)F(\theta_{1},\theta_{2})+o(\varepsilon^{2}).$
Now we use $g_{(\varepsilon,\varepsilon),1}$ to map out $\tilde{\eta}^{(2)}_{[0,\varepsilon]}:=g_{\varepsilon}^{(1)}(\eta_{[0,\varepsilon]}^{(2)})$. We note that the capacity of $\tilde{\eta}^{(2)}_{[0,\varepsilon]}$ is not $\varepsilon$. We compute its capacity, we note that
$\partial_{t}\left(g_{t}^{(1)}\right)^{\prime}(z)|_{t=0}=-\frac{z+\mathrm{e}^{\mathfrak{i}\theta_{1}}}{z-\mathrm{e}^{\mathfrak{i}\theta_{1}}}+z\frac{2\mathrm{e}^{\mathfrak{i}\theta_{1}}}{(z-\mathrm{e}^{\mathfrak{i}\theta_{1}})^{2}}=-\frac{z^{2}-\mathrm{e}^{2\mathfrak{i}\theta_{1}}-2z\mathrm{e}^{\mathfrak{i}\theta_{1}}}{(z-\mathrm{e}^{\mathfrak{i}\theta_{1}})^{2}}.$
Therefore, for small $\varepsilon$,
$\left(g_{\varepsilon}^{(1)}\right)^{\prime}(\mathrm{e}^{\mathfrak{i}\theta_{2}})$ $=1-\frac{\varepsilon}{2\left(\sin((\theta_{2}-\theta_{1})/2)\right)^{2}}+\mathfrak{i}\varepsilon\cot((\theta_{2}-\theta_{1})/2)+O(\varepsilon^{2}).$
So we have
$\left\lvert\left(g_{\varepsilon}^{(1)}\right)^{\prime}(\mathrm{e}^{\mathfrak{i}\theta_{2}})\right\rvert=1-\frac{\varepsilon}{2\left(\sin((\theta_{2}-\theta_{1})/2)\right)^{2}}+o(\varepsilon).$
It follows from (3.7) that the image under $g_{\varepsilon}^{(1)}$ of a set of capacity $\varepsilon$ near $\mathrm{e}^{\mathfrak{i}\theta_{2}}$ has capacity
$\varepsilon^{\prime}=\varepsilon\left(\left\lvert\left(g_{\varepsilon}^{(1)}\right)^{\prime}(\mathrm{e}^{\mathfrak{i}\theta_{2}})\right\rvert^{2}+o(\varepsilon)\right)=\varepsilon\left(1-\frac{\varepsilon}{\left(\sin((\theta_{2}-\theta_{1})/2)\right)^{2}}\right)+o(\varepsilon^{2}).$
In particular, $\tilde{\eta}^{(2)}_{[0,\varepsilon]}$ has capacity $\varepsilon^{\prime}$. Therefore, using the conditions (CI) and (DMP), we have
$\mathbb{E}_{(\theta_{1},\theta_{2})}\left[F(\theta_{(\varepsilon,\varepsilon)}^{(1)},\theta_{(\varepsilon,\varepsilon)}^{(2)})\right]$ (2.5)
$=$ $\left(1+\varepsilon\mathcal{L}_{1}+\frac{\varepsilon^{2}\mathcal{L}_{1}^{2}}{2}\right)\left(1+\varepsilon^{\prime}\mathcal{L}_{2}+\frac{(\varepsilon^{\prime})^{2}\mathcal{L}_{2}^{2}}{2}\right)F(\theta_{1},\theta_{2})+o(\varepsilon^{2})$
$=$ $\left(1+\varepsilon(\mathcal{L}_{1}+\mathcal{L}_{2})+\left(-\varepsilon\delta\mathcal{L}_{2}+\frac{\varepsilon^{2}\mathcal{L}_{1}^{2}}{2}+\frac{\varepsilon^{2}\mathcal{L}_{2}^{2}}{2}+\varepsilon^{2}\mathcal{L}_{1}\mathcal{L}_{2}\right)\right)F(\theta_{1},\theta_{2})+o(\varepsilon^{2})$
where $\delta=\varepsilon\left(\sin((\theta_{2}-\theta_{1})/2)\right)^{-2}$ so that $\varepsilon^{\prime}=\varepsilon(1-\delta)+o(\varepsilon^{2})$.
If we first map out the second curve, then the first curve, and notice that the value of $\delta$ is unchanged, we obtain that the above expectation also equals
$\left(1+\varepsilon(\mathcal{L}_{1}+\mathcal{L}_{2})+\left(-\varepsilon\delta\mathcal{L}_{1}+\frac{\varepsilon^{2}\mathcal{L}_{1}^{2}}{2}+\frac{\varepsilon^{2}\mathcal{L}_{2}^{2}}{2}+\varepsilon^{2}\mathcal{L}_{2}\mathcal{L}_{1}\right)\right)F(\theta_{1},\theta_{2})+o(\varepsilon^{2}).$
Comparing these two expansions, the coefficient of the $\varepsilon^{2}$-order terms have to coincide, we obtain the condition
$\left[\mathcal{L}_{1},\mathcal{L}_{2}\right]:=\mathcal{L}_{1}\mathcal{L}_{2}-\mathcal{L}_{2}\mathcal{L}_{1}=\frac{\delta}{\varepsilon}(\mathcal{L}_{2}-\mathcal{L}_{1})=\frac{\mathcal{L}_{2}-\mathcal{L}_{1}}{\left(\sin((\theta_{2}-\theta_{1})/2)\right)^{2}}$
as claimed. ∎
We note that the condition (CI) implies
$b_{j}(\theta_{1}+a,\theta_{2}+a)=b_{j}(\theta_{1},\theta_{2}),\qquad\forall j=1,2\text{ and }a\in\mathbb{R}.$ (2.6)
Proposition 2.3 (Radial BPZ equations).
Let $\kappa\in(0,\infty)$. Let $b_{1},b_{2}:S^{1}\times S^{1}\setminus\Delta\to\mathbb{R}$ be $C^{2}$ functions as in the condition (MARG). Then (2.4) and (2.6) imply that there exists $\mathcal{Z}:\{(\theta_{1},\theta_{2})\in\mathbb{R}^{2}\,|\,\theta_{1}<\theta_{2}<\theta_{1}+2\pi\}\to\mathbb{R}_{>0}$, called partition function, and a constant $F\in\mathbb{R}$ such that
$b_{j}=\kappa\partial_{j}\log\mathcal{Z},\quad j=1,2,$
and
$\frac{\kappa}{2}\frac{\partial_{11}\mathcal{Z}}{\mathcal{Z}}+\cot(\theta_{21}/2)\frac{\partial_{2}\mathcal{Z}}{\mathcal{Z}}-\frac{h}{2\left(\sin(\theta_{21}/2)\right)^{2}}=$ $F,$ (2.7)
$\frac{\kappa}{2}\frac{\partial_{22}\mathcal{Z}}{\mathcal{Z}}-\cot(\theta_{21}/2)\frac{\partial_{1}\mathcal{Z}}{\mathcal{Z}}-\frac{h}{2\left(\sin(\theta_{21}/2)\right)^{2}}=$ $F,$ (2.8)
where $h=(6-\kappa)/(2\kappa)$ and $\theta_{21}=\theta_{2}-\theta_{1}=-\theta_{12}$.
We note that $\mathcal{Z}$ does not always descend to a function on $(\mathbb{R}/2\pi\mathbb{Z})^{2}$. The radial BPZ equations (or BPZ–Cardy equations) were also derived in the context of conformal field theory in [23, 24].
Proof.
We use the expression (2.3) and obtain
$\left[\mathcal{L}_{1},\mathcal{L}_{2}\right]$ $=\Bigg[-\frac{\kappa}{2\left(\sin(\theta_{12}/2)\right)^{2}}\Bigg]\partial_{11}+\Bigg[\kappa\partial_{1}b_{2}-\kappa\partial_{2}b_{1}\Bigg]\partial_{12}+\Bigg[\frac{\kappa}{2\left(\sin(\theta_{12}/2)\right)^{2}}\Bigg]\partial_{22}$
$+\Bigg[\frac{\kappa-2}{4}$ $\frac{\cot(\theta_{12}/2)}{\left(\sin(\theta_{12}/2)\right)^{2}}-\frac{b_{1}}{2\left(\sin(\theta_{12}/2)\right)^{2}}-\left(\frac{\kappa}{2}\partial_{22}b_{1}+b_{2}\partial_{2}b_{1}+\cot(\theta_{12}/2)\partial_{1}b_{1}\right)\Bigg]\partial_{1}$
$+\Bigg[\frac{\kappa-2}{4}$ $\frac{\cot(\theta_{12}/2)}{\left(\sin(\theta_{12}/2)\right)^{2}}+\frac{b_{2}}{2\left(\sin(\theta_{12}/2)\right)^{2}}+\left(\frac{\kappa}{2}\partial_{11}b_{2}+b_{1}\partial_{1}b_{2}-\cot(\theta_{12}/2)\partial_{2}b_{2}\right)\Bigg]\partial_{2},$
and
$\frac{\mathcal{L}_{2}-\mathcal{L}_{1}}{\left(\sin(\theta_{12}/2)\right)^{2}}=\frac{1}{\left(\sin(\theta_{12}/2)\right)^{2}}\Bigg[\frac{\kappa}{2}(\partial_{22}-\partial_{11})+(\cot(\theta_{12}/2)-b_{1})\partial_{1}+(b_{2}+\cot(\theta_{12}/2))\partial_{2}\Bigg].$
Comparing the coefficients, then (2.4) shows
$\kappa\partial_{1}b_{2}=\kappa\partial_{2}b_{1},$ (2.9)
$\frac{\kappa}{2}\partial_{22}b_{1}+b_{2}\partial_{2}b_{1}+\cot(\theta_{12}/2)\partial_{1}b_{1}-\frac{b_{1}}{2\left(\sin(\theta_{12}/2)\right)^{2}}+\left(\frac{6-\kappa}{4}\right)\frac{\cot(\theta_{12}/2)}{\left(\sin(\theta_{12}/2)\right)^{2}}=0,$ (2.10)
$\frac{\kappa}{2}\partial_{11}b_{2}+b_{1}\partial_{1}b_{2}+\cot(\theta_{21}/2)\partial_{2}b_{2}-\frac{b_{2}}{2\left(\sin(\theta_{21}/2)\right)^{2}}+\left(\frac{6-\kappa}{4}\right)\frac{\cot(\theta_{21}/2)}{\left(\sin(\theta_{21}/2)\right)^{2}}=0.$ (2.11)
Eq. (2.9) shows that there exists a function $\mathcal{Z}:\{(\theta_{1},\theta_{2})\in\mathbb{R}^{2}\,|\,\theta_{1}<\theta_{2}<\theta_{1}+2\pi\}\to\mathbb{R}_{>0}$ such that
$b_{1}=\kappa\,\partial_{1}\log\mathcal{Z}=\kappa\frac{\partial_{1}\mathcal{Z}}{\mathcal{Z}},\quad b_{2}=\kappa\,\partial_{2}\log\mathcal{Z}=\kappa\frac{\partial_{2}\mathcal{Z}}{\mathcal{Z}}.$
Plugging it into (2.10) and (2.11), we have
$\kappa\partial_{1}\left(\frac{\kappa}{2}\frac{\partial_{22}\mathcal{Z}}{\mathcal{Z}}-\cot(\theta_{21}/2)\frac{\partial_{1}\mathcal{Z}}{\mathcal{Z}}-\frac{h}{2\left(\sin(\theta_{21}/2)\right)^{2}}\right)$ $=0,$ (2.12)
$\kappa\partial_{2}\left(\frac{\kappa}{2}\frac{\partial_{11}\mathcal{Z}}{\mathcal{Z}}+\cot(\theta_{21}/2)\frac{\partial_{2}\mathcal{Z}}{\mathcal{Z}}-\frac{h}{2\left(\sin(\theta_{21}/2)\right)^{2}}\right)$ $=0,$ (2.13)
where $h=(6-\kappa)/(2\kappa)$.
Eq. (2.12) and (2.13) imply that, there exist functions $F_{1}$ and $F_{2}$:
$\frac{\kappa}{2}\frac{\partial_{22}\mathcal{Z}}{\mathcal{Z}}-\cot(\theta_{21}/2)\frac{\partial_{1}\mathcal{Z}}{\mathcal{Z}}-\frac{h}{2\left(\sin(\theta_{21}/2)\right)^{2}}=$ $F_{2}(\theta_{2}),$
$\frac{\kappa}{2}\frac{\partial_{11}\mathcal{Z}}{\mathcal{Z}}+\cot(\theta_{21}/2)\frac{\partial_{2}\mathcal{Z}}{\mathcal{Z}}-\frac{h}{2\left(\sin(\theta_{21}/2)\right)^{2}}=$ $F_{1}(\theta_{1}).$
Using the identity
$\frac{\partial_{22}\mathcal{Z}}{\mathcal{Z}}=\partial_{2}\left(\frac{\partial_{2}\mathcal{Z}}{\mathcal{Z}}\right)+\left(\frac{\partial_{2}\mathcal{Z}}{\mathcal{Z}}\right)^{2},$
Eq. (2.6) implies that $F_{1}$ and $F_{2}$ are constants and
$\frac{\kappa}{2}\frac{\partial_{11}\mathcal{Z}}{\mathcal{Z}}+\cot(\theta_{21}/2)\frac{\partial_{2}\mathcal{Z}}{\mathcal{Z}}-\frac{h}{2\left(\sin(\theta_{21}/2)\right)^{2}}=$ $F_{1},$ (2.14)
$\frac{\kappa}{2}\frac{\partial_{22}\mathcal{Z}}{\mathcal{Z}}-\cot(\theta_{21}/2)\frac{\partial_{1}\mathcal{Z}}{\mathcal{Z}}-\frac{h}{2\left(\sin(\theta_{21}/2)\right)^{2}}=$ $F_{2}.$ (2.15)
It remains to show $F_{1}=F_{2}$. Taking the derivative of
$a\mapsto b_{j}(\theta_{1}+a,\theta_{2}+a)=\kappa(\partial_{j}\mathcal{Z}/\mathcal{Z})(\theta_{1}+a,\theta_{2}+a)$
and evaluate at $a=0$, we get from (CI) and (2.6) that
$0=\partial_{1}\left(\frac{\partial_{j}\mathcal{Z}}{\mathcal{Z}}\right)+\partial_{2}\left(\frac{\partial_{j}\mathcal{Z}}{\mathcal{Z}}\right),\qquad\forall j=1,2.$ (2.16)
From this, we have
$\frac{\partial_{11}\mathcal{Z}}{\mathcal{Z}}-\frac{\partial_{22}\mathcal{Z}}{\mathcal{Z}}$ $=\left(\frac{\partial_{1}\mathcal{Z}}{\mathcal{Z}}\right)^{2}-\left(\frac{\partial_{2}\mathcal{Z}}{\mathcal{Z}}\right)^{2}$
$\partial_{1}\left(\frac{\partial_{2}\mathcal{Z}}{\mathcal{Z}}\right)$ $=\partial_{2}\left(\frac{\partial_{1}\mathcal{Z}}{\mathcal{Z}}\right)$
$\partial_{1}\left(\frac{\partial_{1}\mathcal{Z}}{\mathcal{Z}}\right)$ $=\partial_{2}\left(\frac{\partial_{2}\mathcal{Z}}{\mathcal{Z}}\right).$
If we take the difference (2.14) - (2.15), we get
$\frac{\kappa}{2}\left(\left(\frac{\partial_{1}\mathcal{Z}}{\mathcal{Z}}\right)^{2}-\left(\frac{\partial_{2}\mathcal{Z}}{\mathcal{Z}}\right)^{2}\right)+\cot(\theta_{21}/2)\left(\frac{\partial_{2}\mathcal{Z}}{\mathcal{Z}}+\frac{\partial_{1}\mathcal{Z}}{\mathcal{Z}}\right)=F_{1}-F_{2}.$ (2.17)
Note that (2.16) also implies
$0=\partial_{1}\left(\frac{\partial_{1}\mathcal{Z}}{\mathcal{Z}}\right)+\partial_{2}\left(\frac{\partial_{1}\mathcal{Z}}{\mathcal{Z}}\right)=\partial_{1}\left(\frac{\partial_{1}\mathcal{Z}}{\mathcal{Z}}\right)+\partial_{1}\left(\frac{\partial_{2}\mathcal{Z}}{\mathcal{Z}}\right)=\partial_{1}\left(\frac{\partial_{1}\mathcal{Z}}{\mathcal{Z}}+\frac{\partial_{2}\mathcal{Z}}{\mathcal{Z}}\right)$
and
$0=\partial_{2}\left(\frac{\partial_{1}\mathcal{Z}}{\mathcal{Z}}+\frac{\partial_{2}\mathcal{Z}}{\mathcal{Z}}\right).$
Hence there is $\mu\in\mathbb{R}$ such that
$\frac{2\mu}{\kappa}\equiv\frac{\partial_{1}\mathcal{Z}}{\mathcal{Z}}+\frac{\partial_{2}\mathcal{Z}}{\mathcal{Z}}.$ (2.18)
Plugging into (2.17) we get
$\left(\frac{\partial_{1}\mathcal{Z}}{\mathcal{Z}}-\frac{\partial_{2}\mathcal{Z}}{\mathcal{Z}}\right)\mu+\cot(\theta_{21}/2)\frac{2\mu}{\kappa}=F_{1}-F_{2}.$ (2.19)
Combining (2.18) and (2.19), there are two cases.
Case 1: We have either $\mu=0$, then $F_{1}=F_{2}$ as desired.
Case 2: If $\mu\neq 0$, then
$\begin{cases}\dfrac{\partial_{1}\mathcal{Z}}{\mathcal{Z}}=-\frac{1}{\kappa}\cot(\theta_{21}/2)+\frac{\mu}{\kappa}+\frac{F_{1}-F_{2}}{2\mu},\\ \dfrac{\partial_{2}\mathcal{Z}}{\mathcal{Z}}=\frac{1}{\kappa}\cot(\theta_{21}/2)+\frac{\mu}{\kappa}-\frac{F_{1}-F_{2}}{2\mu}.\end{cases}$ (2.20)
We solve (2.20) and obtain that, for some constant $C\in\mathbb{R}$,
$\mathcal{Z}(\theta_{1},\theta_{2})=C\left(\sin(\theta_{21}/2)\right)^{2/\kappa}\exp\left(\frac{\mu_{1}}{\kappa}\theta_{1}+\frac{\mu_{2}}{\kappa}\theta_{2}\right),$ (2.21)
where
$\mu_{1}=\mu+\frac{\kappa(F_{1}-F_{2})}{2\mu},\quad\mu_{2}=\mu-\frac{\kappa(F_{1}-F_{2})}{2\mu}.$
Plugging into (2.14)-(2.15), we obtain $\mu_{1}=\mu_{2}$ which implies $F_{1}=F_{2}$ as desired. Moreover, in this case we have
$F_{1}=F_{2}=\frac{-3+\mu^{2}}{2\kappa},\qquad\mathcal{Z}\propto\mathcal{G}_{\mu},\qquad\text{ when }\mu\neq 0.$
This completes the proof. ∎
2.3 Solutions to commutation relations with interchangeability
We now classify all possible partition functions $\mathcal{Z}$ with the additional condition (INT), which is equivalent to
$b_{1}(\theta_{1},\theta_{2})=b_{2}(\theta_{2},\theta_{1}).$ (2.22)
Theorem 2.4.
Let $\kappa\in(0,8)$ and $\mathcal{Z}$ be a partition function from Proposition 2.3 and assume that (2.22) holds. Then, up to a multiplicative constant, $\mathcal{Z}$ is one of the following functions:
  1. 1.
    $\mathcal{Z}=\mathcal{G}_{\mu}$ for some $\mu\in\mathbb{R}$, where $\mathcal{G}_{\mu}$ is defined in (1.3).
  2. 2.
    $\mathcal{Z}=\mathcal{Z}_{\alpha}$ for some $\alpha<1-\kappa/8$, where $\mathcal{Z}_{\alpha}$ is defined in (1.4).
Throughout this section, we assume that $\mathcal{Z}$ is a partition function from Proposition 2.3 and assume that (2.22) holds.
Lemma 2.5.
There exists $\lambda>0$ such that for all $\theta_{1}<\theta_{2}<\theta_{1}+2\pi$,
$\lambda\mathcal{Z}(\theta_{1},\theta_{2})=\mathcal{Z}(\theta_{2},\theta_{1}+2\pi).$ (2.23)
Proof.
Assumption (2.22) implies that
$\kappa\partial_{1}\log\left(\frac{\mathcal{Z}(\theta_{2},\theta_{1}+2\pi)}{\mathcal{Z}(\theta_{1},\theta_{2})}\right)=b_{2}(\theta_{2},\theta_{1}+2\pi)-b_{1}(\theta_{1},\theta_{2})=b_{2}(\theta_{2},\theta_{1})-b_{1}(\theta_{1},\theta_{2})=0$
and similarly,
$\partial_{2}\log\left(\frac{\mathcal{Z}(\theta_{2},\theta_{1}+2\pi)}{\mathcal{Z}(\theta_{1},\theta_{2})}\right)=0.$
Therefore, $\mathcal{Z}(\theta_{2},\theta_{1}+2\pi)/\mathcal{Z}(\theta_{1},\theta_{2})$ is a constant $\lambda$ which also equals $\mathcal{Z}(\theta_{1}+2\pi,\theta_{2}+2\pi)/Z(\theta_{2},\theta_{1}+2\pi)$. ∎
Proof of Theorem 2.4.
From the proof of Proposition 2.3, combining (2.18) and (2.19), we have two cases: we have either $\mu\neq 0$, then
$\mathcal{Z}\propto\mathcal{G}_{\mu}$
which satisfies the interchangeability condition (2.22).
Or we have $\mu=0$, then
$\partial_{1}\mathcal{Z}=-\partial_{2}\mathcal{Z}.$ (2.24)
Eq. (2.24) implies that $\partial_{a}\mathcal{Z}(\theta_{1}+a,\theta_{2}+a)=0$, in other words, $\mathcal{Z}(\theta_{1},\theta_{2})$ only depends on the difference $\theta=\theta_{2}-\theta_{1}$. Thus we write with a slight abuse of notation
$\mathcal{Z}(\theta_{1},\theta_{2})=\mathcal{Z}(\theta),\quad\theta\in(0,2\pi).$
We want to solve the equations (2.7) and (2.8), they are simplified to the same equation:
$\frac{\kappa}{2}\frac{\mathcal{Z}^{\prime\prime}}{\mathcal{Z}}+\cot(\theta/2)\frac{\mathcal{Z}^{\prime}}{\mathcal{Z}}-\frac{h}{2\left(\sin(\theta/2)\right)^{2}}=F.$ (2.25)
Noticing that
$\lambda^{2}=\frac{\mathcal{Z}(\theta_{1}+2\pi,\theta_{2}+2\pi)}{\mathcal{Z}(\theta_{1},\theta_{2})}=1,$
we obtain $\lambda=1$ and
$\mathcal{Z}(\theta)=\mathcal{Z}(2\pi-\theta).$ (2.26)
To solve (2.25) under the assumption (2.26), we may write
$\theta\in(0,2\pi),\quad u=(\sin(\theta/4))^{2}\in(0,1),\quad\mathcal{Z}(\theta)=C(\sin(\theta/2))^{-2h}\phi(u),$ (2.27)
satisfying for $u\in(0,1)$,
$\begin{cases}u(1-u)\phi^{\prime\prime}+\frac{3\kappa-8}{2\kappa}(1-2u)\phi^{\prime}+\frac{8}{\kappa}\left(\frac{(6-\kappa)(\kappa-2)}{8\kappa}-F\right)\phi=0,\\ \phi(1/2)=1,\quad\phi^{\prime}(1/2)=0,\end{cases}$ (2.28)
where $C$ is a positive constant. Eq. (2.28) has a unique solution $\phi$ in $C^{2}(0,1)$ due to Lemma A.1. More precisely:
  • When
    $\alpha:=\frac{(6-\kappa)(\kappa-2)}{8\kappa}-F<1-\frac{\kappa}{8},\quad\text{ i.e., }\quad F>-\frac{3}{2\kappa},$
    the unique solution $\phi(u)$ is positive for all $u\in(0,1)$.
  • When $\alpha=1-\kappa/8$,
    $\phi(u)=(4u(1-u))^{4/\kappa-1/2}=\left(\sin(\theta/2)\right)^{8/\kappa-1}$
    by (A.1). Hence,
    $\mathcal{Z}(\theta)=C\left(\sin(\theta/2)\right)^{2/\kappa}\propto\mathcal{G}_{0}.$
  • When $\alpha>1-\kappa/8$, Lemma A.1 shows that the unique solution is not always positive. Therefore it does not give any partition function in Proposition 2.3. See Figure 6 for the case when $\kappa=4$.
This completes the proof. ∎
Remark 2.6.
When $\kappa=4$, the partition function $\mathcal{Z}_{\alpha}$ in (1.4) has an explicit expression. We denote $\theta=\theta_{2}-\theta_{1}$.
  • If $\alpha\in[0,1/2)$, we have
    $\mathcal{Z}_{\alpha}(\theta_{1},\theta_{2})=\left(\sin(\theta/2)\right)^{-1/2}\cos\left(\sqrt{\frac{\alpha}{2}}(\theta-\pi)\right).$
  • If $\alpha<0$, we have
    $\mathcal{Z}_{\alpha}(\theta_{1},\theta_{2})=\left(\sin(\theta/2)\right)^{-1/2}\cosh\left(\sqrt{\frac{|\alpha|}{2}}(\theta-\pi)\right).$
See Figure 6.
Figure 6: Plot of $\theta\mapsto\mathcal{Z}_{\alpha}(\theta)$ with $\kappa=4$ and $\theta=\theta_{2}-\theta_{1}$ for different $\alpha$’s. When $\alpha>1/2$, it is not always positive.
Remark 2.7.
When $\kappa\in(0,8)$ and $\alpha=\alpha_{1}(\kappa)$ which is the one-arm exponent for conformal loop ensemble [40]:
$\alpha_{1}(\kappa)=\frac{(3\kappa-8)(8-\kappa)}{32\kappa},$
the partition function $\mathcal{Z}_{\alpha_{1}(\kappa)}$ has an explicit expression: we denote $\theta=\theta_{2}-\theta_{1}$,
$\mathcal{Z}_{\alpha_{1}(\kappa)}(\theta_{1},\theta_{2})=2^{4/\kappa-3/2}(\sin(\theta/2))^{1-6/\kappa}\left((\sin(\theta/4))^{8/\kappa-1}+(\cos(\theta/4))^{8/\kappa-1}\right).$ (2.29)
For $\kappa\in(4,8)$, this is the conjectured partition function for the scaling limit of interfaces in critical random-cluster models with Dobrushin boundary condition conditional on the one-arm event, see [15]. This conjecture holds for percolation (with $\kappa=6$) and FK-Ising model (with $\kappa=16/3$).
Remark 2.8.
Let us give a comment on the $C^{2}$ requirement on $b_{1},b_{2}$ in the assumption (MARG). We assume such $C^{2}$ requirement as part of the axioms at the beginning. But in fact, this is not necessary. Suppose we relax the requirement and only assume $b_{1},b_{2}$ in (MARG) are continuous, we are still able to derive (2.14) and (2.15) as weak solutions. The operators in the left-hand side of radial BPZ equations (2.14) and (2.15) are hypoelliptic, due to a general characterization by Hörmander [18], see also [13, Lemma 5] and [35, Sect. 2.3.3]. Therefore, the weak solutions are strong solutions which are in fact $C^{\infty}$, and consequently $b_{1},b_{2}$ are $C^{\infty}$. Note that such analysis does not work for $\kappa=0$ since the corresponding BPZ equation (4.1) is nonlinear.
3 Identification of locally commuting 2-radial SLEs
The goal of this section is to identify all locally commuting 2-radial SLEs, namely, those whose partition functions are classified in Theorem 2.4: We discuss the 2-sided radial SLE with spiral in Section 3.2 and chordal SLE weighted by conformal radius in Section 3.4.
3.1 Radial SLE
For $\theta\in[0,2\pi)$, suppose $\eta:[0,T]\to\overline{\mathbb{D}}$ is a continuous non-self-crossing curve such that $\eta_{0}=\mathrm{e}^{\mathfrak{i}\theta}$ and $\eta_{(0,T)}\subset\mathbb{D}\setminus\{0\}$. We parameterize the curve by the capacity and denote by $g_{t}$ the corresponding radial Loewner chain as in (1.1). Denote by $\phi_{t}$ the covering conformal map of $g_{t}$, i.e. $g_{t}(\exp(\mathfrak{i}w))=\exp(\mathfrak{i}\phi_{t}(w))$ with $\phi_{0}(w)=w$ for $w\in\mathbb{H}$. Then the radial Loewner equation (1.1) is equivalent to
$\partial_{t}\phi_{t}(w)=\cot\left((\phi_{t}(w)-\xi_{t})/2\right),\quad\phi_{0}(w)=w.$
Radial $\operatorname{SLE}_{\kappa}$ is the radial Loewner chain with $\xi_{t}=\sqrt{\kappa}B_{t}$ where $B$ is one-dimensional Brownian motion. We also call it radial $\operatorname{SLE}_{\kappa}$ in $(\mathbb{D};\mathrm{e}^{\mathfrak{i}\theta};0)$. In the following lemma, we will describe the boundary perturbation property of radial SLE. We fix the parameters:
$\kappa>0,\quad h=\frac{6-\kappa}{2\kappa},\quad\tilde{h}=\frac{(6-\kappa)(\kappa-2)}{8\kappa},\quad c=\frac{(6-\kappa)(3\kappa-8)}{2\kappa}.$ (3.1)
Lemma 3.1 (See [20, Prop. 5] or [17, Prop. 2.2]).
Fix $\kappa\in(0,8)$ and $\theta\in[0,2\pi)$. Suppose $K$ is a compact subset of $\overline{\mathbb{D}}$ such that $\mathbb{D}\setminus K$ is simply connected and contains the origin and that $K$ has a positive distance from $\mathrm{e}^{\mathfrak{i}\theta}$. Suppose $\eta$ is radial $\operatorname{SLE}_{\kappa}$ in $(\mathbb{D};\mathrm{e}^{\mathfrak{i}\theta};0)$ and define $\tau=\inf\{t:\eta_{t}\in K\}$. For $t<\tau$, denote by $g_{t,K}$ the unique conformal map $\mathbb{D}\setminus g_{t}(K)\to\mathbb{D}$ such that $g_{t,K}(0)=0$ and $g_{t,K}^{\prime}(0)>0$. We denote by $\phi_{t,K}$ the covering map of $g_{t,K}$. Then the following process is a local martingale:
$M_{t}=1{\{t<\tau\}}\phi_{t,K}^{\prime}(\xi_{t})^{h}g_{t,K}^{\prime}(0)^{\tilde{h}}\exp\left(\frac{c}{2}m_{t}\right),$
where $m_{t}$ is defined through
$\mathrm{d}m_{t}=-\frac{1}{3}\mathcal{S}\phi_{t,K}(\xi_{t})\mathrm{d}t+\frac{1}{6}\left(1-\phi_{t,K}^{\prime}(\xi_{t})^{2}\right)\mathrm{d}t,$
and $\mathcal{S}\phi=\frac{\phi^{\prime\prime\prime}}{\phi^{\prime}}-\frac{3}{2}\left(\frac{\phi^{\prime\prime}}{\phi^{\prime}}\right)^{2}$ denotes the Schwarzian derivative of $\phi$.
Moreover, when $\kappa\leq 4$, the process $M_{t}$ is a uniformly integrable martingale. The law of radial $\operatorname{SLE}_{\kappa}$ in $(\mathbb{D}\setminus K;\mathrm{e}^{\mathfrak{i}\theta};0)$ is the same as radial $\operatorname{SLE}_{\kappa}$ in $(\mathbb{D};\mathrm{e}^{\mathfrak{i}\theta};0)$ weighted by $M_{t}$.
Remark 3.2.
It is explained in the proof of [20, Prop. 5] that the term $m_{t}=m_{\mathbb{D}}(\eta_{[0,t]},K)$ is the same as the Brownian loop measure of loops that intersect both $\eta_{[0,t]}$ and $K$ when $\eta_{[0,t]}\cap K=\emptyset$.
Fix $\theta_{1},\theta_{2}$ such that $\theta_{1}<\theta_{2}<\theta_{1}+2\pi$. Let $\kappa\in[0,\infty)$, $\rho\in\mathbb{R}$ and $\mu\in\mathbb{R}$. A radial $\operatorname{SLE}_{\kappa}^{\mu}(\rho)$ in $\mathbb{D}$ starting from $\mathrm{e}^{\mathfrak{i}\theta_{1}}$ with force point $\mathrm{e}^{\mathfrak{i}\theta_{2}}$ and spiraling rate $\mu$ is the radial Loewner chain with driving function $\xi_{t}$ that solves the following SDE:
$\begin{cases}\xi_{0}=\theta_{1},V_{0}=\theta_{2},\\ \mathrm{d}\xi_{t}=\sqrt{\kappa}\mathrm{d}B_{t}+\frac{\rho}{2}\cot\left((\xi_{t}-V_{t})/2\right)\mathrm{d}t+\mu\mathrm{d}t,\\ \mathrm{d}V_{t}=\cot\left((V_{t}-\xi_{t})/2\right)\mathrm{d}t.\end{cases}$ (3.2)
The solution to SDE (3.2) exists for all time when $\kappa\in(0,8)$ and $\rho>-2$ and it is generated by a continuous curve from $\mathrm{e}^{\mathfrak{i}\theta_{1}}$ to the origin.
Lemma 3.3.
For $\kappa\in(0,8)$, $\rho>-2$, $\mu\in\mathbb{R}$, radial $\operatorname{SLE}_{\kappa}^{\mu}(\rho)$ in $\mathbb{D}$ is almost surely generated by a continuous curve $\eta$ and $\lim_{t\to\infty}\eta_{t}=0$.
Proof.
This is proved in [33, Prop. 3.30] and [33, Sect. 4]. See also [25]. ∎
Remark 3.4.
The expression of $\mathcal{G}_{\mu}$ in (1.3) is such that the Loewner driving function of radial $\operatorname{SLE}_{\kappa}^{\mu}(2)$ can be rewritten as
$\begin{cases}\xi_{0}=\theta_{1},V_{0}=\theta_{2},\\ \mathrm{d}\xi_{t}=\sqrt{\kappa}\mathrm{d}B_{t}+\kappa\,\partial_{1}\log\mathcal{G}_{\mu}(\xi_{t},V_{t})\mathrm{d}t,\\ \mathrm{d}V_{t}=\cot\left((V_{t}-\xi_{t})/2\right)\mathrm{d}t.\end{cases}$
[…]
3.2 Two-sided radial SLE with spiral
In this section, we define two-sided radial $\operatorname{SLE}_{\kappa}$ with spiral when $\kappa\in(0,8)$ by reweighting two independent radial SLEκ. The two-sided radial SLE analyzed in [25, 16, 28, 17] is a special case where the spiraling rate $\mu=0$.
We use the same notations as in Figure 5.
Lemma 3.5.
Fix $\kappa\in(0,8)$ and $\theta_{1}<\theta_{2}<\theta_{1}+2\pi$. We fix the parameters $h,\tilde{h},c$ as in (3.1). For $\mu\in\mathbb{R}$, we define $\mathcal{G}_{\mu}$ as in (1.3). Let $\mathbb{P}$ denote the probability measure under which $(\eta^{(1)},\eta^{(2)})$ are two independent radial $\operatorname{SLE}_{\kappa}$ in $\mathbb{D}$ starting from $\mathrm{e}^{\mathfrak{i}\theta_{1}}$ and $\mathrm{e}^{\mathfrak{i}\theta_{2}}$ respectively. We define
$M_{\mathbf{t}}(\mathcal{G}_{\mu})=$ $1\left\{\eta^{(1)}_{[0,t_{1}]}\cap\eta^{(2)}_{[0,t_{2}]}=\emptyset\right\}g_{\mathbf{t}}^{\prime}(0)^{\frac{3-\mu^{2}}{2\kappa}-\tilde{h}}\times\prod_{j=1}^{2}\phi_{\mathbf{t},j}^{\prime}\left(\xi_{t_{j}}^{(j)}\right)^{h}g_{\mathbf{t},j}^{\prime}(0)^{\tilde{h}}$
$\times\mathcal{G}_{\mu}\left(\theta_{\mathbf{t}}^{(1)},\theta_{\mathbf{t}}^{(2)}\right)\exp\left(\frac{c}{2}m_{\mathbf{t}}\right),$ (3.3)
where $m_{\mathbf{t}}$ is defined through
$\mathrm{d}m_{\mathbf{t}}=\sum_{j=1}^{2}\left(-\frac{1}{3}\mathcal{S}\phi_{\mathbf{t},j}\left(\xi^{(j)}_{t_{j}}\right)+\frac{1}{6}\left(1-\phi_{\mathbf{t},j}^{\prime}\left(\xi_{t_{j}}^{(j)}\right)^{2}\right)\right)\mathrm{d}t_{j}.$ (3.4)
Then $M_{\mathbf{t}}(\mathcal{G}_{\mu})$ is a two-time-parameter local martingale with respect to $\mathbb{P}$.
Lemma 3.1 and Remark 3.2 show that $m_{\mathbf{t}}=m_{\mathbb{D}}\left(\eta_{[0,t_{1}]}^{(1)},\eta_{[0,t_{2}]}^{(2)}\right)$ is the Brownian loop measure of loops in $\mathbb{D}$ intersecting both $\eta_{[0,t_{1}]}^{(1)}$ and $\eta_{[0,t_{2}]}^{(2)}$ when $\eta_{[0,t_{1}]}^{(1)}\cap\eta_{[0,t_{2}]}^{(2)}=\emptyset$.
We note that $\mathcal{G}_{\mu}$ satisfies the “radial BPZ equations”
$\frac{\kappa}{2}\frac{\partial_{11}\mathcal{G}_{\mu}}{\mathcal{G}_{\mu}}+\cot(\theta_{21}/2)\frac{\partial_{2}\mathcal{G}_{\mu}}{\mathcal{G}_{\mu}}-\frac{h}{2\left(\sin(\theta_{21}/2)\right)^{2}}=$ $\frac{\mu^{2}-3}{2\kappa};$ (3.5)
$\frac{\kappa}{2}\frac{\partial_{22}\mathcal{G}_{\mu}}{\mathcal{G}_{\mu}}-\cot(\theta_{21}/2)\frac{\partial_{1}\mathcal{G}_{\mu}}{\mathcal{G}_{\mu}}-\frac{h}{2\left(\sin(\theta_{21}/2)\right)^{2}}=$ $\frac{\mu^{2}-3}{2\kappa}.$ (3.6)
Relations (3.4), (3.5) and (3.6) play an essential role in the proof of Lemma 3.5.
Proof of Lemma 3.5.
Let us first compute the variations of terms appearing in (3.5). From standard calculations (see e.g. [17, Lem. 3.2]), we have the following variational formula of the capacity parameterizations:
$\begin{split}\frac{\mathrm{d}g^{\prime}_{\mathbf{t},1}(0)}{g^{\prime}_{\mathbf{t},1}(0)}=&\left(\phi^{\prime}_{\mathbf{t},1}\left(\xi^{(1)}_{t_{1}}\right)^{2}-1\right)\mathrm{d}t_{1}+\phi^{\prime}_{\mathbf{t},2}\left(\xi^{(2)}_{t_{2}}\right)^{2}\mathrm{d}t_{2};\\ \frac{\mathrm{d}g^{\prime}_{\mathbf{t},2}(0)}{g^{\prime}_{\mathbf{t},2}(0)}=&\,\phi^{\prime}_{\mathbf{t},1}\left(\xi^{(1)}_{t_{1}}\right)^{2}\mathrm{d}t_{1}+\left(\phi^{\prime}_{\mathbf{t},2}\left(\xi^{(2)}_{t_{2}}\right)^{2}-1\right)\mathrm{d}t_{2};\\ \frac{\mathrm{d}g_{\mathbf{t}}^{\prime}(0)}{g_{\mathbf{t}}^{\prime}(0)}=&\sum_{j=1}^{2}\phi_{\mathbf{t},j}^{\prime}\left(\xi_{t_{j}}^{(j)}\right)^{2}\mathrm{d}t_{j}.\end{split}$ (3.7)
From the assumption that $\eta^{(1)}$ and $\eta^{(2)}$ are two independent radial $\operatorname{SLE}_{\kappa}$ under $\mathbb{P}$, we have that $\xi^{(1)}=\sqrt{\kappa}B^{(1)}+\theta_{1}$ and $\xi^{(2)}=\sqrt{\kappa}B^{(2)}+\theta_{2}$ where $B^{(1)}$ and $B^{(2)}$ are two independent Brownian motions. From this, Itô’s calculus gives:
$\mathrm{d}\theta_{\mathbf{t}}^{(1)}=\mathrm{d}\phi_{\mathbf{t},1}\left(\xi^{(1)}_{t_{1}}\right)$= $\,\phi^{\prime}_{\mathbf{t},1}\left(\xi^{(1)}_{t_{1}}\right)\mathrm{d}\xi_{t_{1}}^{(1)}-\kappa h\phi^{\prime\prime}_{\mathbf{t},1}\left(\xi^{(1)}_{t_{1}}\right)\mathrm{d}t_{1}$
$+\cot\left((\theta_{\mathbf{t}}^{(1)}-\theta_{\mathbf{t}}^{(2)})/2\right)\phi^{\prime}_{\mathbf{t},2}\left(\xi^{(2)}_{t_{2}}\right)^{2}\mathrm{d}t_{2},$ (3.8)
where we used the expansions
$\partial_{t_{1}}\phi_{\mathbf{t},1}(w)$ $=\left(\phi_{\mathbf{t},1}^{(1)}\right)^{\prime}\left(\xi_{t_{1}}^{(1)}\right)^{2}\cot\left((\phi_{\mathbf{t},1}^{(1)}(w)-\theta_{\mathbf{t}}^{(1)})/2\right)-\left(\phi_{\mathbf{t},1}^{(1)}\right)^{\prime}(w)\cot\left((w-\xi_{t_{1}}^{(1)})/2\right)$
$=-3\left(\phi_{\mathbf{t},1}^{(1)}\right)^{\prime\prime}\left(\xi_{t_{1}}^{(1)}\right)+O(w-\xi_{t_{1}}^{(1)}),\quad\text{as }w\to\xi_{t_{1}}^{(1)},$ (3.9)
$\partial_{t_{2}}\phi_{\mathbf{t},1}(w)$ $=\cot\left((\phi_{\mathbf{t},1}^{(1)}(w)-\theta_{\mathbf{t}}^{(2)})/2\right)\phi_{\mathbf{t},2}^{\prime}\left(\xi_{t_{2}}^{(2)}\right)^{2}.$ (3.10)
Similarly, we have
$\mathrm{d}\theta_{\mathbf{t}}^{(2)}=\mathrm{d}\phi_{\mathbf{t},2}\left(\xi^{(2)}_{t_{2}}\right)$= $\phi^{\prime}_{\mathbf{t},2}\left(\xi^{(2)}_{t_{2}}\right)\mathrm{d}\xi_{t_{2}}^{(2)}-\kappa h\phi^{\prime\prime}_{\mathbf{t},2}\left(\xi^{(2)}_{t_{2}}\right)\mathrm{d}t_{2}$
$+\cot\left((\theta_{\mathbf{t}}^{(2)}-\theta_{\mathbf{t}}^{(1)})/2\right)\phi^{\prime}_{\mathbf{t},1}\left(\xi^{(1)}_{t_{1}}\right)^{2}\mathrm{d}t_{1}.$ (3.11)
Taking derivatives with respect to $w$ in (3.9) and (3.10) and let $w\to\xi^{(1)}_{t_{1}}$ we obtain
$\frac{(\partial_{t_{1}}\phi^{\prime}_{\mathbf{t},1})\left(\xi^{(1)}_{t_{1}}\right)}{\phi^{\prime}_{\mathbf{t},1}\left(\xi^{(1)}_{t_{1}}\right)}$= $\frac{1}{2}\frac{\phi^{\prime\prime}_{\mathbf{t},1}\left(\xi^{(1)}_{t_{1}}\right)^{2}}{\phi^{\prime}_{\mathbf{t},1}\left(\xi^{(1)}_{t_{1}}\right)^{2}}-\frac{4}{3}\frac{\phi^{\prime\prime\prime}_{\mathbf{t},1}\left(\xi^{(1)}_{t_{1}}\right)}{\phi^{\prime}_{\mathbf{t},1}\left(\xi^{(1)}_{t_{1}}\right)}-\frac{1}{6}\left(\phi^{\prime}_{\mathbf{t},1}\left(\xi^{(1)}_{t_{1}}\right)^{2}-1\right),$
$\frac{(\partial_{t_{2}}\phi^{\prime}_{\mathbf{t},1})\left(\xi^{(1)}_{t_{1}}\right)}{\phi^{\prime}_{\mathbf{t},1}\left(\xi^{(1)}_{t_{1}}\right)}$= $-\frac{1}{2}\csc^{2}\left((\theta_{\mathbf{t}}^{(1)}-\theta_{\mathbf{t}}^{(2)})/2\right)\phi^{\prime}_{\mathbf{t},2}\left(\xi^{(2)}_{t_{2}}\right)^{2}.$
Itô’s formula gives
$\frac{\mathrm{d}\phi^{\prime}_{\mathbf{t},1}\left(\xi^{(1)}_{t_{1}}\right)}{\phi^{\prime}_{\mathbf{t},1}\left(\xi^{(1)}_{t_{1}}\right)}$= $\frac{\phi^{\prime\prime}_{\mathbf{t},1}\left(\xi^{(1)}_{t_{1}}\right)}{\phi^{\prime}_{\mathbf{t},1}\left(\xi^{(1)}_{t_{1}}\right)}\mathrm{d}\xi_{t_{1}}^{(1)}-\frac{1}{2}\csc^{2}\left((\theta_{\mathbf{t}}^{(1)}-\theta_{\mathbf{t}}^{(2)})/2\right)\phi^{\prime}_{\mathbf{t},2}\left(\xi^{(2)}_{t_{2}}\right)^{2}\mathrm{d}t_{2}$
$+\left(\frac{\kappa}{2}\frac{\phi^{\prime\prime\prime}_{\mathbf{t},1}\left(\xi^{(1)}_{t_{1}}\right)}{\phi^{\prime}_{\mathbf{t},1}\left(\xi^{(1)}_{t_{1}}\right)}+\frac{1}{2}\frac{\phi^{\prime\prime}_{\mathbf{t},1}\left(\xi^{(1)}_{t_{1}}\right)^{2}}{\phi^{\prime}_{\mathbf{t},1}\left(\xi^{(1)}_{t_{1}}\right)^{2}}-\frac{4}{3}\frac{\phi^{\prime\prime\prime}_{\mathbf{t},1}\left(\xi^{(1)}_{t_{1}}\right)}{\phi^{\prime}_{\mathbf{t},1}\left(\xi^{(1)}_{t_{1}}\right)}-\frac{1}{6}\left(\phi^{\prime}_{\mathbf{t},1}\left(\xi^{(1)}_{t_{1}}\right)^{2}-1\right)\right)\mathrm{d}t_{1},$
$\frac{\mathrm{d}\phi^{\prime}_{\mathbf{t},2}\left(\xi^{(2)}_{t_{2}}\right)}{\phi^{\prime}_{\mathbf{t},2}\left(\xi^{(2)}_{t_{2}}\right)}$= $\frac{\phi^{\prime\prime}_{\mathbf{t},2}\left(\xi^{(2)}_{t_{2}}\right)}{\phi^{\prime}_{\mathbf{t},2}\left(\xi^{(2)}_{t_{2}}\right)}\mathrm{d}\xi_{t_{2}}^{(2)}-\frac{1}{2}\csc^{2}\left((\theta_{\mathbf{t}}^{(2)}-\theta_{\mathbf{t}}^{(1)})/2\right)\phi^{\prime}_{\mathbf{t},1}\left(\xi^{(1)}_{t_{1}}\right)^{2}\mathrm{d}t_{1}$
$+\left(\frac{\kappa}{2}\frac{\phi^{\prime\prime\prime}_{\mathbf{t},2}\left(\xi^{(2)}_{t_{2}}\right)}{\phi^{\prime}_{\mathbf{t},2}\left(\xi^{(2)}_{t_{2}}\right)}+\frac{1}{2}\frac{\phi^{\prime\prime}_{\mathbf{t},2}\left(\xi^{(2)}_{t_{2}}\right)^{2}}{\phi^{\prime}_{\mathbf{t},2}\left(\xi^{(2)}_{t_{2}}\right)^{2}}-\frac{4}{3}\frac{\phi^{\prime\prime\prime}_{\mathbf{t},2}\left(\xi^{(2)}_{t_{2}}\right)}{\phi^{\prime}_{\mathbf{t},2}\left(\xi^{(2)}_{t_{2}}\right)}-\frac{1}{6}\left(\phi^{\prime}_{\mathbf{t},2}\left(\xi^{(2)}_{t_{2}}\right)^{2}-1\right)\right)\mathrm{d}t_{2}.$
Now we are ready to prove that $M_{\mathbf{t}}(\mathcal{G}_{\mu})$ is a two-time-parameter local martingale. Combining with (3.4), (3.5) and (3.6), we have
$\frac{\mathrm{d}M_{\mathbf{t}}(\mathcal{G}_{\mu})}{M_{\mathbf{t}}(\mathcal{G}_{\mu})}=$ $\left(\frac{3-\mu^{2}}{2\kappa}-\tilde{h}\right)\frac{\mathrm{d}g_{\mathbf{t}}^{\prime}(0)}{g_{\mathbf{t}}^{\prime}(0)}+\tilde{h}\sum_{j=1}^{2}\frac{\mathrm{d}g^{\prime}_{\mathbf{t},j}(0)}{g^{\prime}_{\mathbf{t},j}(0)}+\frac{c}{2}\mathrm{d}m_{\mathbf{t}}$
$+\sum_{j=1}^{2}\left(\frac{\partial_{j}\mathcal{G}_{\mu}}{\mathcal{G}_{\mu}}\mathrm{d}\theta_{\mathbf{t}}^{(j)}+\frac{\kappa}{2}\frac{\partial_{jj}\mathcal{G}_{\mu}}{\mathcal{G}_{\mu}}\phi^{\prime}_{\mathbf{t},j}\left(\xi^{(j)}_{t_{j}}\right)^{2}\mathrm{d}t_{j}\right)$
$+\sum_{j=1}^{2}\left(h\frac{\mathrm{d}\phi^{\prime}_{\mathbf{t},j}\left(\xi^{(j)}_{t_{j}}\right)}{\phi^{\prime}_{\mathbf{t},j}\left(\xi^{(j)}_{t_{j}}\right)}+\frac{\kappa h(h-1)}{2}\frac{\phi^{\prime\prime}_{\mathbf{t},j}\left(\xi^{(j)}_{t_{j}}\right)^{2}}{\phi^{\prime}_{\mathbf{t},j}\left(\xi^{(j)}_{t_{j}}\right)^{2}}\mathrm{d}t_{j}\right)$
$+\kappa h\sum_{j=1}^{2}\frac{\partial_{j}\mathcal{G}_{\mu}}{\mathcal{G}_{\mu}}\phi^{\prime\prime}_{\mathbf{t},j}\left(\xi^{(j)}_{t_{j}}\right)\mathrm{d}t_{j}$
$=$ $\sum_{j=1}^{2}\left(\frac{\partial_{j}\mathcal{G}_{\mu}}{\mathcal{G}_{\mu}}\phi^{\prime}_{\mathbf{t},j}\left(\xi^{(j)}_{t_{j}}\right)+h\frac{\phi^{\prime\prime}_{\mathbf{t},j}\left(\xi^{(j)}_{t_{j}}\right)}{\phi^{\prime}_{\mathbf{t},j}\left(\xi^{(j)}_{t_{j}}\right)}\right)\mathrm{d}\xi_{t_{j}}^{(j)}.$ (3.12)
In particular, $M_{\mathbf{t}}(\mathcal{G}_{\mu})$ is a two-time-parameter local martingale under $\mathbb{P}$. ∎
Definition 3.6.
Fix $\kappa\in(0,8)$ and $\theta_{1}<\theta_{2}<\theta_{1}+2\pi$. For $\mu\in\mathbb{R}$, we define $\mathcal{G}_{\mu}$ as in (1.3). Let $\mathbb{P}$ denote the probability measure under which $(\eta^{(1)},\eta^{(2)})$ are two independent radial $\operatorname{SLE}_{\kappa}$ in $\mathbb{D}$ starting from $\mathrm{e}^{\mathfrak{i}\theta_{1}}$ and $\mathrm{e}^{\mathfrak{i}\theta_{2}}$ respectively. We define $M_{\mathbf{t}}(\mathcal{G}_{\mu})$ as (3.5) in Lemma 3.5 and denote by $\mathbb{P}(\mathcal{G}_{\mu})$ the probability measure obtained by tilting $\mathbb{P}$ by $M_{\mathbf{t}}(\mathcal{G}_{\mu})$ and call it two-sided radial $\operatorname{SLE}_{\kappa}$ with spiraling rate $\mu$ in $(\mathbb{D};\mathrm{e}^{\mathfrak{i}\theta_{1}},\mathrm{e}^{\mathfrak{i}\theta_{2}};0)$.
From Corollary 3.7 below, it follows that when $\kappa\leq 4$, two-sided radial $\operatorname{SLE}_{\kappa}$ with spiral is well-defined for all time and the two curves do not touch each other before they reach the origin. When $\kappa\in(4,8)$, the above definition for two-sided radial $\operatorname{SLE}_{\kappa}$ with spiral is only defined up to the times the two curves touch each other. When the spiraling rate is 0, we obtain the standard two-sided radial $\operatorname{SLE}$ analyzed in [25, 16, 28, 17].
Corollary 3.7.
Under $\mathbb{P}(\mathcal{G}_{\mu})$, for every $\mathbf{t}=(t_{1},t_{2})$,
  • $g_{\mathbf{t}}(\eta^{(1)})$ is a radial $\operatorname{SLE}_{\kappa}^{\mu}(2)$ in $\mathbb{D}$ starting from $\exp(\mathfrak{i}\theta_{\mathbf{t}}^{(1)})$ with force point at $\exp(\mathfrak{i}\theta_{\mathbf{t}}^{(2)})$,
  • $g_{\mathbf{t}}(\eta^{(2)})$ is a radial $\operatorname{SLE}_{\kappa}^{\mu}(2)$ in $\mathbb{D}$ starting from $\exp(\mathfrak{i}\theta_{\mathbf{t}}^{(2)})$ with force point at $\exp(\mathfrak{i}\theta_{\mathbf{t}}^{(1)})$.
In particular, $\mathbb{P}(\mathcal{G}_{\mu})$ on pairs $(\eta^{(1)},\eta^{(2)})$ satisfies (CI), (DMP), (MARG) and (INT) with
$b_{j}=\kappa\partial_{j}\log\mathcal{G}_{\mu},\quad j=1,2.$ (3.13)
Proof.
Combining (3.12) with (1.3), we have
$\frac{\mathrm{d}M_{\mathbf{t}}(\mathcal{G}_{\mu})}{M_{\mathbf{t}}(\mathcal{G}_{\mu})}=$ $\left(\frac{1}{\kappa}\cot\left((\theta_{\mathbf{t}}^{(1)}-\theta_{\mathbf{t}}^{(2)})/2\right)\phi^{\prime}_{\mathbf{t},1}\left(\xi^{(1)}_{t_{1}}\right)+\frac{\mu}{\kappa}\phi^{\prime}_{\mathbf{t},1}\left(\xi^{(1)}_{t_{1}}\right)+h\frac{\phi^{\prime\prime}_{\mathbf{t},1}\left(\xi^{(1)}_{t_{1}}\right)}{\phi^{\prime}_{\mathbf{t},1}\left(\xi^{(1)}_{t_{1}}\right)}\right)\mathrm{d}\xi_{t_{1}}^{(1)}$
$+\left(\frac{1}{\kappa}\cot\left((\theta_{\mathbf{t}}^{(2)}-\theta_{\mathbf{t}}^{(1)})/2\right)\phi^{\prime}_{\mathbf{t},2}\left(\xi^{(2)}_{t_{2}}\right)+\frac{\mu}{\kappa}\phi^{\prime}_{\mathbf{t},2}\left(\xi^{(2)}_{t_{2}}\right)+h\frac{\phi^{\prime\prime}_{\mathbf{t},2}\left(\xi^{(2)}_{t_{2}}\right)}{\phi^{\prime}_{\mathbf{t},2}\left(\xi^{(2)}_{t_{2}}\right)}\right)\mathrm{d}\xi_{t_{2}}^{(2)}.$
From Girsanov’s theorem and (3.8) and (3.11), under $\mathbb{P}(\mathcal{G}_{\mu})$, we have
$\mathrm{d}\theta_{\mathbf{t}}^{(1)}=$ $\sqrt{\kappa}\phi_{\mathbf{t},1}^{\prime}\left(\xi_{t_{1}}^{(1)}\right)\mathrm{d}\tilde{B}_{t_{1}}^{(1)}+\mu\phi_{\mathbf{t},1}^{\prime}\left(\xi_{t_{1}}^{(1)}\right)^{2}\mathrm{d}t_{1}$
$+\cot\left((\theta_{\mathbf{t}}^{(1)}-\theta_{\mathbf{t}}^{(2)})/2\right)\left(\phi_{\mathbf{t},1}^{\prime}\left(\xi_{t_{1}}^{(1)}\right)^{2}\mathrm{d}t_{1}+\phi_{\mathbf{t},2}^{\prime}\left(\xi_{t_{2}}^{(2)}\right)^{2}\mathrm{d}t_{2}\right),$ (3.14)
$\mathrm{d}\theta_{\mathbf{t}}^{(2)}=$ $\sqrt{\kappa}\phi_{\mathbf{t},2}^{\prime}\left(\xi_{t_{2}}^{(2)}\right)\mathrm{d}\tilde{B}_{t_{2}}^{(2)}+\mu\phi_{\mathbf{t},2}^{\prime}\left(\xi_{t_{2}}^{(2)}\right)^{2}\mathrm{d}t_{2}$
$+\cot\left((\theta_{\mathbf{t}}^{(2)}-\theta_{\mathbf{t}}^{(1)})/2\right)\left(\phi_{\mathbf{t},1}^{\prime}\left(\xi_{t_{1}}^{(1)}\right)^{2}\mathrm{d}t_{1}+\phi_{\mathbf{t},2}^{\prime}\left(\xi_{t_{2}}^{(2)}\right)^{2}\mathrm{d}t_{2}\right),$ (3.15)
where $\tilde{B}^{(1)}$ and $\tilde{B}^{(2)}$ are two independent Brownian motions under $\mathbb{P}(\mathcal{G}_{\mu})$.
Therefore, taking into account the variation of the capacity parametrization (3.7), (3.14) and (3.15) show that under $\mathbb{P}(\mathcal{G}_{\mu})$, for every $\mathbf{t}=(t_{1},t_{2})$, $g_{\mathbf{t}}(\eta^{(1)})$ is a radial $\operatorname{SLE}_{\kappa}^{\mu}(2)$ in $\mathbb{D}$ starting from $\exp(\mathfrak{i}\theta_{\mathbf{t}}^{(1)})$ with force point at $\exp(\mathfrak{i}\theta_{\mathbf{t}}^{(2)})$. Similarly, $g_{\mathbf{t}}(\eta^{(2)})$ is a radial $\operatorname{SLE}_{\kappa}^{\mu}(2)$ in $\mathbb{D}$ starting from $\exp(\mathfrak{i}\theta_{\mathbf{t}}^{(2)})$ with force point at $\exp(\mathfrak{i}\theta_{\mathbf{t}}^{(1)})$. These imply (CI), (DMP), (MARG) and (INT). Eq. (3.13) follows from Remark 3.4. ∎
3.3 Resampling property of two-sided radial SLE with spiral
In this section, we will prove the resampling property of two-sided radial SLE with spiral as we described in Section 1.4 and in Corollary 1.2. We fix $\kappa\in(0,4]$ in the following Theorem 3.8, because we will use the boundary perturbation property in Lemma 3.1 with $\kappa\in(0,4]$ in the proof.
Theorem 3.8 (Resampling property).
Fix $\kappa\in(0,4]$, $\mu\in\mathbb{R}$ and $\theta_{1}<\theta_{2}<\theta_{1}+2\pi$. Suppose $(\eta^{(1)},\eta^{(2)})\sim\mathbb{P}(\mathcal{G}_{\mu})$ is two-sided radial $\operatorname{SLE}_{\kappa}$ with spiraling rate $\mu$ as in Definition 3.6, we have the followings.
  • The marginal law of $\eta^{(1)}$ is radial $\operatorname{SLE}_{\kappa}^{\mu}(2)$ in $\mathbb{D}$ starting from $\mathrm{e}^{\mathfrak{i}\theta_{1}}$ with force point $\mathrm{e}^{\mathfrak{i}\theta_{2}}$ and spiraling rate $\mu$.
  • Given $\eta^{(1)}$, the conditional law of $\eta^{(2)}$ is chordal $\operatorname{SLE}_{\kappa}$ in $\mathbb{D}\setminus\eta^{(1)}$ from $\mathrm{e}^{\mathfrak{i}\theta_{2}}$ to 0.
The same is true when we interchange $\eta^{(1)}$ and $\eta^{(2)}$.
Remark 3.9.
Radial SLE with spiral appears as a flow line in the setup of imaginary geometry [33]. An alternative way to construct a two-sided radial $\operatorname{SLE}_{\kappa}$ with spiraling rate $\mu$ is to take pair of flow lines of
$\Gamma+\frac{(8-\kappa)}{2\sqrt{\kappa}}\arg(\cdot)+\frac{\mu}{\sqrt{\kappa}}\log|\cdot|,$
where $\Gamma$ is a GFF in $\mathbb{D}$ with properly chosen boundary data, and the angles of the two flow lines are also chosen properly. See Figure 2. Using such coupling, one is able to derive the resampling property in Theorem 3.8, see [33, Prop. 3.28]. However, our proof of the resampling property in Section 3.2 does not use the coupling with imaginary geometry. We derive it directly using a refined analysis of the Radon–Nikodym derivative $M_{\mathbf{t}}(\mathcal{G}_{\mu})$ in Definition 3.6.
Proof of Theorem 3.8.
Since $\mathcal{G}_{\mu}(\theta_{1},\theta_{2})=\lambda\,\mathcal{G}_{\mu}(\theta_{2},\theta_{1}+2\pi)$ for some constant $\lambda$ which does not depend on $\theta_{1}$ and $\theta_{2}$, $\eta^{(1)}$ and $\eta^{(2)}$ are interchangeable. Therefore, it suffices to show the bullet points in the statement. The marginal law of $\eta^{(1)}$ is a consequence of Corollary 3.7, and it remains to show the conditional law of $\eta^{(2)}$ given $\eta^{(1)}$.
The law of $\eta^{(1)}_{[0,t_{1}]}$ is the same as radial $\operatorname{SLE}_{\kappa}$ weighted by the following local martingale:
$M_{(t_{1},0)}(\mathcal{G}_{\mu})=$ $\left(g_{t_{1}}^{(1)}\right)^{\prime}(0)^{\frac{3-\mu^{2}}{2\kappa}-\tilde{h}}\times\mathcal{G}_{\mu}\left(\xi_{t_{1}}^{(1)},\phi_{t_{1}}^{(1)}(\theta_{2})\right)\times\left(\phi_{t_{1}}^{(1)}\right)^{\prime}(\theta_{2})^{h}\left(g_{t_{1}}^{(1)}\right)^{\prime}(0)^{\tilde{h}}$.
The law of $(\eta^{(1)},\eta^{(2)})$ is the same as two independent radial $\operatorname{SLE}_{\kappa}$ weighted by the local martingale $M_{\mathbf{t}}(\mathcal{G}_{\mu})$. Therefore, the conditional law of $\eta^{(2)}_{[0,t_{2}]}$ given $\eta^{(1)}_{[0,t_{1}]}$ is the same as a radial $\operatorname{SLE}_{\kappa}$ weighted by
$\frac{M_{(t_{1},t_{2})}(\mathcal{G}_{\mu})}{M_{(t_{1},0)}(\mathcal{G}_{\mu})}=$ $\underbrace{\frac{\phi_{\mathbf{t},2}^{\prime}\left(\xi_{t_{2}}^{(2)}\right)^{h}g_{\mathbf{t},2}^{\prime}(0)^{\tilde{h}}\exp\left(\frac{c}{2}m_{\mathbf{t}}\right)}{\left(\phi_{t_{1}}^{(1)}\right)^{\prime}(\theta_{2})^{h}\left(g_{t_{1}}^{(1)}\right)^{\prime}(0)^{\tilde{h}}}}_{=:P_{\mathbf{t}}}\times g_{\mathbf{t},1}^{\prime}(0)^{\frac{3-\mu^{2}}{2\kappa}}\frac{\mathcal{G}_{\mu}\left(\theta_{\mathbf{t}}^{(1)},\theta_{\mathbf{t}}^{(2)}\right)}{\mathcal{G}_{\mu}\left(\xi_{t_{1}}^{(1)},\phi_{t_{1}}^{(1)}(\theta_{2})\right)}\phi_{\mathbf{t},1}^{\prime}\left(\xi_{t_{1}}^{(1)}\right)^{h}.$
From the boundary perturbation property Lemma 3.1, a radial $\operatorname{SLE}_{\kappa}$ in $(\mathbb{D};\mathrm{e}^{\mathfrak{i}\theta_{2}};0)$ weighted by $P_{\mathbf{t}}$ has the same law as radial $\operatorname{SLE}_{\kappa}$ in $(\mathbb{D}\setminus\eta^{(1)}_{[0,t_{1}]};\mathrm{e}^{\mathfrak{i}\theta_{2}};0)$. Thus, the conditional law of $\eta^{(2)}_{[0,t_{2}]}$ given $\eta^{(1)}_{[0,t_{1}]}$ is the same as a radial $\operatorname{SLE}_{\kappa}$ in $(\mathbb{D}\setminus\eta^{(1)}_{[0,t_{1}]};\mathrm{e}^{\mathfrak{i}\theta_{2}};0)$ weighted by
$g_{\mathbf{t},1}^{\prime}(0)^{\frac{3-\mu^{2}}{2\kappa}}\frac{\mathcal{G}_{\mu}\left(\theta_{\mathbf{t}}^{(1)},\theta_{\mathbf{t}}^{(2)}\right)}{\mathcal{G}_{\mu}\left(\xi_{t_{1}}^{(1)},\phi_{t_{1}}^{(1)}(\theta_{2})\right)}\phi_{\mathbf{t},1}^{\prime}\left(\xi_{t_{1}}^{(1)}\right)^{h}.$ (3.16)
Note that the law of radial $\operatorname{SLE}_{\kappa}(2)$ in $\mathbb{D}\setminus\eta^{(1)}_{[0,t_{1}]}$ from $\mathrm{e}^{\mathfrak{i}\theta_{2}}$ to 0 with force point $\eta^{(1)}_{t_{1}}$ is the same as radial $\operatorname{SLE}_{\kappa}$ in $(\mathbb{D}\setminus\eta^{(1)}_{[0,t_{1}]};\mathrm{e}^{\mathfrak{i}\theta_{2}};0)$ weighted by [41]:
$g_{\mathbf{t},1}^{\prime}(0)^{\frac{3}{2\kappa}}\frac{\mathcal{G}_{0}\left(\theta_{\mathbf{t}}^{(1)},\theta_{\mathbf{t}}^{(2)}\right)}{\mathcal{G}_{0}\left(\xi_{t_{1}}^{(1)},\phi_{t_{1}}^{(1)}(\theta_{2})\right)}\phi_{\mathbf{t},1}^{\prime}\left(\xi_{t_{1}}^{(1)}\right)^{h}.$ (3.17)
Combining (3.16) and (3.17), we see that the conditional law of $\eta^{(2)}_{[0,t_{2}]}$ given $\eta^{(1)}_{[0,t_{1}]}$ is the same as radial $\operatorname{SLE}_{\kappa}(2)$ in $\mathbb{D}\setminus\eta^{(1)}_{[0,t_{1}]}$ from $\mathrm{e}^{\mathfrak{i}\theta_{2}}$ to 0 with force point $\eta^{(1)}_{t_{1}}$ weighted by
$R_{\mathbf{t}}=$ $g_{\mathbf{t},1}^{\prime}(0)^{\frac{-\mu^{2}}{2\kappa}}\frac{\mathcal{G}_{\mu}\left(\theta_{\mathbf{t}}^{(1)},\theta_{\mathbf{t}}^{(2)}\right)}{\mathcal{G}_{\mu}\left(\xi_{t_{1}}^{(1)},\phi_{t_{1}}^{(1)}(\theta_{2})\right)}\frac{\mathcal{G}_{0}\left(\xi_{t_{1}}^{(1)},\phi_{t_{1}}^{(1)}(\theta_{2})\right)}{\mathcal{G}_{0}\left(\theta_{\mathbf{t}}^{(1)},\theta_{\mathbf{t}}^{(2)}\right)}$
$=$ $g_{\mathbf{t},1}^{\prime}(0)^{\frac{-\mu^{2}}{2\kappa}}\exp\left(\frac{\mu}{\kappa}\left(\theta_{\mathbf{t}}^{(1)}+\theta_{\mathbf{t}}^{(2)}\right)-\frac{\mu}{\kappa}\left(\xi_{t_{1}}^{(1)}+\phi_{t_{1}}^{(1)}(\theta_{2})\right)\right).$ (3.18)
We will prove in Lemma 3.10 that the law of radial $\operatorname{SLE}_{\kappa}(2)$ in $\mathbb{D}\setminus\eta^{(1)}_{[0,t_{1}]}$ from $\mathrm{e}^{\mathfrak{i}\theta_{2}}$ to 0 with force point $\eta^{(1)}_{t_{1}}$ converges to chordal $\operatorname{SLE}_{\kappa}$ in $(\mathbb{D}\setminus\eta^{(1)};\mathrm{e}^{\mathfrak{i}\theta_{2}},0)$ as $t_{1}\to\infty$; and we will show in Lemma 3.11 that $R_{\mathbf{t}}\to 1$ almost surely as $t_{1}\to\infty$. Combining these two parts, the conditional law of $\eta^{(2)}_{[0,t_{2}]}$ given $\eta^{(1)}$ is the same as chordal $\operatorname{SLE}_{\kappa}$ as desired. ∎
Lemma 3.10.
Fix $\kappa\in(0,4]$ and $\theta_{1}<\theta_{2}<\theta_{1}+2\pi$. We assume $\eta^{(1)}\in\mathfrak{X}(\mathbb{D};\mathrm{e}^{\mathfrak{i}\theta_{1}};0)$, namely a simple curve in $\mathbb{D}$ from $\mathrm{e}^{\mathfrak{i}\theta_{1}}$ to 0.
  • For $t_{1}\in(0,\infty)$, we denote by $Q_{t_{1}}$ the law of a radial $\operatorname{SLE}_{\kappa}(2)$ in $\mathbb{D}\setminus\eta^{(1)}_{[0,t_{1}]}$ from $\mathrm{e}^{\mathfrak{i}\theta_{2}}$ to 0 with force point $\eta^{(1)}_{t_{1}}$.
  • We denote by $Q_{\infty}$ the conditional law of a chordal $\operatorname{SLE}_{\kappa}$ in $(\mathbb{D}\setminus\eta^{(1)};\mathrm{e}^{\mathfrak{i}\theta_{2}},0)$.
Suppose $U$ is any neighborhood of $\mathrm{e}^{\mathfrak{i}\theta_{2}}$ in $\mathbb{D}\setminus\eta^{(1)}$ such that there is a positive distance between $U$ and $\eta^{(1)}$ and let $\tau$ be the first time $\eta^{(2)}$ exits $U$. Then the law of the curve $\eta^{(2)}$ restricted to $0,\tau]$ under $Q_{t_{1}}$ and $Q_{\infty}$ are absolutely continuous. We have
$\lim_{t_{1}\to\infty}\frac{\mathrm{d}Q_{t_{1}}}{\mathrm{d}Q_{\infty}}\left(\eta^{(2)}_{[0,\tau]}\right)=1,\quad Q_{\infty}-a.s.$
Proof.
The convergence of radial $\operatorname{SLE}_{\kappa}$ in $(\mathbb{D}\setminus\eta^{(1)}_{[0,t_{1}]};\mathrm{e}^{\mathfrak{i}\theta_{2}};0)$ to chordal $\operatorname{SLE}_{\kappa}$ in $(\mathbb{D}\setminus\eta^{(1)};\mathrm{e}^{\mathfrak{i}\theta_{2}},0)$ is well known, see e.g. [21, Lem. 3.2] and [26]. In our setup, we need the convergence of radial $\operatorname{SLE}_{\kappa}(2)$ process.
We fix a sequence of conformal maps $f_{t_{1}}:\mathbb{D}\setminus\eta^{(1)}_{[0,t_{1}]}\to\mathbb{D}$ and a conformal map $f_{\infty}:\mathbb{D}\setminus\eta^{(1)}\to\mathbb{D}$ such that
$f_{t_{1}}(\eta^{(1)}_{t_{1}})=\mathrm{e}^{\mathfrak{i}\theta_{1}},\quad f_{t_{1}}(\mathrm{e}^{\mathfrak{i}\theta_{2}})=\mathrm{e}^{\mathfrak{i}\theta_{2}},\qquad f_{\infty}(0)=\mathrm{e}^{\mathfrak{i}\theta_{1}},\quad f_{\infty}(\mathrm{e}^{\mathfrak{i}\theta_{2}})=\mathrm{e}^{\mathfrak{i}\theta_{2}};$
and that $f_{t_{1}}$ converges to $f_{\infty}$ locally uniformly. Note that $(f_{t_{1}})_{*}(Q_{t_{1}})$ is the same as radial $\operatorname{SLE}_{\kappa}(2)$ in $\mathbb{D}$ from $\mathrm{e}^{\mathfrak{i}\theta_{2}}$ to $w=f_{t_{1}}(0)$ with force point $\mathrm{e}^{\mathfrak{i}\theta_{1}}$, and $(f_{\infty})_{*}(Q_{\infty})$ is the same as chordal $\operatorname{SLE}_{\kappa}$ in $(\mathbb{D};\mathrm{e}^{\mathfrak{i}\theta_{2}},\mathrm{e}^{\mathfrak{i}\theta_{1}})$.
The law of radial $\operatorname{SLE}_{\kappa}(2)$ in $\mathbb{D}$ from $\mathrm{e}^{\mathfrak{i}\theta_{2}}$ to $w$ with force point $\mathrm{e}^{\mathfrak{i}\theta_{1}}$ is the same as the law of radial $\operatorname{SLE}_{\kappa}$ in $(\mathbb{D};\mathrm{e}^{\mathfrak{i}\theta_{2}};0)$ weighted by the local martingale [41]:
$g_{t}^{\prime}(0)^{-\tilde{h}}\phi_{t}^{\prime}(\theta_{1})^{h}\left(\sin((\xi_{t}-\phi_{t}(\theta_{1}))/2)\right)^{\frac{2}{\kappa}}$
$\times|g_{t}^{\prime}(w)|^{\tilde{h}+\frac{3}{2\kappa}}(1-|g_{t}(w)|^{2})^{\frac{(\kappa-8)^{2}}{8\kappa}}|g_{t}(w)-\mathrm{e}^{\mathfrak{i}\xi_{t}}|^{1-\frac{8}{\kappa}}|g_{t}(w)-\mathrm{e}^{\mathfrak{i}\phi_{t}(\theta_{1})}|^{1-\frac{8}{\kappa}}.$ (3.19)
The law of chordal $\operatorname{SLE}_{\kappa}$ in $(\mathbb{D};\mathrm{e}^{\mathfrak{i}\theta_{2}},\mathrm{e}^{\mathfrak{i}\theta_{1}})$ is the same as the law of radial $\operatorname{SLE}_{\kappa}$ in $(\mathbb{D};\mathrm{e}^{\mathfrak{i}\theta_{2}};0)$ weighted by the local martingale [41]:
$g_{t}^{\prime}(0)^{-\tilde{h}}\phi_{t}^{\prime}(\theta_{1})^{h}\left(\sin((\xi_{t}-\phi_{t}(\theta_{1}))/2)\right)^{-2h}.$ (3.20)
Combining (3.3) and (3.20), the law $(f_{t_{1}})_{*}(Q_{t_{1}})$ is the same as $(f_{\infty})_{*}(Q_{\infty})$ weighted by the local martingale
$M_{t}(w)=$ $\left(\frac{\sin((\xi_{t}-\phi_{t}(\theta_{1}))/2)}{\sin((\theta_{2}-\theta_{1})/2)}\right)^{\frac{8}{\kappa}-1}|g_{t}^{\prime}(w)|^{\tilde{h}+\frac{3}{2\kappa}}\left(\frac{1-|g_{t}(w)|^{2}}{1-|w|^{2}}\right)^{\frac{(\kappa-8)^{2}}{8\kappa}}$
$\times\left|\frac{g_{t}(w)-\mathrm{e}^{\mathfrak{i}\xi_{t}}}{w-\mathrm{e}^{\mathfrak{i}\theta_{2}}}\right|^{1-\frac{8}{\kappa}}\left|\frac{g_{t}(w)-\mathrm{e}^{\mathfrak{i}\phi_{t}(\theta_{1})}}{w-\mathrm{e}^{\mathfrak{i}\theta_{1}}}\right|^{1-\frac{8}{\kappa}}.$ (3.21)
As $t_{1}\to\infty$, we have $w=f_{t_{1}}(0)\to\mathrm{e}^{\mathfrak{i}\theta_{1}}$. Thus,
$|g_{t}^{\prime}(w)|\to|g_{t}^{\prime}(\mathrm{e}^{\mathfrak{i}\theta_{1}})|,\quad\frac{1-|g_{t}(w)|^{2}}{1-|w|^{2}}\to|g_{t}^{\prime}(\mathrm{e}^{\mathfrak{i}\theta_{1}})|,$
$\left|\frac{g_{t}(w)-\mathrm{e}^{\mathfrak{i}\xi_{t}}}{w-\mathrm{e}^{\mathfrak{i}\theta_{2}}}\right|\to\frac{\sin((\xi_{t}-\phi_{t}(\theta_{1}))/2)}{\sin((\theta_{2}-\theta_{1})/2)},\quad\left|\frac{g_{t}(w)-\mathrm{e}^{\mathfrak{i}\phi_{t}(\theta_{1})}}{w-\mathrm{e}^{\mathfrak{i}\theta_{1}}}\right|\to|g_{t}^{\prime}(\mathrm{e}^{\mathfrak{i}\theta_{1}})|.$
Consequently, $M_{t}(w)\to 1$ almost surely. These give the desired conclusion. ∎
Lemma 3.11.
Assume the same notations as in the proof of Theorem 3.8 and recall that $R_{\mathbf{t}}$ is defined in (3.3). We have
$\lim_{t_{1}\to\infty}R_{\mathbf{t}}=1,\quad a.s.$
Proof.
We write $I_{t_{1}}$ for the arc in $\partial\mathbb{D}$ that is the image of both sides of $\eta^{(1)}_{[0,t_{1}]}$ under the conformal map $g_{t_{1}}^{(1)}$ extended to the boundary, see Figure 5. It is easy to see the harmonic measure of $\partial\mathbb{D}$ seen from 0 of the domain $\mathbb{D}\setminus\eta^{(1)}_{[0,t_{1}]}$ is decreasing to 0 as $t_{1}\to\infty$. Therefore,
$|I_{t_{1}}^{c}|=|\partial\mathbb{D}\setminus I_{t_{1}}|\xrightarrow[]{t_{1}\to\infty}0$
where $I_{t_{1}}^{c}=\partial\mathbb{D}\setminus I_{t_{1}}$.
Lemma 3.3 shows that $\eta^{(2)}_{[0,t_{2}]}$ is at positive distance from $\eta^{(1)}_{[0,\infty)}$. Hence, there exists $\lambda\in(0,1)$ such that the neighborhood
$U=\left\{z\in\mathbb{D}\,|\,\mathbb{P}_{z}\left(\beta\text{ hits }\eta^{(1)}_{[0,\infty)}\text{ before exiting }\mathbb{D}\right)\geq\lambda\right\}$
satisfies $\eta^{(2)}_{[0,t_{2}]}\cap U=\emptyset$, where $\mathbb{P}_{z}$ is the law of a two-dimensional Brownian motion $\beta$ starting from $z$. Let
$U_{t_{1}}=\left\{z\in\mathbb{D}\,|\,\mathbb{P}_{z}\left(\beta\text{ hits }\eta^{(1)}_{[0,t_{1})}\text{ before exiting }\mathbb{D}\right)\geq\lambda\right\}\subset U.$
The image $\tilde{U}_{t_{1}}:=g_{t_{1}}^{(1)}(U_{t_{1}})$ is bounded by $I_{t_{1}}$ and a circular arc $\alpha_{t_{1}}$ meeting the endpoints of $I_{t_{1}}$ with angle $\lambda/\pi$. Since $|I^{c}_{t_{1}}|\to 0$, the diameter of the domain $\mathbb{D}\setminus\tilde{U}_{t_{1}}$, which contains $\tilde{\eta}^{(2)}_{[0,t_{2}]}=g_{t_{1}}^{(1)}\left(\eta^{(2)}_{[0,t_{2}]}\right)$, converges to 0.
Figure 7: Illustration of the domains $U_{t_{1}}$ and $\tilde{U}_{t_{1}}:=g_{t_{1}}^{(1)}(U_{t_{1}})$ and the rotation map $\rho_{t_{1}}$.
Recall that the map $g_{\mathbf{t},1}$ maps out the curve $\tilde{\eta}^{(2)}_{[0,t_{2}]}=g_{t_{1}}^{(1)}\left(\eta_{[0,t_{2}]}^{(2)}\right)$. If we conjugate $g_{\mathbf{t},1}$ by the rotation $\rho_{t_{1}}:\mathbb{D}\to\mathbb{D}$ such that the image of the mid-point of $\alpha_{t_{1}}$ under $\rho_{t_{1}}$ lies in $(0,1)$ (so that $R(\alpha_{t_{1}})$ is symmetric with respect to the real line and $\rho_{t_{1}}(\alpha_{t_{1}})$ shrinks to the point $1\in\partial\mathbb{D}$), the map $\tilde{g}_{\mathbf{t},1}:=\rho_{t_{1}}\circ g_{\mathbf{t},1}\circ\rho_{t_{1}}^{-1}$ converges in Carathéodory topology (namely, uniformly on compact subsets) to the identity map in $\mathbb{D}$ as $t_{1}\to\infty$. If we Schwarz-reflect $g_{\mathbf{t},1}$ along $\partial\mathbb{D}\setminus g_{t_{1}}^{(1)}(\exp(\mathfrak{i}\theta_{2}))$, we see that the convergence also extends to the boundary, more precisely, we obtain that $\tilde{g}_{\mathbf{t},1}$ converges uniformly on all compact subsets of $\overline{\mathbb{D}}\setminus\{1\}$ (and the map is well-defined on every such compact subset for large enough $t_{1}$), so do the derivatives of $\tilde{g}_{\mathbf{t},1}$ with respect to $z$. Thus, we obtain that $R_{\mathbf{t}}\to 1$ almost surely, which completes the proof. ∎
3.4 Chordal SLE weighted by conformal radius
In this section, we show that the partition functions $\mathcal{Z}_{\alpha}$ correspond to the chordal SLE weighted by the conformal radius to the power $\alpha$.
For this, we first calculate the Laplace transform of the conformal radius of the complement of chordal SLE. Usually, chordal SLE is defined in the upper-half plane. It is more convenient here to describe it in the unit disc $\mathbb{D}$ via a change of coordinate. Fix $\kappa\in(0,8)$ and $\theta_{1}<\theta_{2}<\theta_{1}+2\pi$. Suppose $\gamma$ is chordal $\operatorname{SLE}_{\kappa}$ in $(\mathbb{D};\mathrm{e}^{\mathfrak{i}\theta_{1}},\mathrm{e}^{\mathfrak{i}\theta_{2}})$. We parameterize it by the capacity and define $g_{t},\xi_{t}$ accordingly as in Section 3.1. Denote by $T$ the first time $\gamma$ disconnects $\mathrm{e}^{\mathfrak{i}\theta_{2}}$ from the origin. A chordal $\operatorname{SLE}_{\kappa}$ in $(\mathbb{D};\mathrm{e}^{\mathfrak{i}\theta_{1}},\mathrm{e}^{\mathfrak{i}\theta_{2}})$, up to $T$, has the same law as radial $\operatorname{SLE}_{\kappa}(\kappa-6)$ starting from $\mathrm{e}^{\mathfrak{i}\theta_{1}}$ with force point $\mathrm{e}^{\mathfrak{i}\theta_{2}}$, up to the same time, see [41]. In other words, its driving function $\xi_{t}$ solves the following SDE:
$\begin{cases}\xi_{0}=\theta_{1},V_{0}=\theta_{2},\\ \mathrm{d}\xi_{t}=\sqrt{\kappa}\mathrm{d}B_{t}+\dfrac{\kappa-6}{2}\cot\left((\xi_{t}-V_{t})/2\right)\mathrm{d}t,\\ \mathrm{d}V_{t}=\cot\left((V_{t}-\xi_{t})/2\right)\mathrm{d}t.\end{cases}$ (3.22)
Note that the conformal radius $\mathrm{CR}(\mathbb{D}\setminus\gamma)$ is the same as $\mathrm{e}^{-T}$; thus its Laplace transform can be derived from the SDE (3.22).
Lemma 3.12.
Fix $\kappa\in(0,8)$ and $\theta_{1}<\theta_{2}<\theta_{1}+2\pi$. We denote $\theta=\theta_{2}-\theta_{1}\in(0,2\pi)$. Suppose $\gamma$ is chordal $\operatorname{SLE}_{\kappa}$ in $(\mathbb{D};\mathrm{e}^{\mathfrak{i}\theta_{1}},\mathrm{e}^{\mathfrak{i}\theta_{2}})$ and denote by $\mathbb{E}_{\theta}$ the expectation with respect to $\gamma$. Denote by $\mathrm{CR}(\mathbb{D}\setminus\gamma)$ the conformal radius of $\mathbb{D}\setminus\gamma$ seen from the origin. For $\alpha\in\mathbb{R}$, we define
$\Phi(\kappa,\alpha;u):=\mathbb{E}_{\theta}\left[\mathrm{CR}(\mathbb{D}\setminus\gamma)^{-\alpha}\right],\quad\text{where }u=\left(\sin(\theta/4)\right)^{2}\in(0,1).$ (3.23)
Then $\Phi(\kappa,\alpha;u)$ is finite for $u\in(0,1)$ if only and if $\alpha<1-\kappa/8$. Moreover, when $\alpha<1-\kappa/8$, $\Phi(u)=\Phi(\kappa,\alpha;u)$ satisfies the following ODE
$u(1-u)\Phi^{\prime\prime}+\frac{3\kappa-8}{2\kappa}(1-2u)\Phi^{\prime}+\frac{8\alpha}{\kappa}\Phi=0,$ (3.24)
and the symmetry
$\Phi(u)=\Phi(1-u),\quad u\in(0,1).$ (3.25)
Proof.
We first show that $\Phi(\kappa,\alpha;u)$ is finite as long as $\alpha<1-\kappa/8$. This is done in [36, Proof of Prop. 3.5]. For the reader’s convenience, we summarize its proof here.
When $\alpha\leq 0$, since $\mathrm{CR}(\mathbb{D}\setminus\gamma)\leq 1$ by Schwarz lemma, we obtain immediately that $\Phi(\kappa,\alpha;u)<\infty$.
When $\alpha\in(0,1-\kappa/8)$, we will derive $\Phi$ in terms of hypergeometric functions. We set
$A=1-\frac{4}{\kappa}+\sqrt{\left(1-\frac{4}{\kappa}\right)^{2}+\frac{8\alpha}{\kappa}},\quad B=1-\frac{4}{\kappa}-\sqrt{\left(1-\frac{4}{\kappa}\right)^{2}+\frac{8\alpha}{\kappa}},\quad C=\frac{3}{2}-\frac{4}{\kappa}.$
Assume $C\not\in\mathbb{Z}$ and define
$f_{1}(u):=\;_{2}\mathrm{F}_{1}(A,B,C;u),\quad f_{2}(u):=u^{1-C}\;_{2}\mathrm{F}_{1}(1+A-C,1+B-C,2-C,u),$
where $_{2}\mathrm{F}_{1}$ is the hypergeometric function (see e.g. [1, Eq.(15.1.1)]). Note that $f_{1},f_{2}$ are two linearly independent solutions to ODE (3.24). Let us check the values of $f_{1},f_{2}$ at the endpoints $u=0$ or $u=1$. Since $\kappa\in(0,8)$ and $\alpha\in(0,1-\kappa/8)$ and $C\not\in\mathbb{Z}$, we have
$A<1,\quad B\in(1-8/\kappa,1-4/\kappa],\quad C\in(-\infty,1)\setminus\mathbb{Z},\quad C>A+B.$
From [1, Eq.(15.1.20)], we have
$f_{1}(0)=$ $1,\quad f_{1}(1)=\frac{\Gamma(C)\Gamma(C-A-B)}{\Gamma(C-A)\Gamma(C-B)}=\frac{\cos\left(\pi\sqrt{\left(1-\frac{4}{\kappa}\right)^{2}+\frac{8\alpha}{\kappa}}\right)}{\cos\left(\pi\left(1-\frac{4}{\kappa}\right)\right)};$
$f_{2}(0)=$ $0,\quad f_{2}(1)=\frac{\Gamma(2-C)\Gamma(1-C)}{\Gamma(1-A)\Gamma(1-B)}\in(0,\infty).$
We parameterize $\gamma$ by the capacity, then its driving function $\xi_{t}$ solves SDE (3.22). We denote $\theta_{t}=V_{t}-\xi_{t}$. The process $\theta_{t}$ satisfies the SDE:
$\mathrm{d}\theta_{t}=\sqrt{\kappa}\mathrm{d}B_{t}+\frac{\kappa-4}{2}\cot(\theta_{t}/2)\mathrm{d}t.$ (3.26)
The disconnection time $T$ is the first time that $\theta_{t}$ hits 0 or $2\pi$. Suppose $f$ is an analytic function defined on $(0,1)$. Then $\mathrm{e}^{\alpha t}f\left(\left(\sin(\theta_{t}/4)\right)^{2}\right)$ is a local martingale if and only if $f$ satisfies (3.24). Since $f_{1},f_{2}$ are solutions to this ODE, the processes
$\mathrm{e}^{\alpha t}f_{1}\left(\left(\sin(\theta_{t}/4)\right)^{2}\right)\quad\text{and}\quad\mathrm{e}^{\alpha t}f_{2}\left(\left(\sin(\theta_{t}/4)\right)^{2}\right)$
are local martingales. These martingales are also considered in [40]. Since $f_{1},f_{2}$ are finite at $u=0$ and $u=1$, and the lifetime $T$ has finite expectation, we may conclude that these two local martingales are martingales up to $T$. Then the optional stopping theorem gives
$\begin{cases}\mathbb{E}_{\theta}\left[\mathrm{e}^{\alpha T}1_{\{\theta_{T}=0\}}\right]+f_{1}(1)\mathbb{E}_{\theta}\left[\mathrm{e}^{\alpha T}1_{\{\theta_{T}=2\pi\}}\right]=f_{1}\left(\left(\sin(\theta/4)\right)^{2}\right);\\ f_{2}(1)\mathbb{E}_{\theta}\left[\mathrm{e}^{\alpha T}1_{\{\theta_{T}=2\pi\}}\right]=f_{2}\left(\left(\sin(\theta/4)\right)^{2}\right).\end{cases}$
As $\mathrm{CR}(\mathbb{D}\setminus\gamma)=\mathrm{e}^{-T}$, the above relation gives
$\mathbb{E}_{\theta}\left[\mathrm{CR}(\mathbb{D}\setminus\gamma)^{-\alpha}\right]=f_{1}\left(\left(\sin(\theta/4)\right)^{2}\right)+\frac{1-f_{1}(1)}{f_{2}(1)}f_{2}\left(\left(\sin(\theta/4)\right)^{2}\right).$ (3.27)
In particular, this implies that $\Phi(\kappa,\alpha;u)$ is finite for $u\in(0,1)$ when
$\kappa\in(0,8),\quad C=\frac{3}{2}-\frac{4}{\kappa}\not\in\mathbb{Z},\quad\alpha\in(0,1-\kappa/8).$
As $\Phi(\kappa,\alpha;u)$ is continuous in $\kappa\in(0,8)$ and is increasing in $\alpha$, we conclude that $\Phi(\kappa,\alpha;u)$ is finite for $u\in(0,1)$ when
$\kappa\in(0,8),\quad\alpha<1-\kappa/8.$
Moreover, when $\alpha\in(0,1-\kappa/8)$ and $C=3/2-4/\kappa\not\in\mathbb{Z}$,
$\Phi(u)=\Phi(\kappa,\alpha;u)=\mathbb{E}_{\theta}[\mathrm{CR}(\mathbb{D}\setminus\gamma)^{-\alpha}]=f_{1}(u)+\frac{1-f_{1}(1)}{f_{2}(1)}f_{2}(u)$
satisfies (3.24). In fact, $\Phi$ satisfies (3.24) for all $\kappa\in(0,8)$ and $\alpha<1-\kappa/8$. Note that
$\mathrm{e}^{\alpha t}\Phi\left(\left(\sin(\theta_{t}/4)\right)^{2}\right)=\mathbb{E}_{\theta}\left[\mathrm{CR}(\mathbb{D}\setminus\gamma)^{-\alpha}\;|\;\gamma_{[0,t]}\right]$ (3.28)
is a martingale and $\theta_{t}$ satisfies (3.26). Thus, $\Phi$ is a weak solution for (3.24) and
$\left(\sin((\theta_{2}-\theta_{1})/2)\right)^{-2h}\Phi\left(\left(\sin((\theta_{2}-\theta_{1})/4)\right)^{2}\right)$
is a weak solution to the radial BPZ equations. As the operators in the radial BPZ equations are hypoelliptic, see Remark 2.8, weak solutions are strong solutions. Thus $\Phi$ is a $C^{2}$ solution to (3.24) for all $\kappa\in(0,8)$ and $\alpha<1-\kappa/8$. The symmetry in (3.25) is clear from the definition.
Finally, let us consider the case when $\alpha\geq 1-\kappa/8$. Fix $\kappa\in(0,8)$ with $C\not\in\mathbb{Z}$. As $\alpha\uparrow(1-\kappa/8)$, we have
$A\to 1,\quad B\to(1-8/\kappa),$
$f_{1}(u)\to\;_{2}\mathrm{F}_{1}(1,1-8/\kappa,3/2-4/\kappa;u)\in(-\infty,\infty),\quad$ $f_{1}(1)\to-1,$
$f_{2}(u)\to u^{1-C}\;_{2}\mathrm{F}_{1}(1+4/\kappa,1/2-4/\kappa,1/2+4/\kappa;u)\neq 0,\quad$ $f_{2}(1)\to 0.$
Plugging into (3.27), we see that
$\Phi(\kappa,\alpha;u)\uparrow\infty,\quad\text{as }\alpha\uparrow(1-\kappa/8).$
Note that $\Phi(\kappa,\alpha;u)$ is increasing in $\alpha$. This completes the proof. ∎
Corollary 3.13.
Fix $\kappa\in(0,8)$ and $\theta_{1}<\theta_{2}<\theta_{1}+2\pi$. We denote $\theta=\theta_{2}-\theta_{1}\in(0,2\pi)$. Suppose $\gamma$ is chordal $\operatorname{SLE}_{\kappa}$ in $(\mathbb{D};\mathrm{e}^{\mathfrak{i}\theta_{1}},\mathrm{e}^{\mathfrak{i}\theta_{2}})$ and denote by $\mathbb{E}_{\theta}$ the expectation with respect to $\gamma$. Recall that $\mathcal{Z}_{\alpha}$ is defined in (1.4) for $\alpha<1-\kappa/8$ and $\mathcal{G}_{\mu}$ is defined in (1.3) for $\mu\in\mathbb{R}$. Recall that $h=\frac{6-\kappa}{2\kappa}$ from (3.1). Then we have
$\mathcal{Z}_{\alpha}(\theta_{1},\theta_{2})=\left(\sin(\theta/2)\right)^{-2h}\frac{\mathbb{E}_{\theta}[\mathrm{CR}(\mathbb{D}\setminus\gamma)^{-\alpha}]}{\mathbb{E}_{\pi}[\mathrm{CR}(\mathbb{D}\setminus\gamma)^{-\alpha}]},\quad\text{for }\alpha<1-\kappa/8.$ (3.29)
Moreover, we have
$\frac{\mathbb{E}_{\theta}[\mathrm{CR}(\mathbb{D}\setminus\gamma)^{-\alpha}]}{\mathbb{E}_{\pi}[\mathrm{CR}(\mathbb{D}\setminus\gamma)^{-\alpha}]}=\left(\sin(\theta/2)\right)^{2h}\mathcal{Z}_{\alpha}(\theta_{1},\theta_{2})\to\left(\sin(\theta/2)\right)^{2h}\mathcal{G}_{0}(\theta_{1},\theta_{2})$ (3.30)
as $\alpha\uparrow(1-\kappa/8)$.
Proof.
We denote $u=(\sin(\theta/4))^{2}$. Recall from (1.4) that $\phi_{\alpha}$ is the unique solution to (1.5). Comparing with (3.24) and (3.25), it is clear that
$\phi_{\alpha}(\cdot)=\frac{\Phi(\kappa,\alpha;\cdot)}{\Phi(\kappa,\alpha;1/2)}.$
Thus
$\mathcal{Z}_{\alpha}(\theta_{1},\theta_{2}):=\left(\sin(\theta/2)\right)^{-2h}\phi_{\alpha}(u)=\left(\sin(\theta/2)\right)^{-2h}\frac{\mathbb{E}_{\theta}[\mathrm{CR}(\mathbb{D}\setminus\gamma)^{-\alpha}]}{\mathbb{E}_{\pi}[\mathrm{CR}(\mathbb{D}\setminus\gamma)^{-\alpha}]},$
as desired in (3.29). Moreover, we have
$\frac{\mathbb{E}_{\theta}[\mathrm{CR}(\mathbb{D}\setminus\gamma)^{-\alpha}]}{\mathbb{E}_{\pi}[\mathrm{CR}(\mathbb{D}\setminus\gamma)^{-\alpha}]}$ $=\left(\sin(\theta/2)\right)^{2h}\mathcal{Z}_{\alpha}(\theta_{1},\theta_{2})=\phi_{\alpha}(u)$
which converges to $\phi_{\alpha_{0}}(u)=\left(\sin(\theta/2)\right)^{2h}\mathcal{G}_{0}(\theta_{1},\theta_{2})$ as $\alpha\to\alpha_{0}=1-\kappa/8$ by Lemma A.1. This gives (3.30). ∎
Corollary 3.14.
Fix $\kappa\in(0,8)$ and $\alpha<1-\kappa/8$. Denote by $\mathbb{P}$ the law of $\gamma$ chordal $\operatorname{SLE}_{\kappa}$ in $(\mathbb{D};\mathrm{e}^{\mathfrak{i}\theta_{1}},\mathrm{e}^{\mathfrak{i}\theta_{2}})$ with $\theta_{1}<\theta_{2}<\theta_{1}+2\pi$. Let $\eta^{(1)}$ be $\gamma$ and let $\eta^{(2)}$ be the time-reversal of $\gamma$ and still denote by $\mathbb{P}$ the induced law on $(\eta^{(1)},\eta^{(2)})$. We define $\mathcal{Z}_{\alpha}$ as in (1.4). Denote by $\mathbb{P}(\mathcal{Z}_{\alpha})$ the probability measure obtained by weighting $\mathbb{P}$ by $\mathrm{CR}(\mathbb{D}\setminus\gamma)^{-\alpha}$. Then, under $\mathbb{P}(\mathcal{Z}_{\alpha})$, the family of local laws obtained by restricting the pair $(\eta^{(1)},\eta^{(2)})$ in disjoint neighborhoods satisfies (CI), (DMP), (MARG) and (INT) with
$b_{j}=\kappa\partial_{j}\log\mathcal{Z}_{\alpha},\quad j=1,2.$
More precisely, the driving function of $\eta^{(1)}$ solves the following SDE, up to the first time $\mathrm{e}^{\mathfrak{i}\theta_{2}}$ is disconnected from the origin:
$\begin{cases}\xi_{0}^{(1)}=\theta_{1},V_{0}^{(2)}=\theta_{2},\\ \mathrm{d}\xi_{t}^{(1)}=\sqrt{\kappa}\mathrm{d}\tilde{B}_{t}^{(1)}+\kappa\partial_{1}(\log\mathcal{Z}_{\alpha})(\xi_{t}^{(1)},V_{t}^{(2)})\mathrm{d}t,\\ \mathrm{d}V_{t}^{(2)}=\cot\left((V_{t}^{(2)}-\xi_{t}^{(1)})/2\right)\mathrm{d}t,\end{cases}$ (3.31)
where $\tilde{B}^{(1)}$ is Brownian motion under $\mathbb{P}(\mathcal{Z}_{\alpha})$. Similarly, the driving function of $\eta^{(2)}$ solves the following SDE, up to the first time $\mathrm{e}^{\mathfrak{i}\theta_{1}}$ is disconnected from the origin:
$\begin{cases}V_{0}^{(1)}=\theta_{1},\xi_{0}^{(2)}=\theta_{2},\\ \mathrm{d}\xi_{t}^{(2)}=\sqrt{\kappa}\mathrm{d}\tilde{B}_{t}^{(2)}+\kappa\partial_{2}(\log\mathcal{Z}_{\alpha})(V_{t}^{(1)},\xi_{t}^{(2)})\mathrm{d}t,\\ \mathrm{d}V_{t}^{(1)}=\cot\left((V_{t}^{(1)}-\xi_{t}^{(2)})/2\right)\mathrm{d}t,\end{cases}$ (3.32)
where $\tilde{B}^{(2)}$ is Brownian motion under $\mathbb{P}(\mathcal{Z}_{\alpha})$.
Proof.
The fact that the local laws obtained by restricting the pair $(\eta^{(1)},\eta^{(2)})\sim\mathbb{P}(\mathcal{Z}_{\alpha})$ in disjoint neighborhoods satisfies (CI), (DMP), (MARG) and (INT) follows from the reversibility of SLE (proved in [44, 31, 32]): suppose $\gamma$ is chordal $\operatorname{SLE}_{\kappa}$ in $\mathbb{D}$ from $\mathrm{e}^{\mathfrak{i}\theta_{1}}$ to $\mathrm{e}^{\mathfrak{i}\theta_{2}}$ with $\kappa\in(0,8)$, the time-reversal of $\gamma$ has the same law as chordal $\operatorname{SLE}_{\kappa}$ in $\mathbb{D}$ from $\mathrm{e}^{\mathfrak{i}\theta_{2}}$ to $\mathrm{e}^{\mathfrak{i}\theta_{1}}$. It remains to check (3.31) and (3.32). As the pair $(\eta^{(1)},\eta^{(2)})$ is interchangeable, it suffices to check (3.31).
Denote $\Phi(\cdot)=\Phi(\kappa,\alpha;\cdot)$ as in (3.23). Using the same notations as in the proof of Lemma 3.12, we denote the martingale in (3.28) by
$M_{t}(\alpha):=\mathrm{e}^{\alpha t}\Phi\left(\left(\sin(\theta_{t}/4)\right)^{2}\right).$
Then $\mathbb{P}(\mathcal{Z}_{\alpha})$ is the same as $\mathbb{P}$ tilting by $M_{t}(\alpha)$. Recall from (3.26), under $\mathbb{P}$, we have
$\mathrm{d}\theta_{t}=\sqrt{\kappa}\mathrm{d}B_{t}+\frac{\kappa-4}{2}\cot(\theta_{t}/2)\mathrm{d}t.$
Thus, under $\mathbb{P}$, we have
$\frac{\mathrm{d}M_{t}(\alpha)}{M_{t}(\alpha)}=\frac{\sqrt{\kappa}}{4}\frac{\Phi^{\prime}}{\Phi}\sin(\theta_{t}/2)\mathrm{d}B_{t}.$
Girsanov’s theorem tells that
$\tilde{B}_{t}=B_{t}-\frac{\sqrt{\kappa}}{4}\frac{\Phi^{\prime}}{\Phi}\sin(\theta_{t}/2)\mathrm{d}t$
is Brownian motion under $\mathbb{P}(\mathcal{Z}_{\alpha})$. Combining with (3.29), we obtain (3.31). ∎
Now, we are ready to complete the proof of Theorem 1.1.
Proof of Theorem 1.1.
The conclusion follows from Theorem 2.4, Corollary 3.7 and Corollary 3.14. ∎
3.5 Commuting SLEs without interchangeability
We may also classify the commuting SLEs without interchangeability. Let us first give an example.
Remark 3.15.
Using the same notations as in Lemma 3.12, for $\kappa\in(0,8)$ and $\alpha<1-\kappa/8$, we consider
$\Phi^{L}(\kappa,\alpha;u):=$ $\mathbb{E}_{\theta}\left[\mathrm{CR}(\mathbb{D}\setminus\gamma)^{-\alpha}1_{\{0\text{ is to the left of }\gamma\}}\right];$
$\Phi^{R}(\kappa,\alpha;u):=$ $\mathbb{E}_{\theta}\left[\mathrm{CR}(\mathbb{D}\setminus\gamma)^{-\alpha}1_{\{0\text{ is to the right of }\gamma\}}\right].$
Then $\Phi^{L}(\cdot)=\Phi^{L}(\kappa,\alpha;\cdot)$ and $\Phi^{R}(\cdot)=\Phi^{R}(\kappa,\alpha;\cdot)$ also satisfy the ODE (3.24), but they do not enjoy the symmetry (3.25) anymore. This gives an example of locally commuting 2-radial SLE without interchangeability.
Moreover, we denote $\theta=\theta_{2}-\theta_{1}$ and set
$\mathcal{Z}^{L}(\theta_{1},\theta_{2})=\left(\sin(\theta/2)\right)^{-2h}\Phi^{L}(\kappa,\alpha;\left(\sin(\theta/4)\right)^{2});$
$\mathcal{Z}^{R}(\theta_{1},\theta_{2})=\left(\sin(\theta/2)\right)^{-2h}\Phi^{R}(\kappa,\alpha;\left(\sin(\theta/4)\right)^{2}).$
Then both $\mathcal{Z}^{L}$ and $\mathcal{Z}^{R}$ satisfy (2.14) and (2.15) with
$F=\frac{(6-\kappa)(\kappa-2)}{8\kappa}-\alpha.$
Proposition 3.16.
If one removes the interchangeability condition in Theorem 2.4, one obtains partition functions, $\mathcal{Z}$, of the form:
  1. 1.
    $\mathcal{Z}=\mathcal{G}_{\mu}$ for some $\mu\in\mathbb{R}$, where $\mathcal{G}_{\mu}$ is defined as in (1.3).
  2. 2.
    $\mathcal{Z}=\mathcal{Z}_{\alpha,\beta}$, for $\alpha<1-\kappa/8$ and $\beta\in[0,1]$, where
    $\mathcal{Z}_{\alpha,\beta}(\theta_{1},\theta_{2})=(\sin(\theta_{21}/2))^{-2h}(\beta\Phi^{L}((\sin(\theta_{21}/4))^{2})+(1-\beta)\Phi^{R}((\sin(\theta_{21}/4))^{2})).$
The locally commuting 2-radial $\operatorname{SLE}_{\kappa}$ corresponding to the second case above is obtained analogously to Corollary 3.14. That is, one weights the law of a chordal $\operatorname{SLE}_{\kappa}$ in $(\mathbb{D};\mathrm{e}^{\mathfrak{i}\theta_{1}},\mathrm{e}^{\mathfrak{i}\theta_{2}})$, denoted by $\gamma$, by
[…]
then lets $\eta^{(1)}$ be $\gamma$ and $\eta^{(2)}$ be the time-reversal of $\gamma$, and finally restricts the law of $(\eta^{(1)},\eta^{(2)})$ to disjoint neighborhoods. One sees that the obtained family of local laws satisfies (CI), (DMP), and (MARG) by using the reversibility of SLE: this makes $\eta^{(2)}$ a chordal SLE in $(\mathbb{D};\mathrm{e}^{\mathfrak{i}\theta_{2}},\mathrm{e}^{\mathfrak{i}\theta_{1}})$ weighted by
[…]
Proof of Proposition 3.16.
If we in the proof of Theorem 2.4 assume that $\mu=0$ without assuming interchangeability, then $\mathcal{Z}$ is a positive solution of (2.25). By changing variables by (2.27), we find that $\mathcal{Z}$ corresponds to a positive solution $\phi$ of (3.24). Lemma A.2, shows that (up to a multiplicative constant)
$\phi=\beta\Phi^{L}+(1-\beta)\Phi^{R},\quad\beta\in[0,1],$ (3.33)
if $\alpha<1-\kappa/8$, that there are no positive solutions if $\alpha>1-\kappa/8$, and that there is only one positive solution, corresponding to $\mathcal{Z}=\mathcal{G}_{0}$, if $\alpha=1-\kappa/8$. ∎
4 Semiclassical limits of commutation relation
4.1 Commutation relation when $\kappa=0$
We now consider the commutation relation when $\kappa=0$. In the multichordal SLE case, a similar semiclassical limit of partition functions was considered in [34] and [3].
It is not hard to see that the infinitesimal commutation relation (Proposition 2.2) also holds when $\kappa=0$. However, we need to derive the BPZ equation and classify the partition functions differently, which we summarize in the following proposition.
Proposition 4.1.
We consider an interchangeable and locally commuting $2$-radial $\operatorname{SLE}_{0}$. Let $b_{1},b_{2}:S^{1}\times S^{1}\setminus\Delta\to\mathbb{R}$ be $C^{2}$ functions as in the condition (MARG). Then (2.4) and (2.6) imply that there exists $\mathcal{U}:\{(\theta_{1},\theta_{2})\in\mathbb{R}^{2}\,|\,\theta_{1}<\theta_{2}<\theta_{1}+2\pi\}\to\mathbb{R}$ and a constant $F$ such that
$b_{j}=\partial_{j}\mathcal{U},\quad j=1,2$
and
$\begin{split}(\partial_{2}\mathcal{U})^{2}+2\cot(\theta_{12}/2)\partial_{1}\mathcal{U}-\frac{3}{\left(\sin(\theta_{12}/2)\right)^{2}}&=F,\\ (\partial_{1}\mathcal{U})^{2}+2\cot(\theta_{21}/2)\partial_{2}\mathcal{U}-\frac{3}{\left(\sin(\theta_{21}/2)\right)^{2}}&=F,\end{split}$ (4.1)
where $\theta_{21}=\theta_{2}-\theta_{1}=-\theta_{12}$. The only solutions are
$\mathcal{U}(\theta_{1},\theta_{2})=\mathcal{U}_{\mu}(\theta_{1},\theta_{2}):=2\log\sin(\theta_{21}/2)+\mu(\theta_{1}+\theta_{2})$ (4.2)
for some $\mu\in\mathbb{R}$ or
$\mathcal{U}(\theta_{1},\theta_{2})=-6\log\sin(\theta_{21}/2)$ (4.3)
up to an additive constant.
Remark 4.2.
It follows immediately from the proof that, if one removes the interchangeability assumption from Proposition 4.1, then the only additional solutions $\mathcal{U}$ are of the from
$\mathcal{U}(\theta_{1},\theta_{2})=\mathcal{U}_{\lambda,+}(\theta_{1},\theta_{2}):=-2\log\sin(\theta_{21}/2)+\int_{\theta_{21}}^{\pi}\sqrt{2\lambda+4\cot^{2}(u/2)}\mathrm{d}u,$
$\mathcal{U}(\theta_{1},\theta_{2})=\mathcal{U}_{\lambda,-}(\theta_{1},\theta_{2}):=-2\log\sin(\theta_{21}/2)-\int_{\theta_{21}}^{\pi}\sqrt{2\lambda+4\cot^{2}(u/2)}\mathrm{d}u,$
for $\lambda>0$, up to an additive constant.
Proof.
The proof of Proposition 2.3 up to (2.11) holds verbatim when $\kappa=0$. By conformal invariance $b_{j}$ is invariant under rotation $(\theta_{1},\theta_{2})\mapsto(\theta_{1}+a,\theta_{2}+a)$. Thus, $\hat{b}_{j}:S^{1}\setminus\{0\}\to\mathbb{R}$ defined by $\hat{b}_{j}(\theta)=b_{j}(\theta,0)$, $j=1,2$, satisfies
$\partial_{1}b_{j}(\theta_{1},\theta_{2})=\partial_{1}\hat{b}_{j}^{\prime}(\theta_{12})=-\partial_{2}b_{j}(\theta_{1},\theta_{2}).$
Equations (2.10) and (2.11) give:
$(\cot(\theta/2)-\hat{b}_{2}(\theta))\hat{b}_{1}^{\prime}(\theta)-\frac{\hat{b}_{1}(\theta)}{2\sin^{2}(\theta/2)}+\frac{3\cot(\theta/2)}{2\sin^{2}(\theta/2)}=0$ (4.4)
$(\cot(\theta/2)+\hat{b}_{1}(\theta))\hat{b}_{2}^{\prime}(\theta)-\frac{\hat{b}_{2}(\theta)}{2\sin^{2}(\theta/2)}-\frac{3\cot(\theta/2)}{2\sin^{2}(\theta/2)}=0.$ (4.5)
Taking the difference of these equations yields
$\frac{\mathrm{d}}{\mathrm{d}\theta}((\cot(\theta/2)+\hat{b}_{1}(\theta))(\cot(\theta/2)-\hat{b}_{2}(\theta))+\frac{4\cot(\theta/2)}{\sin^{2}(\theta/2)}=0,$
so that
$(\cot(\theta/2)+\hat{b}_{1}(\theta))(\cot(\theta/2)-\hat{b}_{2}(\theta))=\frac{4}{\sin^{2}(\theta/2)}+C,$ (4.6)
for some $C\in\mathbb{R}$. Thus, multiplying (4.4) by $2\sin^{2}(\theta/2)(\hat{b}_{1}(\theta)+\cot(\theta/2))$ yields
$0=$ $(8+2C\sin^{2}(\theta/2))\hat{b}_{1}^{\prime}(\theta)-(\hat{b}_{1}(\theta)+\cot(\theta/2))(\hat{b}_{1}(\theta)-3\cot(\theta/2))$
$=$ $(8+2C\sin^{2}(\theta/2))f^{\prime}(\theta)-((f(\theta))^{2}+C+4)$
where $f$ is defined by $f(\theta)=\hat{b}_{1}(\theta)-\cot(\theta/2)$. Thus, $f$ solves the separable differential equation
$(8+2C\sin^{2}(\theta/2))f^{\prime}(\theta)=(f(\theta))^{2}+C+4.$ (4.7)
Similarly, we see using (4.5) that $g$ defined by $g(\theta)=\hat{b}_{2}(2\pi-\theta)-\cot(\theta/2)$ also solves (4.7) and by (4.6)
$(2\cot(\theta/2)+f(\theta))(2\cot(\theta/2)-g(2\pi-\theta))=\frac{4}{\sin^{2}(\theta/2)}+C.$ (4.8)
The solutions of (4.7) defined on $(0,2\pi)$ are studied in Lemma B.1.
Case 1: Suppose $C+4>0$. Then both $f$ and $g$ are of the form (B.2) for some constants $D=D_{f},D_{g}\in[-\pi/4,\pi/4].$ From (4.8) it follows that
$(C+4)=f(\pi)g(\pi)=(C+4)\tan(D_{f})\tan(D_{g})$
and thus we must have $D_{g}=D_{f}\in\{\pi/4,-\pi/4\}.$ Using the half-angle formula for the tangent function one can see that these two solutions can be rewritten as
$\begin{cases}f(\theta)=\sqrt{(C+4)+4\cot^{2}(\theta/2)}-\cot(\theta/2)\\ g(\theta)=-\sqrt{(C+4)+4\cot^{2}(\theta/2)}-\cot(\theta/2)\end{cases}$
$\begin{cases}f(\theta)=-\sqrt{(C+4)+4\cot^{2}(\theta/2)}-\cot(\theta/2)\\ g(\theta)=\sqrt{(C+4)+4\cot^{2}(\theta/2)}-\cot(\theta/2).\end{cases}$
Case 2: If $C+4=0$ we have only the solutions
$\begin{cases}f(\theta)=0\\ g(\theta)=0\end{cases}\qquad\text{and}\qquad\begin{cases}f(\theta)=-4\cot(\theta/2)\\ g(\theta)=-4\cot(\theta/2).\end{cases}$
Indeed, if
[…]
for some $D\leq 0$ and $E\geq 0$ then solving (4.8) for $g$ yields
[…]
Therefore, we must have $D=E=0$ (otherwise $g$ is not well-defined on all of $(0,2\pi)$).
Case 3: For $C+4<0$ we have the solutions
$\begin{cases}f(\theta)=\sqrt{-C-4}\\ g(\theta)=\sqrt{-C-4}\end{cases}\quad\text{and}\quad\begin{cases}f(\theta)=-\sqrt{-C-4}\\ g(\theta)=-\sqrt{-C-4}.\end{cases}$
Here we have again used that $f(\pi)g(\pi)=-(C+4).$
From the solutions $(f,g)$ above we obtain corresponding solutions $(\hat{b}_{1},\hat{b}_{2})$ of (4.4) and (4.5) and $(b_{1},b_{2})$ of (2.10) and  (2.11) for $\kappa=0$. We note that these solutions satisfy $\partial_{1}b_{2}=\partial_{2}b_{1}$ (or equivalently $\hat{b}_{1}^{\prime}=-\hat{b}_{2}^{\prime}$, or $f^{\prime}(\theta)=g^{\prime}(2\pi-\theta)$) so there exists a function
$\mathcal{U}:\{(\theta_{1},\theta_{2})\,|\,\theta_{1}<\theta_{2}<\theta_{1}+2\pi\}\to\mathbb{R}$
such that $b_{1}=\partial_{1}\mathcal{U}$ and $b_{2}=\partial_{2}\mathcal{U}$. Up to an additive constant we have that
$\mathcal{U}(\theta_{1},\theta_{2})=\mathcal{U}_{\lambda,+}(\theta_{1},\theta_{2}):=-2\log\sin(\theta_{21}/2)+\int_{\theta_{21}}^{\pi}\sqrt{2\lambda+4\cot^{2}(u/2)}\mathrm{d}u,$
$\mathcal{U}(\theta_{1},\theta_{2})=\mathcal{U}_{\lambda,-}(\theta_{1},\theta_{2}):=-2\log\sin(\theta_{21}/2)-\int_{\theta_{21}}^{\pi}\sqrt{2\lambda+4\cot^{2}(u/2)}\mathrm{d}u,$
for $2\lambda=C+4>0$,
$\mathcal{U}(\theta_{1},\theta_{2})=-6\log\sin(\theta_{21}/2),$
and
$\mathcal{U}(\theta_{1},\theta_{2})=\mathcal{U}_{\mu}(\theta_{1},\theta_{2})=2\log\sin(\theta_{21}/2)+\mu(\theta_{1}+\theta_{2})$
for $\mu=\pm\sqrt{-C-4}\in\mathbb{R}$. It is easily verified that each of these $\mathcal{U}$ satisfy (4.1) for some constant $F$. We note that $\mathcal{U}_{\lambda,\pm}$ does not satisfy the interchangeability condition but that the others do. ∎
We have defined the two-sided radial $\operatorname{SLE}_{\kappa}$ with spiraling rate $\mu$ for $\kappa\in(0,8)$ in Definition 3.6, we now extend the definition to the case $\kappa=0$.
Definition 4.3.
We use the same notations as in Figure 5, we say that a deterministic pair of continuous simple curves $(\eta^{(1)},\eta^{(2)})$ is the two-sided radial $\operatorname{SLE}_{0}$ with spiraling rate $\mu$ in $(\mathbb{D};\mathrm{e}^{\mathfrak{i}\theta_{1}},\mathrm{e}^{\mathfrak{i}\theta_{2}};0)$ if we have for all $\mathbf{t}=(t_{1},t_{2})$,
$\begin{cases}\theta_{(0,0)}^{(1)}=\theta_{1},\quad\theta_{(0,0)}^{(2)}=\theta_{2},\\ \mathrm{d}\theta_{\mathbf{t}}^{(1)}=\mu\phi_{\mathbf{t},1}^{\prime}\left(\xi_{t_{1}}^{(1)}\right)^{2}\mathrm{d}t_{1}+\cot\left((\theta_{\mathbf{t}}^{(1)}-\theta_{\mathbf{t}}^{(2)})/2\right)\left(\phi_{\mathbf{t},1}^{\prime}\left(\xi_{t_{1}}^{(1)}\right)^{2}\mathrm{d}t_{1}+\phi_{\mathbf{t},2}^{\prime}\left(\xi_{t_{2}}^{(2)}\right)^{2}\mathrm{d}t_{2}\right),\\ \mathrm{d}\theta_{\mathbf{t}}^{(2)}=\mu\phi_{\mathbf{t},2}^{\prime}\left(\xi_{t_{2}}^{(2)}\right)^{2}\mathrm{d}t_{2}+\cot\left((\theta_{\mathbf{t}}^{(2)}-\theta_{\mathbf{t}}^{(1)})/2\right)\left(\phi_{\mathbf{t},1}^{\prime}\left(\xi_{t_{1}}^{(1)}\right)^{2}\mathrm{d}t_{1}+\phi_{\mathbf{t},2}^{\prime}\left(\xi_{t_{2}}^{(2)}\right)^{2}\mathrm{d}t_{2}\right).\end{cases}$
To justify such a pair exists, we note that when $t_{2}\equiv 0$, we write $\theta_{\mathbf{t}}^{(1)}=\xi_{t_{1}}^{(1)}$ and $\theta_{\mathbf{t}}^{(2)}=V_{t_{1}}^{(2)}$, then
$\begin{cases}\xi_{0}^{(1)}=\theta_{1},\quad V_{0}^{(2)}=\theta_{2},\\ \mathrm{d}\xi_{t_{1}}^{(1)}=\mu\mathrm{d}t_{1}+\cot\left((\xi_{t_{1}}^{(1)}-V_{t_{1}}^{(2)})/2\right)\mathrm{d}t_{1}=\partial_{1}\mathcal{U}_{\mu}(\xi_{t_{1}}^{(1)},V_{t_{1}}^{(2)})\mathrm{d}t_{1},\\ \mathrm{d}V_{t_{1}}^{(2)}=\cot\left((V_{t_{1}}^{(2)}-\xi_{t_{1}}^{(1)})/2\right)\mathrm{d}t_{1},\end{cases}$ (4.9)
[…]
Since $\mathcal{U}_{\mu}$ satisfies the commutation relation as we showed in Proposition 4.1, we know that $(\eta^{(1)},\eta^{(2)})$ gives a pair in Definition 4.3 if we show that $\theta_{\mathbf{t}}^{(2)}\notin\{\theta_{\mathbf{t}}^{(1)},\theta_{\mathbf{t}}^{(1)}+2\pi\}$ for all $\mathbf{t}$. Indeed, the map $t_{2}\mapsto\theta_{(0,t_{2})}^{(2)}-\theta_{(0,t_{2})}^{(1)}$ is repulsive away from $\{0,2\pi\}$, so is well-defined and for all $t_{2}\geq 0$. Similarly, for fixed $t_{2}$, the function $t_{1}\mapsto\theta_{(t_{1},t_{2})}^{(2)}-\theta_{(t_{1},t_{2})}^{(1)}$ is repulsive away from $\{0,2\pi\}$, so $\theta_{\mathbf{t}}^{(2)}\notin\{\theta_{\mathbf{t}}^{(1)},\theta_{\mathbf{t}}^{(1)}+2\pi\}$ for all $t_{1},t_{2}\geq 0$.
Corollary 4.4.
The only interchangeable and locally commuting $2$-radial $\operatorname{SLE}_{0}$ are the two-sided radial $\operatorname{SLE}_{0}$ with spiraling rate $\mu\in\mathbb{R}$ and the chordal $\operatorname{SLE}_{0}$.
Proof.
When $\mathcal{U}=\mathcal{U}_{\mu}$, the corresponding commuting $\operatorname{SLE}_{0}$ is the two-sided radial $\operatorname{SLE}_{0}$ (with spiraling rate $\mu$).
When $\mathcal{U}(\theta_{1},\theta_{2})=-6\log\sin(\theta_{21}/2)$, the corresponding commuting $\operatorname{SLE}_{0}$ is the chordal $\operatorname{SLE}_{0}$ in $(\mathbb{D};\mathrm{e}^{\mathfrak{i}\theta_{1}},\mathrm{e}^{\mathfrak{i}\theta_{2}})$ and its time-reversal. ∎
Remark 4.5.
We note that when $\mathcal{U}(\theta_{1},\theta_{2})=\mathcal{U}_{0}(\theta_{1},\theta_{2})=2\log\sin(\theta_{21}/2)$, the corresponding commuting $\operatorname{SLE}_{0}$ is the two-sided radial $\operatorname{SLE}_{0}$ (with zero spiraling rate). This is also the geodesic pair in $\mathfrak{X}(\mathbb{D};\mathrm{e}^{\mathfrak{i}\theta_{1}},\mathrm{e}^{\mathfrak{i}\theta_{2}};0)$ studied in [43, 29, 30], or in other words, the concatenation of $\eta^{(1)}$ and $\eta^{(2)}$ is the chord minimizing the chordal Loewner energy among all chords in $(\mathbb{D};\mathrm{e}^{\mathfrak{i}\theta_{1}},\mathrm{e}^{\mathfrak{i}\theta_{2}})$ passing through 0, hence can be viewed as the chordal $\operatorname{SLE}_{0}$ “conditioned” to pass through 0.
We remark that, unlike the $\kappa>0$ case, we do not have the second one-parameter family given by weighting by conformal radius as described in Section 3.4.
4.2 Semiclassical limits of partition functions
To understand the reason why the second one-parameter family given by $\mathcal{Z}_{\alpha}$ disappears when $\kappa=0$, we now take a closer look at the semiclassical limit ($\kappa\to 0$) of $\mathcal{G}_{\mu}$ and $\mathcal{Z}_{\alpha}$.
For $\mu\in\mathbb{R}$, recall that $\mathcal{G}_{\mu}$ is defined in (1.3):
$\mathcal{G}_{\mu}(\theta_{1},\theta_{2})=\left(\sin\left((\theta_{2}-\theta_{1})/2\right)\right)^{2/\kappa}\exp\left(\frac{\mu}{\kappa}(\theta_{1}+\theta_{2})\right).$
Lemma 4.6.
Fix $\mu\in\mathbb{R}$ and $0<\theta_{1}<\theta_{2}<\theta_{1}+2\pi$, we have (1.6):
$\lim_{\kappa\to 0}\kappa\log\mathcal{G}_{\mu}(\theta_{1},\theta_{2})=2\log\sin\left((\theta_{2}-\theta_{1})/2\right)+\mu(\theta_{1}+\theta_{2})$
which is the solution (4.2) in Proposition 4.1.
Proof.
This is immediate from the expression of $\mathcal{G}_{\mu}$. ∎
Lemma 4.6 gives the first part of Proposition 1.3. We will prove its second part below. For $\alpha<1-\kappa/8$, recall that $\mathcal{Z}_{\alpha}$ is defined in (1.4):
$\mathcal{Z}_{\alpha}(\theta_{1},\theta_{2})=\left(\sin\left((\theta_{2}-\theta_{1})/2\right)\right)^{(\kappa-6)/\kappa}\phi_{\alpha}\left(\left(\sin\left((\theta_{2}-\theta_{1})/4\right)\right)^{2}\right),$
where $\phi_{\alpha}$ is the unique solution to (1.5). Before we derive semiclassical limit of $\mathcal{Z}_{\alpha}$, let us first address chordal Loewner energy [34]. Recall that $\mathfrak{X}(\mathbb{D};\mathrm{e}^{\mathfrak{i}\theta_{1}},\mathrm{e}^{\mathfrak{i}\theta_{2}})$ is the space of all continuous curves in $\mathbb{D}$ connecting $\mathrm{e}^{\mathfrak{i}\theta_{1}}$ and $\mathrm{e}^{\mathfrak{i}\theta_{2}}$.
Lemma 4.7.
Fix $\theta_{1}<\theta_{2}<\theta_{1}+2\pi$. Suppose $\gamma\in\mathfrak{X}(\mathbb{D};\mathrm{e}^{\mathfrak{i}\theta_{1}},\mathrm{e}^{\mathfrak{i}\theta_{2}})$. We parameterize it by the capacity and define $g_{t},\xi_{t}$ accordingly as in Section 3.1 and set $T=-\log\mathrm{CR}(\mathbb{D}\setminus\gamma)$. Suppose $t\mapsto\xi_{t}$ is absolutely continuous and denote its derivative by $\dot{\xi}_{t}$. We denote by $t\mapsto V_{t}$ the solution to
$\dot{V}_{t}=\cot\left((V_{t}-\xi_{t})/2\right),\quad V_{0}=\theta_{2}.$ (4.10)
Then the chordal Loewner energy $I(\gamma)$ can be written as
$I(\gamma)=\frac{1}{2}\int_{0}^{T}\left(\dot{\xi}_{s}-3\cot\left((V_{s}-\xi_{s})/2\right)\right)^{2}\mathrm{d}s.$ (4.11)
Furthermore, the infimum of $I(\gamma)$ in $\mathfrak{X}(\mathbb{D};\mathrm{e}^{\mathfrak{i}\theta_{1}},\mathrm{e}^{\mathfrak{i}\theta_{2}})$ is zero and is attained by the hyperbolic geodesic.
Proof.
Eq. (4.11) is a standard calculation by changing coordinates. ∎
Lemma 4.8.
If we choose $\alpha=\alpha(\kappa)$ such that $\alpha=o(1/\kappa)$, then
$\lim_{\kappa\to 0}\kappa\log\mathcal{Z}_{\alpha}(\theta_{1},\theta_{2})=-6\log\sin(\theta_{21}/2)$ (4.12)
which is the solution (4.3) in Proposition 4.1.
If $\alpha\sim-\lambda/\kappa$ for some $\lambda>0$, then the limit $\lim_{\kappa\to 0}\kappa\log\mathcal{Z}_{\alpha}(\theta_{1},\theta_{2})$ exists and equals
$\begin{split}\mathcal{U}^{\lambda}(\theta_{1},\theta_{2})&:=-6\log\sin(\theta_{21}/2)-\inf_{\gamma\in\mathfrak{X}(\mathbb{D};\mathrm{e}^{\mathfrak{i}\theta_{1}},\mathrm{e}^{\mathfrak{i}\theta_{2}})}\left(I(\gamma)-\lambda\log\mathrm{CR}(\mathbb{D}\setminus\gamma)\right)\\ &\quad+\inf_{\gamma\in\mathfrak{X}(\mathbb{D};-1,1)}\left(I(\gamma)-\lambda\log\mathrm{CR}(\mathbb{D}\setminus\gamma)\right),\end{split}$ (4.13)
where $I(\gamma)$ is the chordal Loewner energy of $\gamma$ in $\mathfrak{X}(\mathbb{D};\mathrm{e}^{\mathfrak{i}\theta_{1}},\mathrm{e}^{\mathfrak{i}\theta_{2}})$ and the infimums are attained.
We note that the constraint $\alpha<1-\kappa/8$ implies that we can only choose $\lambda\geq 0$ and the last term in (4.13) is a constant such that $\mathcal{U}^{\lambda}(\theta_{1},\theta_{1}+\pi)=0$ as we have normalized $\mathcal{Z}_{\alpha}$ such that $\mathcal{Z}_{\alpha}(\theta_{1},\theta_{1}+\pi)=1$.
Proof.
For $\alpha<1-\kappa/8$, recall from Lemma 3.12 and (3.29) that $\mathcal{Z}_{\alpha}$ is defined as
$\mathcal{Z}_{\alpha}(\theta_{1},\theta_{2})=\left(\sin(\theta/2)\right)^{(\kappa-6)/\kappa}\frac{\mathbb{E}_{\theta}[\mathrm{CR}(\mathbb{D}\setminus\gamma)^{-\alpha}]}{\mathbb{E}_{\pi}[\mathrm{CR}(\mathbb{D}\setminus\gamma)^{-\alpha}]},$
where $\mathbb{E}_{\theta}$ is the expectation with respect to the law of $\gamma$ which is a chordal $\operatorname{SLE}_{\kappa}$ in $(\mathbb{D};\mathrm{e}^{\mathfrak{i}\theta_{1}},\mathrm{e}^{\mathfrak{i}\theta_{2}})$ and $\theta=\theta_{2}-\theta_{1}$.
Then the result follows from the large deviation principle for chordal SLE as $\kappa\to 0{\scriptstyle+}$ [34]. In fact, the Loewner energy is the large deviation rate function of chordal $\operatorname{SLE}_{0{\scriptstyle+}}$ for the Hausdorff metric and $\gamma\mapsto-\log\mathrm{CR}(\mathbb{D}\setminus\gamma)$ is a continuous function $\mathfrak{X}(\mathbb{D};\mathrm{e}^{\mathfrak{i}\theta_{1}},\mathrm{e}^{\mathfrak{i}\theta_{2}})\to[0,\infty]$. Varadhan’s lemma [11, Lem. 4.3.4 and 4.3.6] shows if $\alpha\sim-\lambda/\kappa$, then
$\lim_{\kappa\to 0}\kappa\log\mathbb{E}_{\theta}[\mathrm{CR}(\mathbb{D}\setminus\gamma)^{-\alpha}]=-\inf_{\gamma\in\mathfrak{X}(\mathbb{D};\mathrm{e}^{\mathfrak{i}\theta_{1}},\mathrm{e}^{\mathfrak{i}\theta_{2}})}\left(I(\gamma)-\lambda\log\mathrm{CR}(\mathbb{D}\setminus\gamma)\right)$
which proves the limit (4.13). Since the large deviation rate function $I$ of chordal SLE0+ is good, the infimum in (4.13) is attained.
Similarly, an easy bound and Varadhan’s lemma also show the limit (4.12). ∎
Proposition 4.9.
For $\theta_{1}<\theta_{2}\leq\theta_{1}+\pi$, we denote $\theta=\theta_{2}-\theta_{1}$ and we have
$\inf_{\gamma\in\mathfrak{X}(\mathbb{D};\mathrm{e}^{\mathfrak{i}\theta_{1}},\mathrm{e}^{\mathfrak{i}\theta_{2}})}\left(I(\gamma)-\lambda\log\mathrm{CR}(\mathbb{D}\setminus\gamma)\right)=\int_{0}^{\theta}\Big(\sqrt{2\lambda+4\cot^{2}(u/2)}-2\cot(u/2)\Big)\,\mathrm{d}u.$ (4.14)
  • If $\theta_{1}<\theta_{2}<\theta_{1}+\pi$, the infimum in (4.14) is attained for a unique curve $\gamma^{*}\in\mathfrak{X}(\mathbb{D};\mathrm{e}^{\mathfrak{i}\theta_{1}},\mathrm{e}^{\mathfrak{i}\theta_{2}})$ whose radial driving function $\xi^{*}_{\cdot}$, defined on $[0,T]$, satisfies
    $\begin{cases}\xi_{0}^{*}=\theta_{1},V_{0}^{*}=\theta_{2},\\ \dot{\xi}^{*}_{t}=\cot((V^{*}_{t}-\xi^{*}_{t})/2)+\sqrt{2\lambda+4\cot^{2}((V^{*}_{t}-\xi^{*}_{t})/2)},\\ \dot{V}^{*}_{t}=\cot((V^{*}_{t}-\xi^{*}_{t})/2),\end{cases}$ (4.15)
    and $\lim_{t\to T-}(V_{t}^{*}-\xi_{t}^{*})=0$.
  • If instead $\theta_{2}=\theta_{1}+\pi,$ then the infimum in (4.14) is attained for two curves, $\gamma^{*}$ and $\gamma^{**}$. One of the corresponding driving functions, say $\xi^{*}_{\cdot}$, satisfies (4.15) so that $\lim_{t\to T-}(V_{t}^{*}-\xi_{t}^{*})=0$, while the other $\xi^{**}_{\cdot},$ satisfies $\dot{\xi}^{**}_{t}=-\dot{\xi}^{*}_{t}$ for all $t\in[0,T)$, so that $\lim_{t\to T-}(V_{t}^{**}-\xi_{t}^{**})=2\pi$.
Proof.
We denote the right-hand side of (4.14) by
$H_{\lambda}(\theta)=\int_{0}^{\theta}\Big(\sqrt{2\lambda+4\cot^{2}(u/2)}-2\cot(u/2)\Big)\mathrm{d}u,$ (4.16)
for $\theta\in[0,\pi]$, and $H_{\lambda}(\theta)=H_{\lambda}(2\pi-\theta)$ for $\theta\in(\pi,2\pi]$. Let $\gamma_{[0,t]}:[0,t]\to\mathbb{D}\setminus\{0\}$ be a simple curve, parametrized by capacity, with an absolutely continuous driving function $t\mapsto\xi_{t}$ and set $V_{t}$ as in (4.10). Define
$J^{\lambda}_{(\theta_{1},\theta_{2})}\left(\gamma_{[0,t]}\right):=\frac{1}{2}\int_{0}^{t}\bigg(\dot{\xi}_{s}-3\cot((V_{s}-\xi_{s})/2)-H^{\prime}_{\lambda}(V_{s}-\xi_{s})\bigg)^{2}\mathrm{d}s.$ (4.17)
Since $H_{\lambda}$ is not differentiable at $\theta=\pi$ (the left and right derivatives differ by a sign), the integrand on the right-hand side of Eq. (4.17) is not necessarily well-defined for $s$ such that $V_{s}-\xi_{s}=\pi$. However, this is not an issue: Let $E=\{s\in[0,t]:V_{s}-\xi_{s}=\pi\}$. In the interior of $E$, we have $\dot{\xi}_{s}=0$ and $\dot{V}_{s}=0$. Hence, the integrand is well-defined and takes the value $2\lambda$ for $s\in E^{\circ}$. Since $\partial E$ has measure zero the integrand need not be well-defined there.
Let us connect $J^{\lambda}_{(\theta_{1},\theta_{2})}$ to the chordal Loewner energy $I$ in Lemma 4.7. We have the following two observations.
  • Denote by $\hat{\gamma}_{t}$ the union of $\gamma_{[0,t]}$ and the hyperbolic geodesic from $\gamma_{t}$ to $\mathrm{e}^{\mathfrak{i}\theta_{2}}$ in $\mathbb{D}\setminus\gamma_{[0,t]}$. Then Lemma 4.7 gives the energy of $\hat{\gamma}_{t}$:
    $I(\hat{\gamma}_{t})=\frac{1}{2}\int_{0}^{t}\bigg(\dot{\xi}_{s}-3\cot((V_{s}-\xi_{s})/2)\bigg)^{2}\mathrm{d}s.$
  • For $H_{\lambda}$, we have
    $H_{\lambda}(V_{t}-\xi_{t})-H_{\lambda}(\theta_{2}-\theta_{1})=\int_{0}^{t}H^{\prime}_{\lambda}(V_{s}-\xi_{s})\Big(\cot((V_{s}-\xi_{s})/2)-\dot{\xi}_{s}\Big)\mathrm{d}s.$
Plugging these two observations into (4.17), we have
$J^{\lambda}_{(\theta_{1},\theta_{2})}\left(\gamma_{[0,t]}\right)=I(\hat{\gamma}_{t})-\lambda\log\mathrm{CR}(\mathbb{D}\setminus\gamma_{t})+H_{\lambda}(V_{t}-\xi_{t})-H_{\lambda}(\theta_{2}-\theta_{1}).$ (4.18)
Suppose $\gamma:(0,T)\to\mathbb{D}\setminus\{0\},$ with $\gamma(0+)=\mathrm{e}^{\mathfrak{i}\theta_{1}}$ and $\gamma(T-)=\mathrm{e}^{\mathfrak{i}\theta_{2}},$ has finite chordal Loewner energy. Then the associated radial driving function, in the capacity parametrization, $\xi_{\cdot}$, is absolutely continuous on $[0,t]$ for all $t\in[0,T)$. Furthermore, a harmonic measure argument shows that
$V_{t}-\xi_{t}\to 0\quad\text{or}\quad V_{t}-\xi_{t}\to 2\pi,\quad\text{as }t\to T$-
(depending on, on which side of the origin $\gamma$ passes). Hence, Eq. (4.18) implies
$I(\gamma)-\lambda\log\mathrm{CR}(\mathbb{D}\setminus\gamma)=H_{\lambda}(\theta_{2}-\theta_{1})+\lim_{t\to T-}J^{\lambda}_{(\theta_{1},\theta_{2})}(\gamma_{t})\geq H_{\lambda}(\theta_{2}-\theta_{1}).$ (4.19)
Now let $\theta_{1}<\theta_{2}\leq\theta_{1}+\pi$, and let $(\xi^{*}_{t},V^{*}_{t})$ be the solution to (4.15). Then,
$\partial_{t}(V^{*}_{t}-\xi^{*}_{t})=-\sqrt{2\lambda+4\cot^{2}((V^{*}_{t}-\xi^{*}_{t})/2)},$
so that there does, indeed, exist $T\in(0,\infty)$ so that $\lim_{t\to T-}(V_{t}^{*}-\xi_{t}^{*})=0$. It follows from (4.19) that $\xi^{*}_{\cdot}$ is the driving function of a simple curve $\gamma^{*}$, with $\gamma^{*}(0+)=\mathrm{e}^{\mathfrak{i}\theta_{1}}$ and $\gamma^{*}(T-)=\mathrm{e}^{\mathfrak{i}\theta_{2}}$, and that
$I(\gamma^{*})-\lambda\log\mathrm{CR}(\mathbb{D}\setminus\gamma^{*})=H_{\lambda}(\theta_{2}-\theta_{1}).$
Thus, the infimum of $I(\gamma)-\lambda\log\mathrm{CR}(\mathbb{D}\setminus\gamma)$ equals $H_{\lambda}(\theta_{2}-\theta_{1})$ and is attained by $\gamma^{*}$. Furthermore, any $\tilde{\gamma}\in\mathfrak{X}(\mathbb{D};\mathrm{e}^{\mathfrak{i}\theta_{1}},\mathrm{e}^{\mathfrak{i}\theta_{2}})$ which minimizes $I(\gamma)-\lambda\log\mathrm{CR}(\mathbb{D}\setminus\gamma)$ must satisfy $J^{\lambda}_{(\theta_{1},\theta_{2})}\left(\tilde{\gamma}_{[0,t]}\right)=0$ for all $t$, or equivalently, its driving function must satisfy
$\dot{\xi}_{t}=3\cot((V_{t}-\xi_{t})/2))+H_{\lambda}^{\prime}(V_{t}-\xi_{t}),\quad\text{a.e.}$ (4.20)
If $\theta_{2}<\theta_{1}+\pi$ the unique continuous solution of (4.20) is $\xi_{\cdot}^{*}$. If $\theta_{2}=\theta_{1}+\pi$, then (4.20) has two continuous solutions, $\xi_{\cdot}^{*}$ and $\xi_{\cdot}^{**}$, where $\dot{\xi}_{t}^{**}=-\dot{\xi}_{t}^{*}$. ∎
Since $\mathcal{Z}_{\alpha}$ is a function of $\theta:=\theta_{2}-\theta_{1}$, so we may write $\mathcal{U}^{\lambda}(\theta)=\mathcal{U}^{\lambda}(\theta_{1},\theta_{2})$. The next corollary explains why we do not find this solution in Proposition 4.1.
Corollary 4.10.
We let $\theta=\theta_{2}-\theta_{1}$ and write $\mathcal{U}^{\lambda}(\theta)=\mathcal{U}^{\lambda}(\theta_{1},\theta_{2})$. Then we have (1.7): for $\theta\in(0,\pi)$,
$\mathcal{U}^{\lambda}(\theta)=\mathcal{U}^{\lambda}(2\pi-\theta)=-2\log\sin(\theta/2)+\int_{\theta}^{\pi}\sqrt{2\lambda+4\cot^{2}(u/2)}\,\mathrm{d}u.$
In particular, it satisfies
$(\mathcal{U}^{\prime}(\theta))^{2}+2\cot(\theta/2)\mathcal{U}^{\prime}(\theta)-\frac{3}{\left(\sin(\theta/2)\right)^{2}}=-3+2\lambda$ (4.21)
and has the left derivative $\sqrt{2\lambda}$ and right derivative $\sqrt{2\lambda}$ at $\theta=\pi$.
In other words, $\mathcal{U}^{\lambda}$ is not differentiable at $\pi$. That is why we do not see it in Proposition 4.1. Moreover, the driving function $\xi^{*}$ in (4.15) satisfies $\dot{\xi}^{*}_{t}=\partial_{1}\mathcal{U}^{\lambda}(\xi^{*}_{t},V^{*}_{t})$ which is analogous to (4.9).
Proof.
The expression (1.7) follows directly from (4.13) and Proposition 4.9. It is straightforward to check that it satisfies (4.21). ∎
Proof of Proposition 1.3 and Proposition 1.4.
This is a collection of Lemma 4.6, Lemma 4.8, Proposition 4.9 and Corollary 4.10. ∎
Remark 4.11.
When $\lambda=2$, Corollary 4.10 is consistent with Remark 2.7: when $\alpha=\alpha_{1}(\kappa)$ as in (2.29), we have $\alpha_{1}(\kappa)\sim-2/\kappa$ and thus
$\mathcal{U}^{\lambda=2}(\theta)=$ $\lim_{\kappa\to 0+}\kappa\log\mathcal{Z}_{\alpha_{1}(\kappa)}(\theta_{1},\theta_{2})$
$=$ $\begin{cases}4\log 2-6\log\sin(\theta/2)+8\log\cos(\theta/4),&\text{if }\theta\in(0,\pi];\\ 4\log 2-6\log\sin(\theta/2)+8\log\sin(\theta/4),&\text{if }\theta\in[\pi,2\pi).\end{cases}$
Then
$\partial_{\theta}\mathcal{U}^{\lambda=2}(\theta)=\begin{cases}-3\cot(\theta/2)-2\tan(\theta/4),&\text{if }\theta\in(0,\pi);\\ -3\cot(\theta/2)+2\cot(\theta/4),&\text{if }\theta\in(\pi,2\pi).\end{cases}$
Note that, when $\theta\in(0,\pi)$, the derivative of the right-hand side of (1.7) is
$-\cot(\theta/2)-\frac{2}{\sin(\theta/2)}=-3\cot(\theta/2)-2\tan(\theta/4).$
Remark 4.12.
Recall from Remark 4.2 that if one removes the assumption of interchangeability from Proposition 4.1, then one obtains, in addition to the solutions (4.2) and (4.3), the solutions $\mathcal{U}_{\lambda,\pm}$, $\lambda>0$. Note that $\mathcal{U}_{\lambda,+}(\theta_{1},\theta_{2})=\mathcal{U}^{\lambda}(\theta_{1},\theta_{2})$ for $\theta_{1}<\theta_{2}\leq\theta_{1}+\pi$ and $\mathcal{U}_{\lambda,-}(\theta_{1},\theta_{2})=\mathcal{U}^{\lambda}(\theta_{1},\theta_{2})$ for $\theta_{1}+\pi\leq\theta_{2}<\theta_{1}+2\pi$. Following Lemma 4.8, one can take the semiclassical limit of $\mathcal{Z}_{\alpha,\beta}$ from Proposition 3.16. For example, if $\lambda>0$ and $\alpha\sim-\lambda/\kappa$, we have
$\lim_{\kappa\to 0}\kappa\log\frac{\mathcal{Z}_{\alpha,1}(\theta_{1},\theta_{2})}{\mathcal{Z}_{\alpha,1}(0,\pi)}=\mathcal{U}_{\lambda,+}(\theta_{1},\theta_{2}).$
This follows from
$\inf_{\begin{subarray}{c}\gamma\in\mathfrak{X}(\mathbb{D};\mathrm{e}^{\mathfrak{i}\theta_{1}},\mathrm{e}^{\mathfrak{i}\theta_{2}})\\ 0\text{ is to the left of }\gamma\end{subarray}}\left(I(\gamma)-\lambda\log\mathrm{CR}(\mathbb{D}\setminus\gamma)\right)=\int_{0}^{\theta_{21}}\Big(\sqrt{2\lambda+4\cot^{2}(u/2)}-2\cot(u/2)\Big)\,\mathrm{d}u,$
which can be seen by examining the proof of Proposition 4.9. By a similar analysis one finds
$\lim_{\kappa\to 0}\kappa\log\frac{\mathcal{Z}_{\alpha,0}(\theta_{1},\theta_{2})}{\mathcal{Z}_{\alpha,0}(0,\pi)}=\mathcal{U}_{\lambda,-}(\theta_{1},\theta_{2}).$
Appendix A Euler’s hypergeometric differential equations
Lemma A.1.
Fix $\kappa\in(0,8)$ and $\alpha\in\mathbb{R}$, we consider Euler’s hypergeometric differential equation (1.5):
$\begin{cases}u(1-u)\phi^{\prime\prime}(u)-\dfrac{3\kappa-8}{2\kappa}(2u-1)\phi^{\prime}(u)+\dfrac{8\alpha}{\kappa}\phi(u)=0,\quad u\in(0,1);\\ \phi(1/2)=1,\quad\phi^{\prime}(1/2)=0.\end{cases}$
There is a unique solution $\phi_{\alpha}$ to (1.5) in $C^{2}(0,1)$ and the solution $\phi_{\alpha}$ is continuous in $\alpha$. Furthermore,
  • when $\alpha=\alpha_{0}:=1-\kappa/8$, we have
    $\phi_{\alpha_{0}}(u)=\left(4u(1-u)\right)^{4/\kappa-1/2};$ (A.1)
  • when $\alpha<1-\kappa/8$, we have $\phi_{\alpha}(u)>0$ for all $u\in(0,1)$;
  • when $\alpha>1-\kappa/8$, there exists $u\in(0,1)$ such that $\phi_{\alpha}(u)\leq 0$.
Proof.
By direct calculation, we see that (A.1) satisfies (1.5) when $\alpha=\alpha_{0}$. We write
$\alpha_{0}=1-\kappa/8,\quad\phi_{\alpha_{0}}(u)=\left(4u(1-u)\right)^{4/\kappa-1/2},\quad\phi(u)=\phi_{\alpha_{0}}(u)f(u).$
Then Eq. (1.5) becomes
$\begin{cases}u(1-u)f^{\prime\prime}(u)-\dfrac{\kappa+8}{2\kappa}(2u-1)f^{\prime}(u)+\dfrac{8}{\kappa}(\alpha-\alpha_{0})f(u)=0,\quad u\in(0,1);\\ f(1/2)=1,\quad f^{\prime}(1/2)=0.\end{cases}$ (A.2)
We write $y_{1}=f$ and $y_{2}=f^{\prime}$, then Eq. (A.2) becomes
$\begin{cases}y_{1}^{\prime}=\mathcal{F}_{1}(u,y_{1},y_{2}):=y_{2};\\ y_{2}^{\prime}=\mathcal{F}_{2}(u,y_{1},y_{2}):=\dfrac{8}{\kappa}\dfrac{\alpha_{0}-\alpha}{u(1-u)}y_{1}+\dfrac{\kappa+8}{2\kappa}\dfrac{(2u-1)}{u(1-u)}y_{2};\\ y_{1}(1/2)=1,\quad y_{2}(1/2)=0.\end{cases}$ (A.3)
The functional $(\mathcal{F}_{1},\mathcal{F}_{2})$ is continuous in $\Lambda=(0,1)\times\mathbb{R}\times\mathbb{R}$ and satisfies a local Lipschitz condition with respect to $(y_{1},y_{2})$ in $\Lambda$. Thus, the initial value problem (A.3) has exactly one solution, and the solution can be extended up to the boundary of $\Lambda$. This gives the unique solution $\phi_{\alpha}$ to (1.5) and the solution $\phi_{\alpha}$ is continuous in $\alpha$.
Next, let us check the positivity of the solution. There are two cases.
Case 1: $\alpha\leq\alpha_{0}$. In this case, $\mathcal{F}_{1}$ is increasing in $y_{2}$ and $\mathcal{F}_{2}$ is increasing in $y_{1}$. Thus we have a comparison principle for the unique solution to (A.3). In particular, the solution $f$ to (A.2) is decreasing in $\alpha$ as long as $\alpha\leq\alpha_{0}$. Consequently, the unique solution $\phi_{\alpha}$ to (1.5) is decreasing in $\alpha$ as long as $\alpha\leq\alpha_{0}$ and $\phi_{\alpha}(u)\geq\phi_{\alpha_{0}}(u)>0$ for all $u\in(0,1)$ when $\alpha\leq\alpha_{0}$.
Case 2: $\alpha>\alpha_{0}$. We prove by contradiction and assume $f(u)\geq 0$ for all $u\in(1/2,1)$. As $f^{\prime}(1/2)=0$ and $f^{\prime\prime}(1/2)=32(\alpha_{0}-\alpha)/\kappa<0$, there exists $\varepsilon>0$ small such that
$\delta:=f^{\prime}(1/2+\varepsilon)<0.$
The ODE in (A.2) implies that
$f^{\prime\prime}(u)\leq\left(\frac{\kappa+8}{2\kappa}\right)\frac{2u-1}{u(1-u)}f^{\prime}(u),\quad u\in(1/2,1).$
Using Grönwall’s inequality, we obtain, for $u\in(1/2+\varepsilon,1)$,
$f^{\prime}(u)\leq f^{\prime}(1/2+\varepsilon)\exp\left(\int_{1/2+\varepsilon}^{u}\left(\frac{\kappa+8}{2\kappa}\right)\frac{2s-1}{s(1-s)}\mathrm{d}s\right)=-\delta\left(\frac{u(1-u)}{1/4-\varepsilon^{2}}\right)^{-\frac{\kappa+8}{2\kappa}}.$
Since $\kappa\in(0,8)$, we have $\frac{\kappa+8}{2\kappa}>1$. The derivative $f^{\prime}$ is not integrable as $u\to 1$. This implies that $f(u)\to$-$\infty$ as $u\to 1$, which contradicts our assumption and completes the proof. ∎
Lemma A.2.
Fix $\kappa\in(0,8)$ and $\alpha\in\mathbb{R}$. Consider Euler’s hypergeometric differential equation:
$u(1-u)\phi^{\prime\prime}(u)-\frac{3\kappa-8}{2\kappa}(2u-1)\phi^{\prime}(u)+\frac{8\alpha}{\kappa}\phi^{\prime}(u)=0,\quad u\in(0,1).$ (A.4)
We have the following:
  • When $\alpha>1-\kappa/8$, there exists, for each solution $\phi$ to (A.4), $u\in(0,1)$ such that $\phi(u)\leq 0.$
  • When $\alpha=1-\kappa/8$, the only positive solution (up to a multiplicative constant) of (A.4) is
    $\phi_{\alpha_{0}}(u)=(4u(1-u))^{4/\kappa-1/2}.$
  • When $\alpha<1-\kappa/8$, the positive solutions of (A.4), are (up to a multiplicative constant) all of the form
    $\phi(u)=\beta\Phi^{L}(u)+(1-\beta)\Phi^{R}(u),\quad\beta\in[0,1],$
    with $\Phi^{L}$ and $\Phi^{R}$ as in Remark 3.15.
Proof.
We first consider $\alpha\geq 1-\kappa/8.$ Let $\phi_{\alpha}$ be as in Lemma A.1, and let $\psi_{\alpha}$ be the solution of (A.4) satisfying $\psi_{\alpha}(1/2)=0$ and $\psi_{\alpha}^{\prime}(1/2)=1$. It is easily verified that
$\psi_{\alpha}(u)=-\psi_{\alpha}(1-u),\quad u\in(0,1).$ (A.5)
Since $\phi_{\alpha}$ and $\psi_{\alpha}$ are linearly independent, any solution of (A.4) can be expressed as a linear combination of them. When $\alpha>1-\kappa/8$, it follows from Lemma A.1 and Eq. (A.5) that there is no positive solution of (A.4). Similarly, when $\alpha=\alpha_{0}=1-\kappa/8$, it follows from Lemma A.1 and Eq. (A.5) that the only positive solutions of (A.4) are $C\phi_{\alpha_{0}}$, with $C>0$.
Now consider $\alpha<1-\kappa/8$. As stated in Remark 3.15, $\Phi^{L}$ and $\Phi^{R}$ are $C^{2}$-solutions to (A.4) ($\Phi^{L},\Phi^{R}\in C^{2}((0,1))$ follows from the same argument as for $\Phi$ in Lemma 3.12 using Remark 2.8). Moreover, $0<\Phi^{L}(u),\Phi^{R}(u)\leq\Phi(u)$, for $u\in(0,1)$, and
$\Phi^{L}(0)=1,\quad\Phi^{L}(1)=0,\quad\text{and}\quad\Phi^{R}(0)=0,\quad\Phi^{R}(1)=1.$
Since $\Phi^{L}$ and $\Phi^{R}$ are linearly independent, any solution $\phi$ of (A.4) can be decomposed as $\phi=C_{L}\Phi^{R}+C_{R}\Phi^{R},$ for some real constants $C_{L}$ and $C_{R}$. If $C_{L}<0$, then $\phi(u)<0$ for $u>0$ small. So, for a positive solution $\phi$ we must have $C_{L}\geq 0$. Similarly, we deduce that $C_{R}\geq 0$, and clearly we may not have $C_{L}=C_{R}=0$. This finishes the proof. ∎
Appendix B Solutions of the separable equation (4.7)
Lemma B.1.
Let $C\in\mathbb{R}$ and consider the separable differential equation
$(8+2C\sin^{2}(\theta/2))f^{\prime}(\theta)=(f(\theta))^{2}+C+4.$ (B.1)
The only $C^{1}$ solutions of (B.1) defined on $(0,2\pi)$ are
  • if $C+4>0$,
[…]
    for any $D\in[-\pi/4,\pi/4],$
  • if $C+4=0$,
[…]
    for any $D\leq 0$ and $E\geq 0$,
  • if $C+4<0$,
    $f(\theta)=\sqrt{-C-4}\quad\text{and}\quad f(\theta)=-\sqrt{-C-4}.$
Proof.
Note that
$\frac{1}{8+2C\sin^{2}(\theta)}=\frac{1}{4+(C+4)\tan^{2}(\theta/2)}\frac{\mathrm{d}}{\mathrm{d}\theta}\tan(\theta/2).$
Thus, we can then rewrite (B.1) as
$\frac{\mathrm{d}f}{f^{2}+(C+4)}=\frac{\mathrm{d}\tan(\theta/2)}{4+(C+4)\tan^{2}(\theta/2)},$ (B.3)
under the assumptions $(f(\theta))^{2}\neq-(C+4),$ $\theta\neq\pi$ and $(C+4)\tan^{2}(\theta/2)\neq-4.$
Case 1: Suppose $C+4>0$. By the Picard-Lindelöf theorem Eq. (4.7) has a unique local solution $f$ around $\theta_{0}\in(0,2\pi)$ for each initial condition $f(\theta_{0})=f_{0}$. Integrating both sides of (B.3) gives
$\arctan(f(\theta)/\sqrt{C+4})=\begin{cases}\frac{\arctan(\frac{1}{2}\sqrt{C+4}\tan(\theta/2))}{2}+D-\pi/4&\theta\in(0,\pi)\\ \frac{\arctan(\frac{1}{2}\sqrt{C+4}\tan(\theta/2))}{2}+D+\pi/4&\theta\in(\pi,2\pi)\end{cases}$
for some $D$. Note that the left-hand side takes values in $(\-\pi/2,\pi/2)$ and that this is true for the right-hand side only if $D\in[-\pi/4,\pi/4]$. Thus the only global solutions of (B.1) in the case $C+4>0$ are the ones given in (B.2).
Case 2: Suppose $C+4=0$. Picard-Lindelöf guarantees existence and uniqueness of local solutions to the initial value problem away from $\theta=\pi$. The solutions on $(0,\pi)$ are of the form $f\equiv 0$ or
$f(\theta)=\frac{1}{D-\frac{1}{4}\tan(\theta/2)}$
for some $D\in\mathbb{R}$. In order for the latter to be defined on all of $(0,\pi)$ we require $D\leq 0$. Similarly, the solutions on $(\pi,2\pi)$ are of the form $f\equiv 0$ or
$f(\theta)=\frac{1}{E-\frac{1}{4}\tan(\theta/2)}$
for some $E\geq 0$. To find solutions on $(0,2\pi)$ we patch the solutions above. For this we note that the left and right derivatives of $f(\theta)=(D-\frac{1}{4}\tan(\theta/2))^{-1}$ at $\pi$ are both $2$, regardless of $D$.
Case 3: Finally, fix $C+4<0$ and let $\theta_{0}$ be the unique solution to $8+2C\sin^{2}(\theta/2)=0$ on $(0,\pi)$. Then Picard-Lindelöf guarantees existence and uniqueness of local solutions to the initial value problem around each $\theta\in(0,2\pi)\setminus\{\theta_{0},2\pi-\theta_{0}\}$. For the initial values $\pm\sqrt{-C-4}$ we have constant solutions. For the other initial values we integrate (B.3) and find solutions of the form
$\log\frac{|f(\theta)-\sqrt{-C-4}|}{|f(\theta)+\sqrt{-C-4}|}=\frac{1}{2}\log\frac{|\sqrt{-C-4}\tan(\theta/2)+2|}{|\sqrt{-C-4}\tan(\theta/2)-2|}+D,$
$D\in\mathbb{R}$. More concretely, if $\theta_{1}\in(0,2\pi)\setminus\{\theta_{0},2\pi-\theta_{0}\}$ and $f(\theta_{1})\in\mathbb{R}\setminus\{\pm\sqrt{-C-4}\}$ then
$f(\theta)=T\bigg(\frac{1}{2}\log\frac{|\sqrt{-C-4}\tan(\theta/2)+2|}{|\sqrt{-C-4}\tan(\theta/2)-2|}+D\bigg)$
for $\theta$ in the connected component of $(0,2\pi)\setminus\{\theta_{0},2\pi-\theta_{0}\}$ containing $\theta_{1}$, where $T$ is the inverse of $\log\frac{|x-\sqrt{-C-4}|}{|x+\sqrt{-C-4}|}$ on the connected component of $\mathbb{R}\setminus\{\pm\sqrt{-C-4}\}$ containing $f(\theta_{1})$. An explicit computation shows that all such solutions satisfy $|f^{\prime}(\theta)|\to\infty$ as $\theta\to\theta_{0}$ or $\theta\to 2\pi-\theta_{0}$ and therefore they can not be extended to $C^{1}$ solutions on $(0,2\pi).$
Acknowledgments.
  • We thank Mo Chen, Yu Feng, Chongzhi Huang, Eveliina Peltola, Fredrik Viklund, and Lu Yang for helpful discussions. We thank Minjae Park for the simulation in Figure 2. We thank the anonymous referees for their constructive comments on a previous version of the paper.
  • E.K. is supported by a grant from the Knut and Alice Wallenberg foundation. E.K. acknowledges the hospitality of the IHES where part of this work was carried out.
  • Y.W. is funded by the European Union (ERC, RaConTeich, 101116694) [^1] Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them..
  • H.W. is supported by Beijing Natural Science Foundation (JQ20001) and New Cornerstone Investigator Program 100001127. Part of this work was carried out during H.W.’s visit at IHES under the support of “K.C. Wong Education Foundation”. H.W. is partly affiliated with Yanqi Lake Beijing Institute of Mathematical Sciences and Applications, Beijing, China.
Conflict of interest. The authors have no competing interests to declare that are relevant to the content of this article.